### v1

# Runtime Analysis of the SMS-EMOA for Many-Objective Optimization

Abstract
The widely used multiobjective optimizer NSGA-II was recently proven to have considerable difficulties in many-objective optimization. In contrast, experimental results in the literature show a good performance of the SMS-EMOA, which can be seen as a steady-state NSGA-II that uses the hypervolume contribution instead of the crowding distance as the second selection criterion.
This paper conducts the first rigorous runtime analysis of the SMS-EMOA for many-objective optimization. To this aim, we first propose a many-objective counterpart, the $m$-objective $m$OJZJ problem, of the bi-objective OJZJ benchmark, which is the first many-objective multimodal benchmark used in a mathematical runtime analysis. We prove that SMS-EMOA computes the full Pareto front of this benchmark in an expected number of $O(M^{2}n^{k})$ iterations, where $n$ denotes the problem size (length of the bit-string representation), $k$ the gap size (a difficulty parameter of the problem), and $M=(2n/m-2k+3)^{m/2}$ the size of the Pareto front. This result together with the existing negative result on the original NSGA-II shows that in principle, the general approach of the NSGA-II is suitable for many-objective optimization, but the crowding distance as tie-breaker has deficiencies.
We obtain three additional insights on the SMS-EMOA. Different from a recent result for the bi-objective OJZJ benchmark, the stochastic population update often does not help for $m$OJZJ. It results in a $1/\Theta(\min\{Mk^{1/2}/2^{k/2},1\})$ speed-up, which is $\Theta(1)$ for large $m$ such as $m>k$. On the positive side, we prove that heavy-tailed mutation still results in a speed-up of order $k^{0.5+k-\beta}$. Finally, we conduct the first runtime analyses of the SMS-EMOA on the bi-objective OneMinMax and LOTZ benchmarks and show that it has a performance comparable to the GSEMO and the NSGA-II.
1 Introduction
The NSGA-II is the most widely-applied multiobjective evolutionary algorithm (MOEA). Non-dominated sorting and crowding distance are its two major features differentiating it from basic MOEAs such as the GSEMO or the $(\mu+1)$ SIBEA. Zheng, Liu, and Doerr [ZLD22] conducted the first runtime analysis of the NSGA-II (see [ZD23a] for the journal version). This work quickly inspired many interesting follow-up results in bi-objective optimization [ZD22, BQ22, DQ23a, DQ23b, DQ23c, DOSS23a, DOSS23b, ${}^{+}$23]. In contrast to these positive results for two objectives, Zheng and Doerr [ZD23b] proved that for $m\geq 3$ objectives the NSGA-II needs at least exponential time (in expectation and with high probability) to cover the full Pareto front of the $m$-objective OneMinMax benchmark, a simple many-objective version of the basic OneMax problem where all search points are Pareto optimal. They claimed that the main reason for this low efficiency is the independent computation of the crowding distance in each objective.
A very recent work showed that the NSGA-III, a successor algorithm of the NSGA-II aimed at better coping with many objectives, can efficiently solve the $3$-objective OneMinMax problem [WD23]. Since apparently practitioners much prefer the NSGA-II (more than 17,000 citations on Google scholar only in the last five years) over the NSGA-III (5,035 citations since its publication in 2013), it remains an interesting question whether there are variants of the NSGA-II which better cope with many objectives.
With the SMS-EMOA, an interesting variant of the NSGA-II was proposed by eume, Naujoks, and Emmerich [BNE07]. This algorithm is a steady-state variant of the NSGA-II (that is, in each iteration only a single offspring is generated and possibly integrated into the population) that further replaces the crowding distance as secondary selection criterion with the classic hypervolume contribution. Many empirical works (see the almost 2,000 papers citing [BNE07]) confirmed the good performance of the SMS-EMOA for many-objective optimization. The only mathematical runtime analysis of the SMS-EMOA was conducted only very recently by Bian et al. [BZLQ23]. They proved that its runtime for the bi-objective OJZJ problem is $O(n^{k+1})$ and proposed a stochastic population update mechanism which was proven to achieve an expected runtime of $O(n^{k+1}\min\{1,n/2^{k/4}\})$.
Our Contributions: This paper conducts the first mathematical runtime analysis of the SMS-EMOA for more than two objectives. We first define the $m$OJZJ benchmark, an $m$-objective counterpart of the bi-objective OJZJ problem [DZ21], which is the problem analyzed in the only previous runtime analysis for the SMS-EMOA [BZLQ23]. We note that the $m$OJZJ problem is the first multimodal many-objective benchmark proposed for a theoretical analysis, to the best of our knowledge. We prove that the SMS-EMOA covers the full Pareto front of this benchmark in an expected number of $O(M^{2}n^{k})$ iterations, where $n$ is the problems size, $k$ the gap size (a difficulty parameter of the problem), $m$ the number of objectives, and $M=(2n/m-2k+3)^{m/2}$ the size of the Pareto front.
We recall that the original NSGA-II needs at least exponential time to optimize the $m$OneMinMax problem, which is a special case of $m$OJZJ with gap size $k=1$. Since the SMS-EMOA employs non-dominated sorting, but replaces the crowding distance in the original NSGA-II by the hypervolume contribution, our result in a similar fashion as the analysis of the NSGA-III in [WD23] suggests that the general approach of the NSGA-II is suitable for many-objective optimization and that it is only the crowding distance as tie-breaker which is not appropriate for more than two objectives.
We then analyze whether the better performance of the SMS-EMOA on the bi-objective OJZJ problem achieved via a new stochastic population update [BZLQ23] extends to the $m$-objective $m$OJZJ problem. Unfortunately, we shall observe that only a speed-up of order $1/\Theta(\min\{Mk^{1/2}/2^{k/2},1\})$ is obtained, which is $\Theta(1)$ when $m$ is large, e.g., $m>k$.
On the positive side, we show that the advantage of heavy-tailed mutation is preserved. We analyze the SMS-EMOA with heavy-tailed mutation on the $m$OJZJ benchmark and prove that a speed-up of order $k^{0.5+k-\beta}$ is achieved. This is the same speed-up as observed for single-objective and bi-objective Jump problems [DLMN17, DZ21, ZD23c]. We note that this is the first theoretical work to support the usefulness of heavy-tailed mutation in many-objective optimization.
Finally, since so far the performance of the SMS-EMOA was only analyzed on the bi-objective OJZJ problem [BZLQ23], we conduct mathematical runtime analyses of the SMS-EMOA also on the two most prominent bi-objective benchmarks OneMinMax and LOTZ. We prove that the SMS-EMOA finds the Pareto fronts of these benchmarks in an expected number of iterations of at most $2e(n+1)n(\ln n+1)$ for OneMinMax and at most $2en^{2}(n+1)$ for LOTZ. These are the same asymptotic runtimes (in terms of fitness evaluations) as known for the GSEMO and the NSGA-II.
This paper is organized as follows. Section 2 briefly introduces some basic concepts for the many-objective optimization and the mathematical notations that will be used in this paper. The many-objective OJZJ benchmark is proposed in Section 3. Sections 4 to 6 analyze the runtime of the original SMS-EMOA, the SMS-EMOA with stochastic population update, and the one with heavy-tailed mutation on this many-objective benchmark. The runtime for two well-known bi-objective functions is conducted in Section 7. Section 8 concludes this work.
2 Preliminaries
In many-objective optimization, one tries to find good solutions for a problem containing several, but at least three objectives. This work considers pseudo-Boolean maximization, hence our problem is described by a function $f=(f_{1},\dots,f_{m}):\{0,1\}^{n}\rightarrow\mathbb{R}^{m}$. Here $m>2$ is the number of objectives and $n\in\mathbb{N}$ is called the problem size. As in bi-objective optimization (that is, the case $m=2$), usually not all solutions are comparable. We say that a solution $x\in\{0,1\}^{n}$ dominates a solution $y$, denoted as $x\succ y$, if $f_{i}(x)\geq f_{i}(y)$ for all $i\in\{1,\dots,m\}$, and at least one of these inequalities is strict. For a given many-objective function $f$, we say that $x\in\{0,1\}^{n}$ is a Pareto optimum if and only if there is no $y\in\{0,1\}^{n}$ that dominates $x$ w.r.t. $f$. The set of all Pareto optima is called Pareto set, and the set of all function values of the Pareto optima is called Pareto front. The goal for an MOEA is to find the full Pareto front as much as possible, that is, to compute a not too large set $S$ of solutions such that $f(S)$ equals or approximates well the Pareto front. As common in the mathematical runtime analysis of MOEAs, we call the runtime of an algorithm the number of function evaluations until its population $P$ covers the full Pareto front, that is, $f(P)$ contains the Pareto front. We refer to [NW10, AD11, Jan13, ZYQ19, DN20] for general introductions to the mathematical runtime analysis of evolutionary algorithms, and to [Bro11] for a discussion of runtime analyses of MOEAs.
Mathematical Notations. For $a,b\in\mathbb{N}$ with $a\leq b$, we denote the set of integers in the interval $[a,b]$ by $[a..b]=\{a,a+1,\dots,b\}$. For a point $x\in\{0,1\}^{n}$, let $x_{[a..b]}:=(x_{a},x_{a+1},\dots,x_{b})$ and $|x|_{1}$ be the number of ones in $x$.
3 The $m$-Objective Jump Function
As mentioned earlier, the NSGA-II was proven to have enormous difficulties in optimizing many-objective problems [ZD23b]. In that paper, the $m$-objective counterpart $m$OneMinMax of the bi-objective OneMinMax benchmark was proposed and analyzed. Since the only runtime work of the SMS-EMOA so far [BZLQ23] analyzed the bi-objective jump problem OJZJ, whose special case with gap size $k=1$ is (essentially) the OneMinMax problem, we shall propose and work with an $m$-objective version of OJZJ as well. Again, its special case $k=1$ will be (essentially) equal to the $m$OneMinMax problem. With this, our results are comparable both the ones in [ZD23b] and [BZLQ23].
3.1 $m$OJZJ
We first recall the definition of the $m$OneMinMax problem. For the ease of presentation, we only consider even numbers $m$ of objectives here. In the $m$OneMinMax problem, the bit string (of length $n$) is divided into $m/2$ blocks of equal length $2n/m$. On each of these, a bi-objective OneMinMax problem is defined. We note that this general block construction goes back to the seminal paper of [LTZ04].
Definition 1 ([ZD23b]).
Let $m$ be the number of objectives and be even, and the problem size $n$ be a multiple of $m/2$. Let $n^{\prime}=2n/m\in\mathbb{N}$. For any $x=(x_{1},\dots,x_{n})$, the $m$-objective function mOneMinMax $f:\{0,1\}^{n}\rightarrow\mathbb{R}^{m}$ is defined by
$f_{i}(x)$ =$=\begin{cases}\textsc{OneMax}(\bar{x}_{[\frac{i-1}{2}n^{\prime}+1..\frac{i+1}{2}n^{\prime}]}),&\text{if $i$ is odd}\\ \textsc{OneMax}({x}_{[\frac{i-2}{2}n^{\prime}+1..\frac{i}{2}n^{\prime}]}),&\text{else,}\end{cases}$
where $\bar{x}=(1-x_{1},\dots,1-x_{n})$ and the function $\textsc{OneMax}:\{0,1\}^{n^{\prime}}\rightarrow\mathbb{R}$ is defined by
$\textsc{OneMax}(y)=\sum_{i=1}^{n^{\prime}}y_{i}$
for any $y\in\{0,1\}^{n^{\prime}}$.
In order to include $m$OneMinMax as a special case of $m$OJZJ, we define the $m$-objective OJZJ in a similar manner, that is, we divide the $n$ bit positions into $m/2$ blocks and define a OJZJ problem in each block.
Definition 2.
[…]
$f_{i}(x)$ =$=\begin{cases}\textsc{Jump}_{n^{\prime},k}(x_{[\frac{i-1}{2}n^{\prime}+1..\frac{i+1}{2}n^{\prime}]}),&\text{if $i$ is odd}\\ \textsc{Jump}_{n^{\prime},k}(\bar{x}_{[\frac{i-2}{2}n^{\prime}+1..\frac{i}{2}n^{\prime}]}),&\text{else,}\end{cases}$
where $\bar{x}=(1-x_{1},\dots,1-x_{n})$ and the function $\textsc{Jump}_{n^{\prime},k}:\{0,1\}^{n^{\prime}}\rightarrow\mathbb{R}$ is defined by
[…]
for any $y\in\{0,1\}^{n^{\prime}}$.
We note that the function $\textsc{Jump}_{n^{\prime},k}$ used in the definition above is the famous Jump benchmark. It was first defined by Droste, Jansen, and Wegener [DJW02] and has quickly become the most employed multimodal benchmark in the theory of randomized search heuristics, leading to many fundamental results on how these algorithms cope with local optima [JW02, ${}^{+}$18, Doe21, BBD21, HS22, RW22, Wit23, DEJK24].
[…]
3.2 Characteristics
We now give more details on this $m$OJZJ function. Let $B_{i}:=[(i-1)n^{\prime}+1..in^{\prime}]$, $i\in[1..m/2]$, be the $i$-th block in the partition of $n$ bit positions. From Definition 2, we know that the bit values in the block $B_{i}$ only influence the objectives $f_{2i-1}$ and $f_{2i}$. Figure 1 plots the objective values of $f_{2i-1}$ and $f_{2i}$ relative to the number of ones in this block. Obviously, $m$OJZJ is multimodal with respect to the definition of multimodality of multiobjective problems in [DZ21].
Figure 1: The objective values of $f_{2i-1}$ and $f_{2i}$ in $m$OJZJ w.r.t. $|x_{B_{i}}|_{1}$, the number of ones in the block $B_{i}$.
It is not difficult to see that the Pareto set is $S^{*}=\{x\in\{0,1\}^{n}\mid|x_{B_{i}}|_{1}\in[k..n-k]\cup\{0,n\},i\in[1..m/2]\}$. Hence, the Pareto front is
[…]
${}\mid a_{1},\dots,a_{m/2}\in[2k..n^{\prime}]\cup\{k,n^{\prime}+k\}\}$
and the Pareto front size is $M:=|F^{*}|=(n^{\prime}-2k+3)^{m/2}$.
We recall at this point that we defined the runtime of an MOEA as the time to cover the full Pareto front $F^{*}$, that is, the first time to reach a population $P$ with $F^{*}\subseteq f(P)$.
Note that for $k>1$, not all points are Pareto optimal and not all sets of mutually non-dominated points are the subsets of the Pareto front. However, we have the following result showing that the maximum number of mutually non-dominated points is at most the size of the Pareto front.
Lemma 3.
Let $k\leq n^{\prime}/2$ and $P$ be a set of solutions such that any two are not weakly dominating (w.r.t. $m$OJZJ). Then $|P|\leq M$.
Proof.
Note that all objective values of any non-Pareto optimal solution are less than $2k$. Since each Pareto optimum has its all objective values of at least $2k$, we know that any non-Pareto optimal solution is dominated by all Pareto optima excluding the ones with some block all zeros or all ones (called extreme-block Pareto optima). Hence, $P$ excluding extreme-block Pareto optima, denoted as $P^{\prime}$, cannot include both Pareto optima and non-Pareto optimal solutions. If $P$ includes only Pareto optima, then we know that $|P|\leq M$.
We now consider that $P^{\prime}$ includes only non-Pareto optimal solutions, that is, for any $x\in P^{\prime}$, there exists a block $B_{i}:=[(i-1)n^{\prime}+1..in^{\prime}],i\in[1..m/2]$ such that $|x_{B_{i}}|_{1}\in[1..k-1]\cup[n^{\prime}-k+1..n^{\prime}-1]$. Note that for any block $B_{i}$, any two solutions $x,y$ are comparable w.r.t. $(f_{2i-1},f_{2i})$ if $(|x_{B_{i}}|_{1}\in[1..k-1],|y_{B_{i}}|_{1}\in[1..k-1])$ or $(|x_{B_{i}}|_{1}\in[n^{\prime}-k+1..n^{\prime}-1],|y_{B_{i}}|_{1}\in[n^{\prime}-k+1..n^{\prime}-1])$. Hence, with the bit values on other blocks fixed, there are at most two solutions in $P^{\prime}$ that are not weakly dominated. We also note that $(f_{2i-1}(z),f_{2i}(z))>(f_{2i-1}(x),f_{2i}(x))$ holds if $(z_{B_{i}}=1^{n^{\prime}},|x_{B_{i}}|_{1}\in[1..k-1])$ or $(z_{B_{i}}=0^{n^{\prime}},|x_{B_{i}}|_{1}\in[n^{\prime}-k+1..n^{\prime}-1])$. Hence, with the bit values on other blocks fixed, there are at most two solutions in $P$ that are not weakly dominated. Then $m/2$ blocks results in at most $2^{m/2}$ non-weakly-dominated solutions, that is, $|P|\leq 2^{m/2}<3^{m/2}\leq M$. ∎
For ease of reading of the proof sketches in the following sections, we call a Pareto optimum $x$ an inner Pareto optimum if for all blocks $B_{i}=[(i-1)n^{\prime}+1..in^{\prime}],i\in[1..m/2]$, we have $x_{B_{i}}\notin\{0^{n^{\prime}},1^{n^{\prime}}\}$. We call $u=(u_{1},n^{\prime}+2k-u_{1},\dots,u_{m/2},n^{\prime}+2k-u_{m/2})$ and $v=(v_{1},n^{\prime}+2k-v_{1},\dots,v_{m/2},n^{\prime}+2k-v_{m/2})$ neighbors if there exists $i\in[1..m/2]$ such that $u_{i}-v_{i}=1$ and $u_{j}=v_{j}$ for all $j\neq i$.
4 SMS-EMOA Can Optimize $m$OJZJ
As discussed before, the original NSGA-II with non-dominated sorting and crowding distance needs at least exponential runtime to cover the full Pareto front of the $m$OneMinMax [ZD23b], which is a special case of $m$OJZJ. Motivated by the question whether there is a simple efficient variant of NSGA-II for many objectives, this section will analyze the runtime of the SMS-EMOA, which still uses the non-dominated sorting but replaces the crowding distance by hypervolume measure.
4.1 Algorithm Description
The SMS-EMOA is a steady-state variant (that is, the offspring size is smaller than the parent size) of the NSGA-II. It maintains a fixed population size $\mu$. For each iteration, only one offspring $x^{\prime}$ is generated from the parent population $P_{t}$. From the $\mu+1$ individuals in the combined population $R_{t}=P_{t}\cup\{x^{\prime}\}$, one individual will be removed. It first sorts the combined population into several fronts $F_{1},\dots,F_{i^{*}}$, where $F_{i}$ contains all mutually non-dominated individuals in $R_{t}\setminus(\bigcup_{j=1}^{i-1}F_{j})$. This is called non-dominated sorting. Different from the original NSGA-II that uses the crowding distance as the secondary criterion to remove the individual, SMS-EMOA removes the one with the smallest hypervolume contribution in the critical front, that is, $F_{i^{*}}$. The hypervolume of the set $S$ w.r.t. a reference point $r$ is defined as
$\operatorname{HV}_{r}(S)=\mathcal{L}\mathopen{}\mathclose{{}\left(\,\bigcup_{u\in S}\{h\in\mathbb{R}^{m}\mid r\leq h\leq f(u)\}}\right),$
where $\mathcal{L}$ is the Lebesgue measure. The hypervolume contribution of an individual $x\in F_{i^{*}}$ is calculated by
$\Delta_{r}(x,F_{i^{*}}):=\operatorname{HV}_{r}(F_{i^{*}})-\operatorname{HV}_{r}(F_{i^{*}}\setminus\{x\}).$
Algorithm 1 shows the general procedure of the SMS-EMOA.
1:Initialize $P_{0}$ by generating $\mu$ solutions uniformly at random from $\{0,1\}^{n}$
2:for $t=0,1,2,\dots,$ do
3:   Select a solution $x$ uniformly at random from $P_{t}$
4:   Generate $x^{\prime}$ by flipping each bit of $x$ independently with probability $1/n$
5:   %Original Population Update
6:   Use fast-non-dominated-sort() [DPAM02] to divide $R_{t}=P_{t}\cup\{x^{\prime}\}$ into $F_{1},\dots,F_{i^{*}}$
7:   Calculate $\Delta_{r}(z,F_{i^{*}})$ for all $z\in F_{i^{*}}$ and find $D=\arg\min_{z\in F_{i^{*}}}{\Delta_{r}(z,F_{i^{*}})}$
8:   Uniformly at random pick $z^{\prime}\in D$ and $P_{t+1}=R_{t}\setminus\{z^{\prime}\}$
9:end for
Algorithm 1 SMS-EMOA
4.2 Runtime of SMS-EMOA
We now establish the runtime of the SMS-EMOA to cover the full Pareto front of the $m$OJZJ. We first give the following lemma to show that for a large enough population size, the function value of any point with rank one in the combined population will survive to the next population. The key of the proof is that all $F_{1}$ points will survive if there is a point with a larger rank, and that at least two points have the same function value if all points are in $F_{1}$ and the population size is large enough.
Lemma 4.
Consider using SMS-EMOA with $\mu\geq M$ to optimize the $m$OJZJ problem. For any $x\in F_{1}$ for the current generation $t$, we have that there is $y$ with $f(y)=f(x)$ in $P_{t+1}$. Specifically, any Pareto front point will be maintained in all future populations once it is reached.
Proof.
Let $x\in F_{1}$. If $\cup_{i\geq 2}F_{i}\neq\emptyset$, then the removal will happen in $\cup_{i\geq 2}F_{i}$, and $x\in P^{t+1}$.
If all solutions are in $F_{1}$, from Lemma 3, we know that $|f(F_{1})|\leq M$. Hence, there exist $z_{1},z_{2}\in F_{1}$ such that $f(z_{1})=f(z_{2})$. Let $S_{x}=\{z\in R_{t}\mid f(x)=f(z)\}$. Then $|S_{x}|\geq 1$. If $|S_{x}|=1$, then $\Delta(x,F_{1})>0=\Delta(z_{1},F_{1})$, and thus $x\in P_{t+1}$. If $|S_{x}|>1$, since the survival selection only removes one solution, then there exists $y\in S_{x}$, such that $y\in P_{t+1}$.
Since any Pareto optimum is always in $F_{1}$, the latter part of this lemma holds. ∎
Recall that the inner Pareto optimum is a Pareto optimum with $x_{B_{i}}\notin\{0^{n^{\prime}},1^{n^{\prime}}\}$ for all blocks $B_{i}=[(i-1)n^{\prime}+1..in^{\prime}],i\in[1..m/2]$, as defined before. Since not all points are inner Pareto optimum, the following lemma establishes the runtime to reach one inner Pareto optimum for the first time. The key of the proof is that at most $k$ specific bit flipping of a point that is not an inner Pareto optimum will result in an inner Pareto optimum.
Lemma 5.
Let $k\leq n^{\prime}/2$. Consider using SMS-EMOA with $\mu\geq M$ to optimize the $m$OJZJ problem. Then after at most $\frac{1}{2}em^{k+1}k^{k}$ iterations in expectation, the population (and also the populations afterwards) contains at least one inner Pareto optimum.
Proof.
This lemma is already proved if the initial population has at least one inner Pareto optimum. Now we consider that the initial population contains no inner Pareto optimum. Consider any population $P$ without inner Pareto optimum. Then we know that there exists a block $B_{i}:=[(i-1)n^{\prime}+1..in^{\prime}],i\in[1..m/2]$ such that for all $x\in P$, $A_{i}:=|x_{B_{i}}|_{1}\in[0..k-1]\cup[n^{\prime}-k+1..n^{\prime}]$. We regard it a success if $y$ with $|y_{B_{i}}|_{1}\in[k..n^{\prime}-k]$ is generated. Then the probability to generate such $y$ from $x$ with $A_{i}<k$ is at least
$\binom{n^{\prime}-A_{i}}{k-A_{i}}$ ${}\mathopen{}\mathclose{{}\left(1-\frac{1}{n}}\right)^{n-(k-A_{i})}\mathopen{}\mathclose{{}\left(\frac{1}{n}}\right)^{k-A_{i}}\geq\frac{1}{e}\binom{n^{\prime}-A_{i}}{k-A_{i}}\mathopen{}\mathclose{{}\left(\frac{1}{n}}\right)^{k-A_{i}}\geq\frac{1}{e}\mathopen{}\mathclose{{}\left(\frac{n^{\prime}-A_{i}}{n(k-A_{i})}}\right)^{k-A_{i}}$
$\geq{}\frac{1}{e}\mathopen{}\mathclose{{}\left(\frac{n^{\prime}-k}{nk}}\right)^{k}\geq\frac{1}{e}\mathopen{}\mathclose{{}\left(\frac{n^{\prime}}{2nk}}\right)^{k}=\frac{1}{e}\mathopen{}\mathclose{{}\left(\frac{1}{mk}}\right)^{k},$
where the last inequality uses $k\leq n^{\prime}/2$ and the last equality uses $n^{\prime}=2m/n$. We also know that the probability to generate such $y$ from $x$ with $A_{i}>n^{\prime}-k$ is at least
$\binom{A_{i}}{A_{i}-(n^{\prime}-k)}$ ${}\mathopen{}\mathclose{{}\left(1-\frac{1}{n}}\right)^{n-(A_{i}-(n^{\prime}-k))}\mathopen{}\mathclose{{}\left(\frac{1}{n}}\right)^{A_{i}-(n^{\prime}-k)}\geq\frac{1}{e}\mathopen{}\mathclose{{}\left(\frac{A_{i}}{n(A_{i}-(n^{\prime}-k))}}\right)^{A_{i}-(n^{\prime}-k)}$
$\geq{}\frac{1}{e}\mathopen{}\mathclose{{}\left(\frac{n^{\prime}-k}{nk}}\right)^{k}\geq\frac{1}{e}\mathopen{}\mathclose{{}\left(\frac{n^{\prime}}{2nk}}\right)^{k}=\frac{1}{e}\mathopen{}\mathclose{{}\left(\frac{1}{mk}}\right)^{k},$
where the last two inequalities use $n^{\prime}-k<A_{i}\leq n^{\prime}$ and $k\leq n^{\prime}/2$ respectively, and the last equality uses $n^{\prime}=2m/n$. Note that $A_{i}\in[0..k-1]\cup[n^{\prime}-k+1..n^{\prime}]$ holds for all $x\in P$ before such $y$ is generated. Then we know that $y$ will be generated in expected $e(mk)^{k}$ iterations. Since such $y$ cannot be dominated by any solution in $P$, we know that $y\in F_{1}$. From Lemma 4, we know that once $y$ is generated, $f(y)$ will be covered by the next population and by all future populations, and thus future populations will contain at least one solution $z$ with $|z_{B_{i}}|_{1}\in[k..n^{\prime}-k]$.
Since an inner Pareto optimum is reached when $y$ with $|y_{B_{1}}|_{1},\dots,|y_{B_{m/2}}|_{1}\in[k..n^{\prime}-k]$ is generated, we know that the expected iterations to reach an inner Pareto optimum is at most $\frac{m}{2}e(mk)^{k}=\frac{1}{2}em^{k+1}k^{k}$. ∎
We consider the next stage of covering all inner Pareto front points. The following lemma shows the runtime. The key of the proof is that one uncovered Pareto front point can be easily generated from a neighbor Pareto front if it is covered.
Lemma 6.
Let $k\leq n^{\prime}/2$. Consider using SMS-EMOA with $\mu\geq M$ to optimize the $m$OJZJ problem. Assume that the current population contains at least one inner Pareto optimum. Then after at most $en\mu M$ iterations in expectation, all inner Pareto front points are covered.
Proof.
Let $u\in F^{*}$ be an uncovered inner Pareto front point that has some neighbor covered by the population $P$. Since such a neighbor is also a Pareto front point, we know that it is maintained in all future generations. Since the probability to generate an offspring with its function value of $u$ is at least $\mathopen{}\mathclose{{}\left(1-\frac{1}{n}}\right)^{n-1}\frac{1}{n\mu}\geq\frac{1}{en\mu}$, we know that the expected iterations to witness such an offspring is at most $n\mu$ in expectation. From Lemma 4, we know that once such an offspring is generated, $u$ will be covered in all future populations. Noting that there are at most $M$ inner Pareto front points, we know that in expected at most $en\mu M$ iterations, all Pareto front points are covered. ∎
The last stage is to cover the remaining Pareto front points. The following lemma shows the runtime. The key of the proof is that we can divide the Pareto front into several levels and any individual in the current $i$-th level can be easily generated from a point in the $(i-1)$-th level.
Lemma 7.
Let $k\leq n^{\prime}/2$. Consider using SMS-EMOA with $\mu\geq M$ to optimize the $m$OJZJ problem. Assume that the current population covers all inner Pareto front points. Then after at most $eM\mu n^{k}$ iterations in expectation, the full Pareto front is covered.
Proof.
We call $(k,n^{\prime}+k)$ or $(n^{\prime}+k,k)$ extreme block. We divide the Pareto front points into $m/2+1$ levels, where each point in $i$-th level contains $i$ extreme blocks. Obviously, all inner Pareto front points belong to the 00-th level. Note that any point in $j$-th level ($j\geq 1$) can be generated by a certain point in $(j-1)$-th level via flipping specific $k$ bits with other bits unchanged. Hence, the probability to generate any point in $j$-th level condition on that all $(j-1)$-th level points are covered is at least
$\frac{1}{\mu}\mathopen{}\mathclose{{}\left(1-\frac{1}{n}}\right)^{n-k}\mathopen{}\mathclose{{}\left(\frac{1}{n}}\right)^{k}\geq\frac{1}{e\mu n^{k}}.$
From Lemma 4, we know that such a Pareto front point is reached in at most $e\mu n^{k}$ in expectation, and will survive in all future populations. Consider the process of covering all non-inner Pareto front points in the order of covering the $1$-st level, …, $(m/2)$-th level. Since the 00-th level points are already covered from our assumption, and since there are at most $M$ non-inner Pareto front points, we know that the remaining levels are covered in at most $eM\mu n^{k}$ iterations in expectation. ∎
Summing up the runtime of all stages from Lemmas 5 to 7, we have the following theorem for the runtime of the full coverage of the Pareto front.
Theorem 8.
Let $k\leq n^{\prime}/2$. Consider using SMS-EMOA with $\mu\geq M$ to optimize the $m$OJZJ problem. Then after at most $\frac{1}{2}em^{k+1}k^{k}+n\mu M+eM\mu n^{k}$ iterations ($\mu+\frac{1}{2}em^{k+1}k^{k}+n\mu M+eM\mu n^{k}$ function evaluations) in expectation, the full Pareto front is covered.
Comparing with the NSGA-II’s exponential lower bound to cover the special case of $m$OJZJ, Theorem 8 shows the efficiency of the SMS-EMOA for many objectives, and indicates that for the NSGA-II, the non-dominated sorting is cool, but the crowding distance as the secondary criterion is deficient.
4.3 Runtime of GSMEO
For an overall comparison, we also show the runtime of the GSMEO on the $m$OJZJ. The GSEMO is a multiobjective counterpart of the $(1+1)$ EA. The initial population contains a uniformly and randomly generated solution. Then in each iteration, a solution is uniformly at random picked from the population to generate an offspring with standard bit-wise mutation. If this offspring cannot be dominated by any solution in the population, it will enter into the population and removes all solutions weakly dominated by it. It is not difficult to see all solutions in the population of the GSEMO are mutually non-dominated. Hence, the population size of the GSEMO is at most $M$ from Lemma 3.
The proof ideas are similar to Lemmas 5 to 7. The only difference is that SMS-EMOA needs Lemma 4 to ensure that all function values covered in $F_{1}$ will remain covered by the next population, but for GSEMO, all such function values are surely covered by the next population as $F_{1}$ contains all mutually non-dominated individuals whose function values will not be removed in the survival selection. Together with the upper bound $M$ of the population size, we have the following theorem.
Theorem 9.
Let $k\leq n^{\prime}/2$. Consider using GSEMO to optimize the $m$OJZJ problem. Then after at most $\frac{1}{2}em^{k+1}k^{k}+nM^{2}+eM^{2}n^{k}$ iterations in expectation, the full Pareto front is covered.
5 Reduced Impact of Stochastic Population Updates
Bian et al. [BZLQ23] proposed a stochastic population update mechanism for the SMS-EMOA and proved that, somewhat unexpectedly given the state of the art, it can lead to significant performance gains. More precisely, they proved that the classic SMS-EMOA with reasonable population size $\mu=\Theta(n)$ solves the OJZJ problem in an expected number of $O(n^{k+2})\cap\Omega(n^{k+1})$ function evaluations, whereas for the SMS-EMOA with stochastic population update, $O(n^{k+2}\min\{1,2^{-k/4}n\})$ suffice. Hence, for $k=\omega(\log n)$, a super-polynomial speed-up is shown.
In this section, we extend the analysis of [BZLQ23] to $m$ objectives. Unfortunately, we will observe that the larger population sizes necessary here reduce the impact of the stochastic population update. As in [BZLQ23], we have no proven lower bounds for the SMS-EMOA with stochastic population update, but our upper-bound proofs suggest that the reduced effect of the stochastic population update on the runtime guarantee is real, that is, the impact on the true runtime is diminishing with the larger population sizes necessary in the many-objective setting.
5.1 Stochastic Population Update
The rough idea of the stochastic selection proposed by Bian et al. [BZLQ23] is that a random half of the individuals survive into the next generation regardless of their quality. The individual to be discarded is chosen from the remaining individuals according to non-dominated sorting and hypervolume contribution as in the classic SMS-EMOA. This approach resembles the random mixing of two acceptance operators in the Move Acceptance Hyper-Heuristic studied recently [LOW23, DDLS23].
More precisely, after generating the offspring $x^{\prime}$, the SMS-EMOA with stochastic selection chooses from $P_{t}\cup\{x^{\prime}\}$ a set $R^{\prime}$ of $\lfloor(\mu+1)/2\rfloor$ solutions randomly with replacement. The individual $z^{\prime}$ to be removed from $P_{t}\cup\{x^{\prime}\}$ is then determined via non-dominated sorting and hypervolume contribution applied to $R^{\prime}$ only. We note that with this mechanism, even the worst solution can enter the next generation with probability at least $1/2$.
5.2 Runtime
To establish the runtime result of the SMS-EMOA with stochastic population update, we separate the key facts into several lemmas for a clear demonstration. The following two lemmas 10 and 11 respectively show the general survival probability of each individual in the combined population, and the survival of the individual with rank one.
Lemma 10.
Consider using SMS-EMOA with stochastic population update to optimize the $m$OJZJ problem. For any $x\in R_{t}$ for the current generation $t$, we have that there is $y$ with $f(y)=f(x)$ in $P_{t+1}$ with probability at least $1/2$.
Proof.
We simply calculate the probability of $x$ being removed after the population update. Since $x$ is possible to be removed only if it is selected in $R^{\prime}$, we know that it will be removed with probability at most $\lfloor|R^{\prime}|/2\rfloor/|R^{\prime}|\leq 1/2$, and hence $f(x)$ will be maintained in $P_{t+1}$ with probability at least $1/2$. ∎
Lemma 11.
Consider using SMS-EMOA with stochastic population update and with $\mu\geq 2(M+1)$ to optimize the $m$OJZJ problem. For any $x\in F_{1}$ for the current generation $t$, we have that there is $y$ with $f(y)=f(x)$ in $P_{t+1}$. Specifically, any Pareto front point will be maintained in all future populations once it is reached.
Proof.
Let $x\in F_{1}$. If $\cup_{i\geq 2}F_{i}\neq\emptyset$, then the removal will happen in $\cup_{i\geq 2}F_{i}$, and $x\in P^{t+1}$.
If all solutions are in $F_{1}$, from Lemma 3, we know that $|f(F_{1})|\leq M$. Since there are $|R^{\prime}|\geq\lfloor(\mu+1)/2\rfloor\geq M+1>M$, we know that there exist $z_{1},z_{2}\in F_{1}$ such that $f(z_{1})=f(z_{2})$. Let $S_{x}=\{z\in R^{\prime}_{t}\mid f(x)=f(z)\}$. Then $|S_{x}|\geq 1$. If $|S_{x}|=1$, then $\Delta(x,F_{1})>0=\Delta(z_{1},F_{1})$, and thus $x\in P_{t+1}$. If $|S_{x}|>1$, since the survival selection only removes one solution, then there exists $y\in S_{x}$, such that $y\in P_{t+1}$.
Since any Pareto optimum is always in $F_{1}$, the latter part of this lemma holds. ∎
Recalling the proof in Lemmas 5 and 6, we know that these lemmas hold merely relies on that $f(F_{1})$ value will be maintained once it is reached. Hence, from Lemma 11, we have the following results for obtaining a population containing at least one inner Pareto optimum, and for covering all inner Pareto front points starting from a population with at least one inner Pareto optimum.
Lemma 12.
Let $k\leq n^{\prime}/2$. Consider using SMS-EMOA with stochastic population update and with $\mu\geq 2(M+1)$ to optimize the $m$OJZJ problem. Then
  • after at most $\frac{1}{2}em^{k+1}k^{k}$ iterations in expectation, the population (and also the populations afterwards) contains at least one inner Pareto optimum;
  • after at most $en\mu M$ iterations in expectation, all inner Pareto front points are covered starting from a population with at least one inner Pareto optimum.
Now we consider the runtime for the full coverage of the Pareto front after all inner Pareto front points are covered. Similar to the proof of Lemma 7 for the original SMS-EMOA, we partition the Pareto front points into $m/2+1$ levels based on the number of extreme blocks inside. The difference here is that we will consider the effect of the gap points, similar to [BZLQ23].
Lemma 13.
Let $k\leq n^{\prime}/2$. Consider using SMS-EMOA with stochastic population update and with $\mu\geq 2(M+1)$ to optimize the $m$OJZJ problem. Assume that the current population covers all inner Pareto front points. Then after at most $O(\min\{\frac{\mu k^{1/2}}{2^{k/2}},1\}M\mu n^{k})$ iterations in expectation, the full Pareto front is covered.
Proof.
We use the same definitions of the $j$-th ($j\in[0..m/2]$) level where each Pareto front point contains $j$ extreme blocks, and note that 00-th level points are covered. Let $P_{t}$ be the current population. We consider generating any point in $j$-th level condition on that all $(j-1)$-th level points are covered. Let $v$ be an uncovered Pareto front point in the $j$-th level. Then there exists a block $B$ that $v$ has the value of $(k,n^{\prime}+k)$ or $(n^{\prime}+k,k)$. W.l.o.g., let the block value of $v$ is $v_{B}=(k,n^{\prime}+k)$. Besides, there exists a set of points $U$ in the $(j-1)$-th level such that each $u\in U$ is the same as $v$ except that $u$ has the value of $(a,n^{\prime}+2k-a)$ with $a\in[2k..n^{\prime}]$ in the above block $B$. Let $S_{v}=\{f(z)\mid z\in P,z\text{\leavevmode\nobreak\ is the same as\leavevmode\nobreak\ }v\text{\leavevmode\nobreak\ except\leavevmode\nobreak\ }z_{B}\neq(k,n^{\prime}+k)\}$. Obviously, $y:=\min_{z\in S_{v}}z_{B}\leq 2k$. Let $A_{1}$ denote the event $k<y\leq\frac{3}{2}k$, $A_{2}$ be the event $y\geq\frac{7}{4}k$, and $A_{3}$ be $\frac{3}{2}k<y<\frac{7}{4}k$. Consider the potential function
$h(P_{t})=\begin{cases}0,&{}\text{if $v\in P_{t}$},\\ e\mu n^{k/2},&{}\text{if $A_{1}$},\\ e\mu n^{k/2}+1,&{}\text{if $A_{2}\cup A_{3}$}.\end{cases}$
For $A_{1}$, we pessimistically consider the decrease of $h$ by generating $v$ and the increase of $h$ by kicking out a corresponding solution of $y$. Hence, we have
$E[h$ ${}(P_{t})-h(P_{t+1})\mid P_{t},A_{1}]\geq\frac{1}{\mu}\frac{1}{n^{y-k}}\mathopen{}\mathclose{{}\left(1-\frac{1}{n}}\right)^{n-(y-k)}e\mu n^{k/2}-\frac{1}{2}\cdot 1\geq 1-\frac{1}{2}=\frac{1}{2}.$
For $A_{2}$, $h$ will not increase, and there are $y-k\in[\frac{3}{4}k,k]$ ones in the block of a corresponding solution of $y$. We pessimistically consider the decrease of $h$ by flipping $y-k-\frac{1}{2}k$ ones and keeping other bits unchanged of the corresponding solution. Hence, we have
$E[h$ ${}(P_{t})-h(P_{t+1})\mid P_{t},A_{2}]\geq\frac{1}{\mu}\frac{\binom{y-k}{y-k-k/2}}{n^{y-k-k/2}}\mathopen{}\mathclose{{}\left(1-\frac{1}{n}}\right)^{n-(y-k-k/2)}\frac{1}{2}$
[…]
where the last inequality uses $\frac{3}{4}k\leq y-k\leq k$. For $A_{3}$, $h$ will not increase, and there are $y-k\in(\frac{1}{2},\frac{3}{4}k)$ ones in the block of a corresponding solution of $y$. We pessimistically consider the decrease of $h$ by flipping $\frac{1}{4}k$ ones and keeping other bits unchanged of the corresponding solution. Hence, we have
$E[h$ ${}(P_{t})-h(P_{t+1})\mid P_{t},A_{3}]\geq\frac{1}{\mu}\frac{\binom{y-k}{k/4}}{n^{k/4}}\mathopen{}\mathclose{{}\left(1-\frac{1}{n}}\right)^{n-k/4}\frac{1}{2}\geq\frac{1}{2e\mu}\frac{\binom{k/2}{k/4}}{n^{k/4}}.$
Therefore, we have
$E[h$ ${}(P_{t+1})-h(P_{t})\mid P_{t},v\notin P_{t}]=\sum_{i=1}^{3}\Pr[A_{i}]E[h(P_{t+1})-h(P_{t})\mid P_{t},A_{i}]$
[…]
where the third inequality uses $\binom{k/2}{k/4}\leq\binom{\frac{3}{4}k}{k/2}$ and $n^{k/2}\geq n^{k/4}$, and the last inequality uses $k\geq 1$. Since the maximum $h$ is $e\mu n^{k/2}+1$, we know that the Pareto front point $v$ will be reached in at most
$(e\mu n^{k/2}+1)\frac{{4e\mu n^{k/2}\sqrt{k}}}{4^{k/4}}=O\mathopen{}\mathclose{{}\left(\frac{\mu k^{1/2}}{2^{k/2}}\mu n^{k}}\right)$
in expectation, and will survive in all future populations. Consider the process of covering all non-inner Pareto front points in the order of covering the $1$-st level, …, $(m/2)$-th level. Since the 00-th level points are already covered from our assumption, and since there are at most $M$ non-inner Pareto front points, we know that the remaining levels are covered in at most $O(\frac{\mu k^{1/2}}{2^{k/2}}M\mu n^{k})$ iterations in expectation. With the same argument in the proof of Lemma 7, we know that the expected number of iterations to cover the full Pareto front is at most $eM\mu n^{k}$. Hence, we proved this lemma. ∎
Combining Lemmas 12 and 13, we have the runtime of the SMS-EMOA with stochastic population update in the following theorem.
Theorem 14.
Let $k\leq n^{\prime}/2$. Consider using SMS-EMOA with $\mu\geq 2(M+1)$ and stochastic population update to optimize the $m$OJZJ problem. Then after at most $\frac{1}{2}em^{k+1}k^{k}+en\mu M+O(\min\{\frac{\mu k^{1/2}}{2^{k/2}},1\}M\mu n^{k})$ iterations in expectation, the full Pareto front is covered.
Theorem 14 shows that the best runtime for the SMS-EMOA with stochastic population update is $O(\min\{\frac{Mk^{1/2}}{2^{k/2}},1\}M^{2}n^{k})$ when we choose $\mu=\Theta(M)$. Recall that the runtime of the original SMS-EMOA is $O(M^{2}n^{k})$ for $\mu=\Theta(M)$ from Theorem 8. We know the asymptotic speed-up is $\Theta(\max\{\frac{2^{k/2}}{Mk^{1/2}},1\})$ with $M=(2n/m-2k+3)^{m/2}$. Hence, when the objective number is not small, like at least the same size of $k$, then the speed-up is only $\Theta(1)$. We note that it is not contradictory to the nice speed-up obtained for two objectives in [BZLQ23] but adds a limitation situation for the stochastic population update.
6 Heavy-Tailed Mutation Helps
In the previous section, we saw that the advantages of a stochastic population update do not well generalize from bi-objective optimization to many-objective optimization. We now regard another design choice that so far was only analyzed in bi-objective optimization, namely a heavy-tailed mutation operator. We shall prove that the $k^{\Omega(k)}$ factor speed-up observed in bi-objective optimization extends to many objectives.
6.1 Heavy-Tailed Mutation
Different from the standard bit-wise mutation, the heavy-tailed mutation flips each bit of the solution with probability $\alpha/n$ independently, where $\alpha$ follows a power-law distribution with parameter $\beta$ defined as follows.
Definition 15.
Let $n\in\mathbb{N}$ and $\beta>1$. We say $\alpha$ follows the power-law distribution if for all $i\in[1..n/2]$, $\Pr[\alpha=i]=\mathopen{}\mathclose{{}\left(\sum_{j=1}^{n/2}k^{-\beta}}\right)^{-1}i^{-\beta}$.
This power-law distribution has a good feature that a point with Hamming distance of $j$ of a given bit-string will be generated with a probability of $\Omega(j^{-\beta})$. It will facilitate the jump with a larger size and facilitate escaping from the local optimum. The heavy-tailed mutation has resulted in an asymptotic performance speed-up by a factor of $k^{\Omega(k)}$ of the $(1+1)$ EA on the single-objective Jump function with gap size $k$ [DLMN17]. For the bi-objective OJZJ, the speed-up of $k^{\Omega(k)}$ is also witnessed when it is applied to the GSEMO [DZ21] and to the NSGA-II [DQ23a]. Several other positive theoretical results exist for this heavy-tailed mutation, or more generally, other heavy-tailed parameter choises [FQW18, WQT18, QGWF21, COY21, DELQ22, ABD22, DGI22, DR23]. Prior to this work, no theoretical analysis of the heavy-tailed mutation operator in many-objective optimization existed.
6.2 Runtime
We now analyze the runtime of the SMS-EMOA using the heavy-tailed mutation operator instead of bit-wise mutation, the standard choice for this algorithm.
Since the heavy-tailed mutation does not change the survival selection of the original SMS-EMOA, we easily have the following corollary about the survival of the individuals with rank one from Lemma 4.
Corollary 16.
Consider using SMS-EMOA with heavy-tailed mutation and with $\mu\geq M$ to optimize the $m$OJZJ problem. For any $x\in F_{1}$ for the current generation $t$, we have that there is $y$ with $f(y)=f(x)$ in $P_{t+1}$. Specifically, any Pareto front point will be maintained in all future populations once it is reached.
With similar proof ideas in Lemma 5 to 7 but noticing the probability of generating a solution with a specific Hamming distance, we obtain the following lemmas for reaching the first inner Pareto optimum (Lemma 17), covering all inner Pareto front points (Lemma 18), and covering the remaining Pareto front points (Lemma 19).
Lemma 17.
Let $k\leq n^{\prime}/2$. Consider using SMS-EMOA with heavy-tailed mutation and with $\mu\geq M$ to optimize the $m$OJZJ problem. Then after at most $C_{n/2}^{\beta}\binom{n}{k}mk^{\beta}$ iterations in expectation, the population (and also the populations afterwards) contains at least one inner Pareto optimum.
Proof.
We use similar arguments of Lemma 5 except following probability calculations. The probability to generate a $y$ with $|y_{B_{i}}|_{1}\in[k..n^{\prime}-k]$ from $x$ with $A_{i}<k$ or with $A_{i}>n^{\prime}-k$ is at least $\binom{n}{k}^{-1}P_{k}^{\beta}\geq\binom{n}{k}^{-1}\mathopen{}\mathclose{{}\left(C_{n/2}^{\beta}}\right)^{-1}k^{-\beta}$. Hence such $y$ is generated in expected $\binom{n}{k}C_{n/2}^{\beta}k^{\beta}$ iterations. Hence the expected iterations to reach an inner Pareto optimum is at most $\frac{m}{2}\binom{n}{k}C_{n/2}^{\beta}k^{\beta}=\binom{n}{k}C_{n/2}^{\beta}mk^{\beta}$ iterations. ∎
Lemma 18.
Let $k\leq n^{\prime}/2$. Consider using SMS-EMOA with heavy-tailed mutation and with $\mu\geq M$ to optimize the $m$OJZJ problem. Assume that the current population contains at least one inner Pareto optimum. Then after at most $\frac{e\beta}{\beta-1}n\mu M$ iterations in expectation, all inner Pareto front points are covered.
Proof.
We use similar arguments of Lemma 6 except following probability calculation. The probability to generate an offspring with its function value of an uncovered inner Pareto front point from a neighbor covered by the population is at least $\frac{1}{n\mu}P_{1}^{\beta}=\frac{\beta-1}{e\beta n\mu}$. ∎
Lemma 19.
Let $k\leq n^{\prime}/2$. Consider using SMS-EMOA with heavy-tailed mutation and with $\mu\geq M$ to optimize the $m$OJZJ problem. Assume that the current population covers all inner Pareto front points. Then after at most $\binom{n}{k}C_{n/2}^{\beta}\mu Mk^{\beta}$ iterations in expectation, the full Pareto front is covered.
Proof.
We use similar arguments of Lemma 7 except following probability calculation. The probability to generate any point in $j$-th level condition on that all $(j-1)$-th level points are covered is at least
$\frac{1}{\mu}\binom{n}{k}^{-1}P_{k}^{\beta}\geq\binom{n}{k}^{-1}\mathopen{}\mathclose{{}\left(C_{n/2}^{\beta}}\right)^{-1}\frac{1}{\mu}k^{-\beta}.\qed$
Considering the whole process for the full coverage of the Pareto front from Lemmas 17 to 19, we have the following runtime result for the SMS-EMOA with heavy-tailed mutation.
Theorem 20.
Let $k\leq n^{\prime}/2$. Consider using SMS-EMOA with $\mu\geq M$ to optimize the $m$OJZJ problem. Then after at most $O(M\mu k^{\beta-0.5-k}n^{k})$ iterations or function evaluations in expectation, the full Pareto front is covered.
Theorem 20 shows the best asymptotic upper bound of the runtime of the SMS-EMOA with heavy-tailed mutation is $O(M^{2}n^{k}k^{\beta-0.5-k})$. Comparing with the runtime of $O(M^{2}n^{k})$ for $\mu=\Theta(M)$ for the original SMS-EMOA from Theorem 8, we see a speed-up of $k^{k+0.5-\beta}$, which is also the speed-up reached for the single objective and two objectives.
7 SMS-EMOA Also Performs Well for Bi-objective Optimization
The above sections discussed the runtime of the SMS-EMOA for many objectives, and showed that it performs well for many objectives, which is different from the original NSGA-II. Since the only theory paper [BZLQ23] merely considers its performance for the bi-objective OJZJ, in this section, we will analyze its runtime for other well-known bi-objective benchmarks to have a more comprehensive theoretical understanding.
7.1 OneMinMax and LOTZ
OneMinMax [GL10] (also the CountingOnesCountingZeroes [${}^{+}$02]) and LOTZ [LTZ04] (also the weighted version WLPTNO [QYZ13]) are the bi-objective counterpart of the well-analyzed single-objective OneMax and LeadingOnes respectively. For OneMinMax, two objectives count the number of ones and zeros in a given bit-string respectively. For LOTZ, the first objective counts the continuous ones starting from the first bit position, and the second one counts the continuous zeros starting from the last bit position. Both benchmarks with a simple and clear structure facilitate the fundamental theoretical understanding of the MOEAs. We note that there are other benchmarks [QTZ16, LZZZ16, DOSS23b], but here we focus on the two kinds of most popular benchmarks for analysis.
Here are formal definitions of the OneMinMax and LOTZ.
Definition 21 ([GL10]).
Let $n\in\mathbb{N}$ be the problem size. The OneMinMax function :$:\{0,1\}^{n}\rightarrow\mathbb{R}$ is defined by
$\textsc{OneMinMax}(x)=\mathopen{}\mathclose{{}\left(\sum_{i=1}^{n}(1-x_{i}),\sum_{i=1}^{n}x_{i}}\right)$
for any $x=(x_{1},\dots,x_{n})\in\{0,1\}^{n}$.
Definition 22 ([LTZ04]).
Let $n\in\mathbb{N}$ be the problem size. The LOTZ function :$:\{0,1\}^{n}\rightarrow\mathbb{R}$ is defined by
$\textsc{LOTZ}(x)=\mathopen{}\mathclose{{}\left(\sum_{i=1}^{n}\prod_{j=1}^{i}x_{j},\sum_{i=n}^{1}\prod_{j=n}^{i}x_{j}}\right)$
for any $x=(x_{1},\dots,x_{n})\in\{0,1\}^{n}$.
7.2 Runtime
We now establish the runtime of the SMS-EMOA on the OneMinMax and LOTZ. Note that the Pareto front sizes for OneMinMax and LOTZ are $n+1$. Hence, with the similar proof of Lemma 4, we know that for $\mu\geq n+1$ and for a specific function value $u$ reached by $F_{1}$, one individual in the non-empty $\cup_{i\geq 2}F_{i}$ or one individual $x$ in $F_{1}$ with $\Delta(x,F_{1})=0$ will be removed and $u$ will remain covered in the next generation. We have the following formal description.
Lemma 23.
Consider using SMS-EMOA with $\mu\geq n+1$ to optimize the LOTZ or OneMinMax problem. For any $x\in F_{1}$ for the current generation $t$, we have that there is $y$ with $f(y)=f(x)$ in $P_{t+1}$. Specifically, any Pareto front point will be maintained in all future populations once it is reached.
Theorem 24 shows the runtime of SMS-EMOA on OneMinMax. The proof considers the probability to reach an uncovered Pareto front point from a specific covered one. From Lemma 23 that any covered Pareto front point will be kept, we then turn such probability to the runtime argument.
Theorem 24.
Consider using SMS-EMOA with $\mu\geq n+1$ to optimize the OneMinMax problem with problem size $n$. Then after at most $2e\mu n(\ln n+1)$ iterations in expectation, the full Pareto front is covered.
Proof.
Since each search point in $\{0,1\}^{n}$ is Pareto optimal, from Lemma 23, we know that any Pareto front point covered by the population $P_{t}$ will be maintained in all future generations. Assume that $(i,n-i)$ for some $i\in[0..n]$ is covered in $P_{t}$. Then the probability to reach a solution with function value $(i+1,n-i-1)$ is at least
$\frac{1}{\mu}\frac{n-i}{n}\mathopen{}\mathclose{{}\left(1-\frac{1}{n}}\right)^{n-1}\geq\frac{n-i}{e\mu n},$
and the probability to reach a solution with function value $(i-1,n-i+1)$ is at least
$\frac{1}{\mu}\frac{i}{n}\mathopen{}\mathclose{{}\left(1-\frac{1}{n}}\right)^{n-1}\geq\frac{i}{e\mu n}.$
Thus the expected numbers of iterations to cover $(i+1,n-i-1)$ and $(i-1,n-i+1)$ are at most $\frac{e\mu n}{n-i}$ and at most $\frac{e\mu n}{i}$ respectively. Hence, the expected number of iterations to cover the full Pareto front is
$\Bigg{(}\sum_{j=i}^{1}\frac{e\mu n}{n-j}\Bigg{)}+\mathopen{}\mathclose{{}\left(\sum_{j=i}^{n-1}\frac{e\mu n}{j}}\right)\leq 2\sum_{j=1}^{n-1}\frac{e\mu n}{j}\leq 2e\mu n(\ln n+1).\qed$
We now establish the runtime of SMS-EMOA on LOTZ, see Theorem 25. In the proof, the process is divided into two stages. The first stage is to cover a Pareto front point for the first time and the second stage is to cover the full Pareto front starting from one covered Pareto front point. For the first stage, we pessimistically consider the time to reach $1^{n}$ and for the second stage, we consider the time to reach $01^{i}0^{n-i},i=n-1,\dots,$ 0 one by one.
Theorem 25.
Consider using SMS-EMOA with $\mu\geq n+1$ to optimize the LOTZ problem with problem size $n$. Then after at most $2e\mu n^{2}$ iterations in expectation, the full Pareto front is covered.
Proof.
We first consider the time to reach $1^{n}$ in the first stage. Let $S=\arg\max_{x\in P_{t}}f_{1}(x)$ and $S^{\prime}=\arg\max_{x\in R_{t}}f_{1}(x)$. Then there is at least one individual $x\in S^{\prime}$ such that $x\in F_{1}$ (If not, there is $y\notin S^{\prime}$, such that $x\prec y$ for all $x\in S^{\prime}$. That is, $f_{1}(y)\geq f_{1}(x),f_{2}(y)\geq f_{2}(x),$ and at least one of the inequality is strict. Since $y\notin S$, we know that $f_{1}(y)<f_{1}(x)$, which contradicts $f_{1}(y)\geq f_{1}(x)$). From Lemma 23, we know that $f(x)$ (also $f_{1}(x)$) will be covered by $P_{t+1}$. Hence, to reach $1^{n}$, we only need to reach such $S^{\prime}$ with $f_{1}(S^{\prime})=n$. We calculate
$\Pr[f_{1}(S^{\prime})>f_{1}(S)\mid f_{1}(S)]\geq\frac{1}{\mu}\mathopen{}\mathclose{{}\left(1-\frac{1}{n}}\right)^{n-1}\frac{1}{n}\geq\frac{1}{e\mu n}$
where we consider choosing one individual of $S$ and flipping the critical bit value from 00 to $1$. Since $f_{1}(S)$ will not decrease, we know that the expected iterations to witness such $S^{\prime}$ is at most $e\mu n$. Since there are at most $n$ such $S^{\prime}$ with $f_{1}(S^{\prime})>f_{1}(S)$ to reach $1^{n}$, we know that $1^{n}$ will be reached in at most $e\mu n^{2}$ in expectation.
Once $1^{n}$ is reached, we consider the time to reach $1^{n-1}0,\dots,10^{n-1},0^{n}$ one by one. Assume that $1^{i}0^{n-i}$ with $i\in[1..n]$ is reached. Then to reach $1^{i-1}0^{n-i+1}$, we only need to choose $1^{i}0^{n-i}$ from $P_{t}$, and flip the $i$-th bit from $1$ to 00 with other bits unchanged, which happens with probability
$\frac{1}{\mu}\mathopen{}\mathclose{{}\left(1-\frac{1}{n}}\right)^{n-1}\frac{1}{n}\geq\frac{1}{e\mu n}.$
From Lemma 23, we know that any Pareto front point will be maintained in all future generations once reached. Hence, $1^{i}0^{n-i}\in P_{\tau}$ for all $\tau\geq t$. Thus, the expected number of iterations to reach $1^{i-1}0^{n-i+1}$ is at most $e\mu n$. Since we have $n$ other Pareto optima to reach, we know that the expected number of iterations to cover all remaining Pareto front points is at most $e\mu n^{2}$.
In summary, the full Pareto front will be covered in at most $2e\mu n^{2}$ iterations in expectation. ∎
Note that the runtime of the GSEMO and NSGA-II are $O(n^{2}\ln n)$ and $O(n^{3})$ for the OneMinMax and LOTZ respectively. Hence, Theorems 24 and 25 shows the SMS-EMOA achieves the same asymptotic runtime as GSEMO and NSGA-II.
8 Conclusion
Motivated by the observed difficulty of the NSGA-II for many objectives, this paper resorted to the SMS-EMOA, a variant of the steady-state NSGA-II, and proved that, different from the NSGA-II, it efficiently solves the $m$OJZJ problem. Noting that the SMS-EMOA also employs non-dominated sorting, but replaces the crowding distance with the hypervolume, this result together with the one of Wietheger and Doerr [WD23] supports the conclusion that non-dominated sorting is a good building block for MOEAs, but the crowding distance has deficiencies for more than two objectives.
We also showed that the stochastic population update proposed in [BZLQ23] for the bi-objective SMS-EMOA becomes less effective for many objectives. All these results in a very rigorous manner support the general knowledge that multi-objective optimization becomes increasingly harder with growing numbers of objectives.
On the positive side, we showed that the advantages of heavy-tailed mutation, previously observed in single- and bi-objective optimization, remain when increasing the number of objectives.
Given that our first result shows advantages of the SMS-EMOA and there is only one previous work performing rigorous runtime analyses for this algorithm, we extended our knowledge in this direction by proving competitive runtime guarantees for this algorithm on the two most prominent bi-objective benchmarks.
Acknowledgments
This work was supported by National Natural Science Foundation of China (Grant No. 62306086), Science, Technology and Innovation Commission of Shenzhen Municipality (Grant No. GXWD20220818191018001), Guangdong Basic and Applied Basic Research Foundation (Grant No. 2019A1515110177). This research benefited from the support of the FMJH Program Gaspard Monge for optimization and operations research and their interactions with data science.

### v2

@@
 Finally, we conduct the first runtime analyses of the SMS-EMOA on the bi-objective OneMinMax and LOTZ benchmarks and show that it has a performance comparable to the GSEMO and the NSGA-II.
-1 Introduction
+Introduction
 The NSGA-II is the most widely-applied multiobjective evolutionary algorithm (MOEA).
 Non-dominated sorting and crowding distance are its two major features differentiating it from basic MOEAs such as the GSEMO or the $(\mu+1)$ SIBEA.
-Zheng, Liu, and Doerr [ZLD22] conducted the first runtime analysis of the NSGA-II (see [ZD23a] for the journal version).
-This work quickly inspired many interesting follow-up results in bi-objective optimization [ZD22, BQ22, DQ23a, DQ23b, DQ23c, DOSS23a, DOSS23b, ${}^{+}$23].
-In contrast to these positive results for two objectives, Zheng and Doerr [ZD23b] proved that for $m\geq 3$ objectives the NSGA-II needs at least exponential time (in expectation and with high probability) to cover the full Pareto front of the $m$-objective OneMinMax benchmark, a simple many-objective version of the basic OneMax problem where all search points are Pareto optimal.
+Zheng, Liu, and Doerr (2022) conducted the first runtime analysis of the NSGA-II (see (Zheng and Doerr 2023a) for the journal version).
+This work quickly inspired many interesting follow-up results in bi-objective optimization (Zheng and Doerr 2022; Bian and Qian 2022; Doerr and Qu 2023a, b, c; Dang et al.
+2023a, b; Cerf et al.
+2023).
+In contrast to these positive results for two objectives, Zheng and Doerr (2023b) proved that for $m\geq 3$ objectives the NSGA-II needs at least exponential time (in expectation and with high probability) to cover the full Pareto front of the $m$-objective OneMinMax benchmark, a simple many-objective version of the basic OneMax problem where all search points are Pareto optimal.
 They claimed that the main reason for this low efficiency is the independent computation of the crowding distance in each objective.
-A very recent work showed that the NSGA-III, a successor algorithm of the NSGA-II aimed at better coping with many objectives, can efficiently solve the $3$-objective OneMinMax problem [WD23].
+A very recent work showed that the NSGA-III, a successor algorithm of the NSGA-II aimed at better coping with many objectives, can efficiently solve the $3$-objective OneMinMax problem (Wietheger and Doerr 2023).
 Since apparently practitioners much prefer the NSGA-II (more than 17,000 citations on Google scholar only in the last five years) over the NSGA-III (5,035 citations since its publication in 2013), it remains an interesting question whether there are variants of the NSGA-II which better cope with many objectives.
-With the SMS-EMOA, an interesting variant of the NSGA-II was proposed by eume, Naujoks, and Emmerich [BNE07].
+With the SMS-EMOA, an interesting variant of the NSGA-II was proposed by Beume, Naujoks, and Emmerich (2007).
 This algorithm is a steady-state variant of the NSGA-II (that is, in each iteration only a single offspring is generated and possibly integrated into the population) that further replaces the crowding distance as secondary selection criterion with the classic hypervolume contribution.
-Many empirical works (see the almost 2,000 papers citing [BNE07]) confirmed the good performance of the SMS-EMOA for many-objective optimization.
-The only mathematical runtime analysis of the SMS-EMOA was conducted only very recently by Bian et al.
-[BZLQ23].
-They proved that its runtime for the bi-objective OJZJ problem is $O(n^{k+1})$ and proposed a stochastic population update mechanism which was proven to achieve an expected runtime of $O(n^{k+1}\min\{1,n/2^{k/4}\})$.
+Many empirical works (see the almost 2,000 papers citing (Beume, Naujoks, and Emmerich 2007)) confirmed the good performance of the SMS-EMOA for many-objective optimization.
+The first mathematical runtime analysis of the SMS-EMOA was conducted very recently by Bian et al.
+(2023), who proved that its expected runtime on the bi-objective OJZJ problem is $O(n^{k+1})$.
+They also proposed a stochastic population update mechanism and proved that it has the often superior runtime of $O(n^{k+1}\min\{1,n/2^{k/4}\})$.
+Zheng et al.
+(2024) proved that the SMS-EMOA has an expected runtime of $O(n^{4})$ on the bi-objective DLTB problem.
 Our Contributions: This paper conducts the first mathematical runtime analysis of the SMS-EMOA for more than two objectives.
-We first define the $m$OJZJ benchmark, an $m$-objective counterpart of the bi-objective OJZJ problem [DZ21], which is the problem analyzed in the only previous runtime analysis for the SMS-EMOA [BZLQ23].
+We first define the $m$OJZJ benchmark, an $m$-objective counterpart of the bi-objective OJZJ problem (Zheng and Doerr 2023d), which is the problem analyzed in the first runtime analysis for the SMS-EMOA (Bian et al.
+2023).
 We note that the $m$OJZJ problem is the first multimodal many-objective benchmark proposed for a theoretical analysis, to the best of our knowledge.
@@
 We recall that the original NSGA-II needs at least exponential time to optimize the $m$OneMinMax problem, which is a special case of $m$OJZJ with gap size $k=1$.
-Since the SMS-EMOA employs non-dominated sorting, but replaces the crowding distance in the original NSGA-II by the hypervolume contribution, our result in a similar fashion as the analysis of the NSGA-III in [WD23] suggests that the general approach of the NSGA-II is suitable for many-objective optimization and that it is only the crowding distance as tie-breaker which is not appropriate for more than two objectives.
-We then analyze whether the better performance of the SMS-EMOA on the bi-objective OJZJ problem achieved via a new stochastic population update [BZLQ23] extends to the $m$-objective $m$OJZJ problem.
+Since the SMS-EMOA employs non-dominated sorting, but replaces the crowding distance in the original NSGA-II by the hypervolume contribution, our result in a similar fashion as the analysis of the NSGA-III in (Wietheger and Doerr 2023) suggests that the general approach of the NSGA-II is suitable for many-objective optimization and that it is only the crowding distance as tie-breaker which is not appropriate for more than two objectives.
+We then analyze whether the better performance of the SMS-EMOA on the bi-objective OJZJ problem achieved via a new stochastic population update (Bian et al.
+2023) extends to the $m$-objective $m$OJZJ problem.
 Unfortunately, we shall observe that only a speed-up of order $1/\Theta(\min\{Mk^{1/2}/2^{k/2},1\})$ is obtained, which is $\Theta(1)$ when $m$ is large, e.g., $m>k$.
@@
 We analyze the SMS-EMOA with heavy-tailed mutation on the $m$OJZJ benchmark and prove that a speed-up of order $k^{0.5+k-\beta}$ is achieved.
-This is the same speed-up as observed for single-objective and bi-objective Jump problems [DLMN17, DZ21, ZD23c].
+This is the same speed-up as observed for single-objective and bi-objective Jump problems (Doerr et al.
+2017; Zheng and Doerr 2023d).
 We note that this is the first theoretical work to support the usefulness of heavy-tailed mutation in many-objective optimization.
-Finally, since so far the performance of the SMS-EMOA was only analyzed on the bi-objective OJZJ problem [BZLQ23], we conduct mathematical runtime analyses of the SMS-EMOA also on the two most prominent bi-objective benchmarks OneMinMax and LOTZ.
+Finally, since so far the performance of the SMS-EMOA was only analyzed on the bi-objective OJZJ problem (Bian et al.
+2023), we conduct mathematical runtime analyses of the SMS-EMOA also on the two most prominent bi-objective benchmarks OneMinMax and LOTZ.
 We prove that the SMS-EMOA finds the Pareto fronts of these benchmarks in an expected number of iterations of at most $2e(n+1)n(\ln n+1)$ for OneMinMax and at most $2en^{2}(n+1)$ for LOTZ.
 These are the same asymptotic runtimes (in terms of fitness evaluations) as known for the GSEMO and the NSGA-II.
-This paper is organized as follows.
-Section 2 briefly introduces some basic concepts for the many-objective optimization and the mathematical notations that will be used in this paper.
-The many-objective OJZJ benchmark is proposed in Section 3.
-Sections 4 to 6 analyze the runtime of the original SMS-EMOA, the SMS-EMOA with stochastic population update, and the one with heavy-tailed mutation on this many-objective benchmark.
-The runtime for two well-known bi-objective functions is conducted in Section 7.
-Section 8 concludes this work.
-2 Preliminaries
+Preliminaries
 In many-objective optimization, one tries to find good solutions for a problem containing several, but at least three objectives.
@@
 As common in the mathematical runtime analysis of MOEAs, we call the runtime of an algorithm the number of function evaluations until its population $P$ covers the full Pareto front, that is, $f(P)$ contains the Pareto front.
-We refer to [NW10, AD11, Jan13, ZYQ19, DN20] for general introductions to the mathematical runtime analysis of evolutionary algorithms, and to [Bro11] for a discussion of runtime analyses of MOEAs.
+We refer to (Neumann and Witt 2010; Auger and Doerr 2011; Jansen 2013; Zhou, Yu, and Qian 2019; Doerr and Neumann 2020) for general introductions to the mathematical runtime analysis of evolutionary algorithms, and to (Brockhoff 2011) for a discussion of runtime analyses of MOEAs.
 Mathematical Notations.
@@
 For a point $x\in\{0,1\}^{n}$, let $x_{[a..b]}:=(x_{a},x_{a+1},\dots,x_{b})$ and $|x|_{1}$ be the number of ones in $x$.
-3 The $m$-Objective Jump Function
-As mentioned earlier, the NSGA-II was proven to have enormous difficulties in optimizing many-objective problems [ZD23b].
+The $m$-Objective Jump Function
+As mentioned earlier, the NSGA-II was proven to have enormous difficulties in optimizing many-objective problems (Zheng and Doerr 2023b).
 In that paper, the $m$-objective counterpart $m$OneMinMax of the bi-objective OneMinMax benchmark was proposed and analyzed.
-Since the only runtime work of the SMS-EMOA so far [BZLQ23] analyzed the bi-objective jump problem OJZJ, whose special case with gap size $k=1$ is (essentially) the OneMinMax problem, we shall propose and work with an $m$-objective version of OJZJ as well.
+Since the first runtime work of the SMS-EMOA so far (Bian et al.
+2023) analyzed the bi-objective jump problem OJZJ, whose special case with gap size $k=1$ is (essentially) the OneMinMax problem, we shall propose and work with an $m$-objective version of OJZJ as well.
 Again, its special case $k=1$ will be (essentially) equal to the $m$OneMinMax problem.
-With this, our results are comparable both the ones in [ZD23b] and [BZLQ23].
-3.1 $m$OJZJ
+With this, our results are comparable both the ones in (Zheng and Doerr 2023b) and (Bian et al.
+2023).
+$m$OJZJ
 We first recall the definition of the $m$OneMinMax problem.
@@
 On each of these, a bi-objective OneMinMax problem is defined.
-We note that this general block construction goes back to the seminal paper of [LTZ04].
-Definition 1 ([ZD23b]).
+We note that this general block construction goes back to the seminal paper of (Laumanns, Thiele, and Zitzler 2004).
+Definition 1 (Zheng and Doerr (2023b)).
 Let $m$ be the number of objectives and be even, and the problem size $n$ be a multiple of $m/2$.
@@
 We note that the function $\textsc{Jump}_{n^{\prime},k}$ used in the definition above is the famous Jump benchmark.
-It was first defined by Droste, Jansen, and Wegener [DJW02] and has quickly become the most employed multimodal benchmark in the theory of randomized search heuristics, leading to many fundamental results on how these algorithms cope with local optima [JW02, ${}^{+}$18, Doe21, BBD21, HS22, RW22, Wit23, DEJK24].
+It was first defined by Droste, Jansen, and Wegener (2002) and has quickly become the most employed multimodal benchmark in the theory of randomized search heuristics, leading to many fundamental results on how these algorithms cope with local optima (Jansen and Wegener 2002; Dang et al.
+2018; Doerr 2021; Benbaki, Benomar, and Doerr 2021; Hevia Fajardo and Sudholt 2022; Rajabi and Witt 2022; Witt 2023; Doerr et al.
+2024).
 […]
-3.2 Characteristics
+Characteristics
 We now give more details on this $m$OJZJ function.
@@
 Figure 1 plots the objective values of $f_{2i-1}$ and $f_{2i}$ relative to the number of ones in this block.
-Obviously, $m$OJZJ is multimodal with respect to the definition of multimodality of multiobjective problems in [DZ21].
+Obviously, $m$OJZJ is multimodal with respect to the definition of multimodality of multiobjective problems in (Zheng and Doerr 2023d).
 Figure 1: The objective values of $f_{2i-1}$ and $f_{2i}$ in $m$OJZJ w.r.t. $|x_{B_{i}}|_{1}$, the number of ones in the block $B_{i}$.
@@
 However, we have the following result showing that the maximum number of mutually non-dominated points is at most the size of the Pareto front.
+Due to the limited space, all mathematical proofs could only be sketched or had to be omitted completely.
+They can be found in the preprint (Zheng and Doerr 2023c).
 Lemma 3.
-Let $k\leq n^{\prime}/2$ and $P$ be a set of solutions such that any two are not weakly dominating (w.r.t. $m$OJZJ). Then $|P|\leq M$.
-Proof.
-Note that all objective values of any non-Pareto optimal solution are less than $2k$.
-Since each Pareto optimum has its all objective values of at least $2k$, we know that any non-Pareto optimal solution is dominated by all Pareto optima excluding the ones with some block all zeros or all ones (called extreme-block Pareto optima).
-Hence, $P$ excluding extreme-block Pareto optima, denoted as $P^{\prime}$, cannot include both Pareto optima and non-Pareto optimal solutions.
-If $P$ includes only Pareto optima, then we know that $|P|\leq M$.
-We now consider that $P^{\prime}$ includes only non-Pareto optimal solutions, that is, for any $x\in P^{\prime}$, there exists a block $B_{i}:=[(i-1)n^{\prime}+1..in^{\prime}],i\in[1..m/2]$ such that $|x_{B_{i}}|_{1}\in[1..k-1]\cup[n^{\prime}-k+1..n^{\prime}-1]$.
-Note that for any block $B_{i}$, any two solutions $x,y$ are comparable w.r.t.
-$(f_{2i-1},f_{2i})$ if $(|x_{B_{i}}|_{1}\in[1..k-1],|y_{B_{i}}|_{1}\in[1..k-1])$ or $(|x_{B_{i}}|_{1}\in[n^{\prime}-k+1..n^{\prime}-1],|y_{B_{i}}|_{1}\in[n^{\prime}-k+1..n^{\prime}-1])$.
-Hence, with the bit values on other blocks fixed, there are at most two solutions in $P^{\prime}$ that are not weakly dominated.
-We also note that $(f_{2i-1}(z),f_{2i}(z))>(f_{2i-1}(x),f_{2i}(x))$ holds if $(z_{B_{i}}=1^{n^{\prime}},|x_{B_{i}}|_{1}\in[1..k-1])$ or $(z_{B_{i}}=0^{n^{\prime}},|x_{B_{i}}|_{1}\in[n^{\prime}-k+1..n^{\prime}-1])$.
-Hence, with the bit values on other blocks fixed, there are at most two solutions in $P$ that are not weakly dominated.
-Then $m/2$ blocks results in at most $2^{m/2}$ non-weakly-dominated solutions, that is, $|P|\leq 2^{m/2}<3^{m/2}\leq M$.
-∎
+Let $k\leq n^{\prime}/2$ and $P$ be a set of solutions such that any two are not weakly dominated (w.r.t. $m$OJZJ). Then $|P|\leq M$.
 For ease of reading of the proof sketches in the following sections, we call a Pareto optimum $x$ an inner Pareto optimum if for all blocks $B_{i}=[(i-1)n^{\prime}+1..in^{\prime}],i\in[1..m/2]$, we have $x_{B_{i}}\notin\{0^{n^{\prime}},1^{n^{\prime}}\}$.
 We call $u=(u_{1},n^{\prime}+2k-u_{1},\dots,u_{m/2},n^{\prime}+2k-u_{m/2})$ and $v=(v_{1},n^{\prime}+2k-v_{1},\dots,v_{m/2},n^{\prime}+2k-v_{m/2})$ neighbors if there exists $i\in[1..m/2]$ such that $u_{i}-v_{i}=1$ and $u_{j}=v_{j}$ for all $j\neq i$.
-4 SMS-EMOA Can Optimize $m$OJZJ
-As discussed before, the original NSGA-II with non-dominated sorting and crowding distance needs at least exponential runtime to cover the full Pareto front of the $m$OneMinMax [ZD23b], which is a special case of $m$OJZJ.
-Motivated by the question whether there is a simple efficient variant of NSGA-II for many objectives, this section will analyze the runtime of the SMS-EMOA, which still uses the non-dominated sorting but replaces the crowding distance by hypervolume measure.
-4.1 Algorithm Description
+The SMS-EMOA Can Optimize $m$OJZJ
+As discussed before, the original NSGA-II with non-dominated sorting and crowding distance needs at least exponential runtime to cover the full Pareto front of $m$OneMinMax, $m\geq 3$, (Zheng and Doerr 2023b), which is a special case of $m$OJZJ.
+In this section, we analyze the runtime of the SMS-EMOA on this problem and show that it does not encounter such problems.
+Algorithm Description
 The SMS-EMOA is a steady-state variant (that is, the offspring size is smaller than the parent size) of the NSGA-II.
-It maintains a fixed population size $\mu$.
-For each iteration, only one offspring $x^{\prime}$ is generated from the parent population $P_{t}$.
-From the $\mu+1$ individuals in the combined population $R_{t}=P_{t}\cup\{x^{\prime}\}$, one individual will be removed.
-It first sorts the combined population into several fronts $F_{1},\dots,F_{i^{*}}$, where $F_{i}$ contains all mutually non-dominated individuals in $R_{t}\setminus(\bigcup_{j=1}^{i-1}F_{j})$.
-This is called non-dominated sorting.
-Different from the original NSGA-II that uses the crowding distance as the secondary criterion to remove the individual, SMS-EMOA removes the one with the smallest hypervolume contribution in the critical front, that is, $F_{i^{*}}$.
-The hypervolume of the set $S$ w.r.t.
-a reference point $r$ is defined as
-$\operatorname{HV}_{r}(S)=\mathcal{L}\mathopen{}\mathclose{{}\left(\,\bigcup_{u\in S}\{h\in\mathbb{R}^{m}\mid r\leq h\leq f(u)\}}\right),$
-where $\mathcal{L}$ is the Lebesgue measure. The hypervolume contribution of an individual $x\in F_{i^{*}}$ is calculated by
+Like the NSGA-II, the SMS-EMOA works with a population of fixed size $\mu$.
+However, in each iteration, only one offspring $x^{\prime}$ is generated from the parent population $P_{t}$.
+From the $\mu+1$ individuals in the combined parent and offspring population $R_{t}=P_{t}\cup\{x^{\prime}\}$, a single individual is removed.
+To this aim, the SMS-EMOA like the NSGA-II uses non-dominated sorting, that is, it partitions $R_{t}$ into fronts $F_{1},\dots,F_{i^{*}}$, where $F_{i}$ contains all non-dominated individuals in $R_{t}\setminus(\bigcup_{j=1}^{i-1}F_{j})$.
+Different from the original NSGA-II that now uses the crowding distance as the secondary criterion for the removal of individuals, the SMS-EMOA removes the individual with the smallest hypervolume contribution in the critical front (which here always is the last front $F_{i^{*}}$).
+The hypervolume of a set $S$ of individuals w.r.t. a reference point $r$ in the objective space is defined as
+$\operatorname{HV}_{r}(S)=\mathcal{L}\left(\,\bigcup_{u\in S}\{h\in\mathbb{R}^{m}\mid r\leq h\leq f(u)\}\right),$
+where $\mathcal{L}$ is the Lebesgue measure. The hypervolume contribution of an individual $x\in F_{i^{*}}$ is calculated via
 $\Delta_{r}(x,F_{i^{*}}):=\operatorname{HV}_{r}(F_{i^{*}})-\operatorname{HV}_{r}(F_{i^{*}}\setminus\{x\}).$
-Algorithm 1 shows the general procedure of the SMS-EMOA.
+Algorithm 1 gives the pseudocode of the SMS-EMOA.
 1:Initialize $P_{0}$ by generating $\mu$ solutions uniformly at random from $\{0,1\}^{n}$
@@
 4:   Generate $x^{\prime}$ by flipping each bit of $x$ independently with probability $1/n$
-5:   %Original Population Update
-6:   Use fast-non-dominated-sort() [DPAM02] to divide $R_{t}=P_{t}\cup\{x^{\prime}\}$ into $F_{1},\dots,F_{i^{*}}$
-7:   Calculate $\Delta_{r}(z,F_{i^{*}})$ for all $z\in F_{i^{*}}$ and find $D=\arg\min_{z\in F_{i^{*}}}{\Delta_{r}(z,F_{i^{*}})}$
-8:   Uniformly at random pick $z^{\prime}\in D$ and $P_{t+1}=R_{t}\setminus\{z^{\prime}\}$
-9:end for
+5:   Use fast-non-dominated-sort() (Deb et al. 2002) to partition $R_{t}=P_{t}\cup\{x^{\prime}\}$ into $F_{1},\dots,F_{i^{*}}$
+6:   Calculate $\Delta_{r}(z,F_{i^{*}})$ for all $z\in F_{i^{*}}$ and find $D=\arg\min_{z\in F_{i^{*}}}{\Delta_{r}(z,F_{i^{*}})}$
+7:   Uniformly at random pick $z^{\prime}\in D$ and set $P_{t+1}=R_{t}\setminus\{z^{\prime}\}$
+8:end for
 Algorithm 1 SMS-EMOA
-4.2 Runtime of SMS-EMOA
-We now establish the runtime of the SMS-EMOA to cover the full Pareto front of the $m$OJZJ.
-We first give the following lemma to show that for a large enough population size, the function value of any point with rank one in the combined population will survive to the next population.
-The key of the proof is that all $F_{1}$ points will survive if there is a point with a larger rank, and that at least two points have the same function value if all points are in $F_{1}$ and the population size is large enough.
+Runtime of SMS-EMOA
+We now analyze the runtime of the SMS-EMOA, that is, the time until its population covers the full Pareto front of $m$OJZJ.
+We start by proving the following result ensuring that certain individuals survive in the population or are replaced by at least as good individuals.
+We formulate this result in manner more general than what we need since we expect this general version to be useful in other analyses of the SMS-EMOA.
+The key to the proof is that (i) if there is a point with a rank larger than one, then all solutions with rank one (that is, in $F_{1}$) will survive, and (ii) if $F_{1}$ contains two or more solutions with same function value, then only such a solution can be removed from $F_{1}$.
+If the population size is large enough, by the pigeon-hole principle, one of these two cases comes true.
 Lemma 4.
-Consider using SMS-EMOA with $\mu\geq M$ to optimize the $m$OJZJ problem.
-For any $x\in F_{1}$ for the current generation $t$, we have that there is $y$ with $f(y)=f(x)$ in $P_{t+1}$.
-Specifically, any Pareto front point will be maintained in all future populations once it is reached.
-Proof.
-Let $x\in F_{1}$. If $\cup_{i\geq 2}F_{i}\neq\emptyset$, then the removal will happen in $\cup_{i\geq 2}F_{i}$, and $x\in P^{t+1}$.
-If all solutions are in $F_{1}$, from Lemma 3, we know that $|f(F_{1})|\leq M$.
-Hence, there exist $z_{1},z_{2}\in F_{1}$ such that $f(z_{1})=f(z_{2})$.
-Let $S_{x}=\{z\in R_{t}\mid f(x)=f(z)\}$.
-Then $|S_{x}|\geq 1$.
-If $|S_{x}|=1$, then $\Delta(x,F_{1})>0=\Delta(z_{1},F_{1})$, and thus $x\in P_{t+1}$.
-If $|S_{x}|>1$, since the survival selection only removes one solution, then there exists $y\in S_{x}$, such that $y\in P_{t+1}$.
-Since any Pareto optimum is always in $F_{1}$, the latter part of this lemma holds. ∎
-Recall that the inner Pareto optimum is a Pareto optimum with $x_{B_{i}}\notin\{0^{n^{\prime}},1^{n^{\prime}}\}$ for all blocks $B_{i}=[(i-1)n^{\prime}+1..in^{\prime}],i\in[1..m/2]$, as defined before.
-Since not all points are inner Pareto optimum, the following lemma establishes the runtime to reach one inner Pareto optimum for the first time.
-The key of the proof is that at most $k$ specific bit flipping of a point that is not an inner Pareto optimum will result in an inner Pareto optimum.
+Consider any $m$-objective optimization problem.
+Let $M\in\mathbb{N}$ be such that any set $S$ of pairwise non-dominated solutions having different objective values satisfies $|S|\leq M$.
+Consider solving this problem via the SMS-EMOA with population size $\mu\geq M$ and using a reference point $r$ such that $\operatorname{HV}_{r}(\{x\})>0$ for any individual $x\in\{0,1\}^{n}$.
+Then the following is true.
+If at some time $t$ the combined parent and offspring population $R_{t}$ of the SMS-EMOA contains some solution $x$ (and thus in particular if $x\in P_{t}$), then at any later time $s>t$ the parent population $P_{s}$ contains a solution $y$ such that $y\succeq x$.
+In particular, if $R_{t}$ contains a Pareto optimum $x$, then all future generations contain a solution $y$ with $f(y)=f(x)$.
+Recall that an inner Pareto optimum is a Pareto optimum with $x_{B_{i}}\notin\{0^{n^{\prime}},1^{n^{\prime}}\}$ for all blocks $B_{i}=[(i-1)n^{\prime}+1..in^{\prime}],i\in[1..m/2]$, as defined before.
+While it is very likely that at least one initial solution is an inner Pareto optimum, this does not happen with probability one, and hence in the following lemma we estimate the time until the population contains an inner Pareto optimum.
+Since this time usually is much smaller than the time to generate all remaining Pareto optima, we do not care that this is estimate could easily be improved.
+The key to the proof is to note that changing a block $x_{B_{i}}$ to a bit-string with between $k$ and $n^{\prime}-k$ ones is relatively easy.
+By Lemma 4, an individual with this $f_{i}$ value will remain in the population.
+Hence a total of at most $m$ such block changes (applied to the right individual) suffice to obtain an inner Pareto optimum.
 Lemma 5.
@@
 Consider using SMS-EMOA with $\mu\geq M$ to optimize the $m$OJZJ problem.
-Then after at most $\frac{1}{2}em^{k+1}k^{k}$ iterations in expectation, the population (and also the populations afterwards) contains at least one inner Pareto optimum.
-Proof.
-This lemma is already proved if the initial population has at least one inner Pareto optimum.
-Now we consider that the initial population contains no inner Pareto optimum.
-Consider any population $P$ without inner Pareto optimum.
-Then we know that there exists a block $B_{i}:=[(i-1)n^{\prime}+1..in^{\prime}],i\in[1..m/2]$ such that for all $x\in P$, $A_{i}:=|x_{B_{i}}|_{1}\in[0..k-1]\cup[n^{\prime}-k+1..n^{\prime}]$.
-We regard it a success if $y$ with $|y_{B_{i}}|_{1}\in[k..n^{\prime}-k]$ is generated.
-Then the probability to generate such $y$ from $x$ with $A_{i}<k$ is at least
-$\binom{n^{\prime}-A_{i}}{k-A_{i}}$ ${}\mathopen{}\mathclose{{}\left(1-\frac{1}{n}}\right)^{n-(k-A_{i})}\mathopen{}\mathclose{{}\left(\frac{1}{n}}\right)^{k-A_{i}}\geq\frac{1}{e}\binom{n^{\prime}-A_{i}}{k-A_{i}}\mathopen{}\mathclose{{}\left(\frac{1}{n}}\right)^{k-A_{i}}\geq\frac{1}{e}\mathopen{}\mathclose{{}\left(\frac{n^{\prime}-A_{i}}{n(k-A_{i})}}\right)^{k-A_{i}}$
-$\geq{}\frac{1}{e}\mathopen{}\mathclose{{}\left(\frac{n^{\prime}-k}{nk}}\right)^{k}\geq\frac{1}{e}\mathopen{}\mathclose{{}\left(\frac{n^{\prime}}{2nk}}\right)^{k}=\frac{1}{e}\mathopen{}\mathclose{{}\left(\frac{1}{mk}}\right)^{k},$
-where the last inequality uses $k\leq n^{\prime}/2$ and the last equality uses $n^{\prime}=2m/n$. We also know that the probability to generate such $y$ from $x$ with $A_{i}>n^{\prime}-k$ is at least
-$\binom{A_{i}}{A_{i}-(n^{\prime}-k)}$ ${}\mathopen{}\mathclose{{}\left(1-\frac{1}{n}}\right)^{n-(A_{i}-(n^{\prime}-k))}\mathopen{}\mathclose{{}\left(\frac{1}{n}}\right)^{A_{i}-(n^{\prime}-k)}\geq\frac{1}{e}\mathopen{}\mathclose{{}\left(\frac{A_{i}}{n(A_{i}-(n^{\prime}-k))}}\right)^{A_{i}-(n^{\prime}-k)}$
-$\geq{}\frac{1}{e}\mathopen{}\mathclose{{}\left(\frac{n^{\prime}-k}{nk}}\right)^{k}\geq\frac{1}{e}\mathopen{}\mathclose{{}\left(\frac{n^{\prime}}{2nk}}\right)^{k}=\frac{1}{e}\mathopen{}\mathclose{{}\left(\frac{1}{mk}}\right)^{k},$
-where the last two inequalities use $n^{\prime}-k<A_{i}\leq n^{\prime}$ and $k\leq n^{\prime}/2$ respectively, and the last equality uses $n^{\prime}=2m/n$.
-Note that $A_{i}\in[0..k-1]\cup[n^{\prime}-k+1..n^{\prime}]$ holds for all $x\in P$ before such $y$ is generated.
-Then we know that $y$ will be generated in expected $e(mk)^{k}$ iterations.
-Since such $y$ cannot be dominated by any solution in $P$, we know that $y\in F_{1}$.
-From Lemma 4, we know that once $y$ is generated, $f(y)$ will be covered by the next population and by all future populations, and thus future populations will contain at least one solution $z$ with $|z_{B_{i}}|_{1}\in[k..n^{\prime}-k]$.
-Since an inner Pareto optimum is reached when $y$ with $|y_{B_{1}}|_{1},\dots,|y_{B_{m/2}}|_{1}\in[k..n^{\prime}-k]$ is generated, we know that the expected iterations to reach an inner Pareto optimum is at most $\frac{m}{2}e(mk)^{k}=\frac{1}{2}em^{k+1}k^{k}$.
-∎
-We consider the next stage of covering all inner Pareto front points.
-The following lemma shows the runtime.
-The key of the proof is that one uncovered Pareto front point can be easily generated from a neighbor Pareto front if it is covered.
+Then after at most $e\mu(mk)^{k}(1+\ln m)$ iterations in expectation, the population (and also the populations afterwards) contains at least one inner Pareto optimum.
+In the next lemma, we consider the stage of covering all inner Pareto front points once at least one such point is in the population.
+The key to the proof is that as long as we have not yet discovered the full inner Pareto front, there always exists a missing Pareto front point that is a neighbor of a point that is already covered by the population.
+Hence choosing the right parent and flipping the right single bit suffices to cover the desired point on the Pareto front.
 Lemma 6.
@@
 Then after at most $en\mu M$ iterations in expectation, all inner Pareto front points are covered.
-Proof.
-Let $u\in F^{*}$ be an uncovered inner Pareto front point that has some neighbor covered by the population $P$.
-Since such a neighbor is also a Pareto front point, we know that it is maintained in all future generations.
-Since the probability to generate an offspring with its function value of $u$ is at least $\mathopen{}\mathclose{{}\left(1-\frac{1}{n}}\right)^{n-1}\frac{1}{n\mu}\geq\frac{1}{en\mu}$, we know that the expected iterations to witness such an offspring is at most $n\mu$ in expectation.
-From Lemma 4, we know that once such an offspring is generated, $u$ will be covered in all future populations.
-Noting that there are at most $M$ inner Pareto front points, we know that in expected at most $en\mu M$ iterations, all Pareto front points are covered.
-∎
 The last stage is to cover the remaining Pareto front points.
 The following lemma shows the runtime.
-The key of the proof is that we can divide the Pareto front into several levels and any individual in the current $i$-th level can be easily generated from a point in the $(i-1)$-th level.
+The key of the proof is that we can divide the Pareto front into several levels and any individual in the $i$-th level can be generated from a point in the $(i-1)$-th level by flipping the right $k$ bits.
 Lemma 7.
@@
 Then after at most $eM\mu n^{k}$ iterations in expectation, the full Pareto front is covered.
-Proof.
-We call $(k,n^{\prime}+k)$ or $(n^{\prime}+k,k)$ extreme block.
-We divide the Pareto front points into $m/2+1$ levels, where each point in $i$-th level contains $i$ extreme blocks.
-Obviously, all inner Pareto front points belong to the 00-th level.
-Note that any point in $j$-th level ($j\geq 1$) can be generated by a certain point in $(j-1)$-th level via flipping specific $k$ bits with other bits unchanged.
-Hence, the probability to generate any point in $j$-th level condition on that all $(j-1)$-th level points are covered is at least
-$\frac{1}{\mu}\mathopen{}\mathclose{{}\left(1-\frac{1}{n}}\right)^{n-k}\mathopen{}\mathclose{{}\left(\frac{1}{n}}\right)^{k}\geq\frac{1}{e\mu n^{k}}.$
-From Lemma 4, we know that such a Pareto front point is reached in at most $e\mu n^{k}$ in expectation, and will survive in all future populations.
-Consider the process of covering all non-inner Pareto front points in the order of covering the $1$-st level, …, $(m/2)$-th level.
-Since the 00-th level points are already covered from our assumption, and since there are at most $M$ non-inner Pareto front points, we know that the remaining levels are covered in at most $eM\mu n^{k}$ iterations in expectation.
-∎
 Summing up the runtime of all stages from Lemmas 5 to 7, we have the following theorem for the runtime of the full coverage of the Pareto front.
@@
 Consider using SMS-EMOA with $\mu\geq M$ to optimize the $m$OJZJ problem.
-Then after at most $\frac{1}{2}em^{k+1}k^{k}+n\mu M+eM\mu n^{k}$ iterations ($\mu+\frac{1}{2}em^{k+1}k^{k}+n\mu M+eM\mu n^{k}$ function evaluations) in expectation, the full Pareto front is covered.
-Comparing with the NSGA-II’s exponential lower bound to cover the special case of $m$OJZJ, Theorem 8 shows the efficiency of the SMS-EMOA for many objectives, and indicates that for the NSGA-II, the non-dominated sorting is cool, but the crowding distance as the secondary criterion is deficient.
-4.3 Runtime of GSMEO
-For an overall comparison, we also show the runtime of the GSMEO on the $m$OJZJ.
+Then after at most $e\mu(mk)^{k}(1+\ln m)+e\mu Mn+e\mu Mn^{k}$ iterations ($\mu+e\mu(mk)^{k}(1+\ln m)+e\mu Mn+e\mu Mn^{k}$ function evaluations) in expectation, the full Pareto front is covered.
+Since $k\leq n^{\prime}/2=n/m$ and $M\geq 2^{m/2}$, we easily see that both runtime expressions are $O(\mu Mn^{k})$, even when allowing $k$, $m$ and $\mu$ to depend on $n$.
+Runtime of the GSEMO on $m$OJZJ
+Since our arguments above can easily be extended to analyze the runtime of the GSEMO on the $m$OJZJ, we quickly do so for reasons of completeness.
 The GSEMO is a multiobjective counterpart of the ${(1+1)}$ EA.
-The initial population contains a uniformly and randomly generated solution.
-Then in each iteration, a solution is uniformly at random picked from the population to generate an offspring with standard bit-wise mutation.
-If this offspring cannot be dominated by any solution in the population, it will enter into the population and removes all solutions weakly dominated by it.
-It is not difficult to see all solutions in the population of the GSEMO are mutually non-dominated.
-Hence, the population size of the GSEMO is at most $M$ from Lemma 3.
-The proof ideas are similar to Lemmas 5 to 7.
-The only difference is that SMS-EMOA needs Lemma 4 to ensure that all function values covered in $F_{1}$ will remain covered by the next population, but for GSEMO, all such function values are surely covered by the next population as $F_{1}$ contains all mutually non-dominated individuals whose function values will not be removed in the survival selection.
-Together with the upper bound $M$ of the population size, we have the following theorem.
+The initial population consists of a single randomly generated solution.
+In each iteration, a solution is uniformly at random picked from the population to generate an offspring via standard bit-wise mutation.
+If this offspring is not dominated by any solution in the population, it is added to the population and all solutions weakly dominated by it are removed.
+It is not difficult to see that all solutions in the population of the GSEMO are mutually non-dominated.
+Hence, the population size of the GSEMO is at most $M$ by Lemma 3.
+When analyzing the runtime of the GSEMO on $m$OJZJ, the main difference is that SMS-EMOA requires a statement like Lemma 4 to ensure that previous progress is not lost via unlucky selection decisions.
+For the GSEMO, this property follows immediately from the selection mechanism, which keeps all non-dominated solutions.
+Together with the upper bound $M$ of the population size, we obtain the following theorem.
 Theorem 9.
 Let $k\leq n^{\prime}/2$.
-Consider using GSEMO to optimize the $m$OJZJ problem.
-Then after at most $\frac{1}{2}em^{k+1}k^{k}+nM^{2}+eM^{2}n^{k}$ iterations in expectation, the full Pareto front is covered.
-5 Reduced Impact of Stochastic Population Updates
-Bian et al.
-[BZLQ23] proposed a stochastic population update mechanism for the SMS-EMOA and proved that, somewhat unexpectedly given the state of the art, it can lead to significant performance gains.
-More precisely, they proved that the classic SMS-EMOA with reasonable population size $\mu=\Theta(n)$ solves the OJZJ problem in an expected number of $O(n^{k+2})\cap\Omega(n^{k+1})$ function evaluations, whereas for the SMS-EMOA with stochastic population update, $O(n^{k+2}\min\{1,2^{-k/4}n\})$ suffice.
-Hence, for $k=\omega(\log n)$, a super-polynomial speed-up is shown.
-In this section, we extend the analysis of [BZLQ23] to $m$ objectives.
+Consider using the GSEMO to optimize the $m$OJZJ problem.
+Then after an expected number of at most $eM(mk)^{k}(1+\ln m)+eM^{2}n+eM^{2}n^{k}$ iterations (or $1+eM(mk)^{k}(1+\ln m)+eM^{2}n+eM^{2}n^{k}$ fitness evaluations), the full Pareto front is covered.
+Reduced Impact of Stochastic Population Update
+Bian, Zhou, Li, and Qian (2023) proposed a stochastic population update mechanism for the SMS-EMOA and proved that, somewhat unexpectedly given the state of the art, it can lead to significant performance gains.
+More precisely, it was proven that the classic SMS-EMOA with reasonable population size $\mu=\Theta(n)$ solves the bi-objective OJZJ problem in an expected number of $O(n^{k+2})\cap\Omega(n^{k+1})$ function evaluations, whereas for the SMS-EMOA with stochastic population update, $O(n^{k+2}\min\{1,2^{-k/4}n\})$ suffice.
+Hence, for $k=\omega(\log n)$, a super-polynomial speed-up was shown.
+In this section, we extend the analysis of (Bian et al.
+2023) to $m$ objectives.
 Unfortunately, we will observe that the larger population sizes necessary here reduce the impact of the stochastic population update.
-As in [BZLQ23], we have no proven lower bounds for the SMS-EMOA with stochastic population update, but our upper-bound proofs suggest that the reduced effect of the stochastic population update on the runtime guarantee is real, that is, the impact on the true runtime is diminishing with the larger population sizes necessary in the many-objective setting.
-5.1 Stochastic Population Update
+As in (Bian et al.
+2023), we have no proven tight lower bounds for the SMS-EMOA with stochastic population update, but our upper-bound proofs suggest that the reduced effect of the stochastic population update on the runtime guarantee is real, that is, the impact on the true runtime is diminishing with the larger population sizes necessary in the many-objective setting.
+Stochastic Population Update
 The rough idea of the stochastic selection proposed by Bian et al.
-[BZLQ23] is that a random half of the individuals survive into the next generation regardless of their quality.
+(2023) is that a random half of the individuals survive into the next generation regardless of their quality.
 The individual to be discarded is chosen from the remaining individuals according to non-dominated sorting and hypervolume contribution as in the classic SMS-EMOA.
-This approach resembles the random mixing of two acceptance operators in the Move Acceptance Hyper-Heuristic studied recently [LOW23, DDLS23].
+This approach resembles the random mixing of two acceptance operators in the Move Acceptance Hyper-Heuristic studied recently (Lehre and Özcan 2013; Lissovoi, Oliveto, and Warwicker 2023; Doerr et al.
+2023).
 More precisely, after generating the offspring $x^{\prime}$, the SMS-EMOA with stochastic selection chooses from $P_{t}\cup\{x^{\prime}\}$ a set $R^{\prime}$ of $\lfloor(\mu+1)/2\rfloor$ solutions randomly with replacement.
 The individual $z^{\prime}$ to be removed from $P_{t}\cup\{x^{\prime}\}$ is then determined via non-dominated sorting and hypervolume contribution applied to $R^{\prime}$ only.
-We note that with this mechanism, even the worst solution can enter the next generation with probability at least $1/2$.
-5.2 Runtime
-To establish the runtime result of the SMS-EMOA with stochastic population update, we separate the key facts into several lemmas for a clear demonstration.
-The following two lemmas 10 and 11 respectively show the general survival probability of each individual in the combined population, and the survival of the individual with rank one.
+We note that with this mechanism, even the worst solution enters the next generation with probability at least $1/2$.
+Runtime
+To analyze the runtime of the SMS-EMOA with stochastic population update, we first derive and formulate separately two insights on the survival of solutions.
+The following Lemmas 10 is an elementary consequence of the algorithm definition, whereas Lemma 11 builds on Lemma 4.
 Lemma 10.
-Consider using SMS-EMOA with stochastic population update to optimize the $m$OJZJ problem.
-For any $x\in R_{t}$ for the current generation $t$, we have that there is $y$ with $f(y)=f(x)$ in $P_{t+1}$ with probability at least $1/2$.
-Proof.
-We simply calculate the probability of $x$ being removed after the population update.
-Since $x$ is possible to be removed only if it is selected in $R^{\prime}$, we know that it will be removed with probability at most $\lfloor|R^{\prime}|/2\rfloor/|R^{\prime}|\leq 1/2$, and hence $f(x)$ will be maintained in $P_{t+1}$ with probability at least $1/2$.
-∎
+Consider using the SMS-EMOA with stochastic population update to optimize the $m$OJZJ problem.
+For any iteration $t$ and $x\in R_{t}$, we have that there is a $y$ with $f(y)=f(x)$ in $P_{t+1}$ with probability at least $1/2$.
 Lemma 11.
-Consider using SMS-EMOA with stochastic population update and with $\mu\geq 2(M+1)$ to optimize the $m$OJZJ problem.
-For any $x\in F_{1}$ for the current generation $t$, we have that there is $y$ with $f(y)=f(x)$ in $P_{t+1}$.
-Specifically, any Pareto front point will be maintained in all future populations once it is reached.
-Proof.
-Let $x\in F_{1}$. If $\cup_{i\geq 2}F_{i}\neq\emptyset$, then the removal will happen in $\cup_{i\geq 2}F_{i}$, and $x\in P^{t+1}$.
-If all solutions are in $F_{1}$, from Lemma 3, we know that $|f(F_{1})|\leq M$.
-Since there are $|R^{\prime}|\geq\lfloor(\mu+1)/2\rfloor\geq M+1>M$, we know that there exist $z_{1},z_{2}\in F_{1}$ such that $f(z_{1})=f(z_{2})$.
-Let $S_{x}=\{z\in R^{\prime}_{t}\mid f(x)=f(z)\}$.
-Then $|S_{x}|\geq 1$.
-If $|S_{x}|=1$, then $\Delta(x,F_{1})>0=\Delta(z_{1},F_{1})$, and thus $x\in P_{t+1}$.
-If $|S_{x}|>1$, since the survival selection only removes one solution, then there exists $y\in S_{x}$, such that $y\in P_{t+1}$.
-Since any Pareto optimum is always in $F_{1}$, the latter part of this lemma holds. ∎
-Recalling the proof in Lemmas 5 and 6, we know that these lemmas hold merely relies on that $f(F_{1})$ value will be maintained once it is reached.
-Hence, from Lemma 11, we have the following results for obtaining a population containing at least one inner Pareto optimum, and for covering all inner Pareto front points starting from a population with at least one inner Pareto optimum.
+Consider using the SMS-EMOA with stochastic population update and with $\mu\geq 2(M+1)$ to optimize the $m$OJZJ problem.
+If at some time $t$ the combined parent and offspring population $R_{t}$ of the SMS-EMOA contains some solution $x$, then at any later time $s>t$ the parent population $P_{s}$ contains a solution $y$ such that $y\succeq x$.
+Since the proofs of Lemmas 5 and 6 for the classic SMS-EMOA mostly relied on elementary properties of standard bit-wise mutation and on the survival guarantee of Lemma 4, we can now use the survival guarantee of Lemma 11 to obtain analogous results for the SMS-EMOA with stochastic selection (at the price of requiring essentially twice the population size).
+This yields the following estimates for the time to obtain at least one inner Pareto optimum and the time to cover all inner Pareto front points starting from a population with at least one inner Pareto optimum.
 Lemma 12.
-Let $k\leq n^{\prime}/2$. Consider using SMS-EMOA with stochastic population update and with $\mu\geq 2(M+1)$ to optimize the $m$OJZJ problem. Then
-  • after at most $\frac{1}{2}em^{k+1}k^{k}$ iterations in expectation, the population (and also the populations afterwards) contains at least one inner Pareto optimum;
-  • after at most $en\mu M$ iterations in expectation, all inner Pareto front points are covered starting from a population with at least one inner Pareto optimum.
+Let $k\leq n^{\prime}/2$. Consider using the SMS-EMOA with stochastic population update and with $\mu\geq 2(M+1)$ to optimize the $m$OJZJ problem. Then
+  • after at most $e\mu(mk)^{k}(1+\ln m)$ iterations in expectation, the population (and also the populations afterwards) contains at least one inner Pareto optimum;
+  • after another at most $en\mu M$ iterations in expectation, all inner Pareto front points are covered.
 Now we consider the runtime for the full coverage of the Pareto front after all inner Pareto front points are covered.
-Similar to the proof of Lemma 7 for the original SMS-EMOA, we partition the Pareto front points into $m/2+1$ levels based on the number of extreme blocks inside.
-The difference here is that we will consider the effect of the gap points, similar to [BZLQ23].
+Similar to the proof of Lemma 7 for the original SMS-EMOA, we partition the Pareto front points into $m/2+1$ levels based on the number of extreme blocks they contain.
+The difference here is that we consider the effect of the gap points, similar to Bian et al.
+(2023).
 Lemma 13.
 Let $k\leq n^{\prime}/2$.
-Consider using SMS-EMOA with stochastic population update and with $\mu\geq 2(M+1)$ to optimize the $m$OJZJ problem.
+Consider using the SMS-EMOA with stochastic population update and with $\mu\geq 2(M+1)$ to optimize the $m$OJZJ problem.
 Assume that the current population covers all inner Pareto front points.
 Then after at most $O(\min\{\frac{\mu k^{1/2}}{2^{k/2}},1\}M\mu n^{k})$ iterations in expectation, the full Pareto front is covered.
-Proof.
-We use the same definitions of the $j$-th ($j\in[0..m/2]$) level where each Pareto front point contains $j$ extreme blocks, and note that 00-th level points are covered.
-Let $P_{t}$ be the current population.
-We consider generating any point in $j$-th level condition on that all $(j-1)$-th level points are covered.
-Let $v$ be an uncovered Pareto front point in the $j$-th level.
-Then there exists a block $B$ that $v$ has the value of $(k,n^{\prime}+k)$ or $(n^{\prime}+k,k)$.
-W.l.o.g., let the block value of $v$ is $v_{B}=(k,n^{\prime}+k)$.
-Besides, there exists a set of points $U$ in the $(j-1)$-th level such that each $u\in U$ is the same as $v$ except that $u$ has the value of $(a,n^{\prime}+2k-a)$ with $a\in[2k..n^{\prime}]$ in the above block $B$.
-Let $S_{v}=\{f(z)\mid z\in P,z\text{\leavevmode\nobreak\ is the same as\leavevmode\nobreak\ }v\text{\leavevmode\nobreak\ except\leavevmode\nobreak\ }z_{B}\neq(k,n^{\prime}+k)\}$.
-Obviously, $y:=\min_{z\in S_{v}}z_{B}\leq 2k$.
-Let $A_{1}$ denote the event $k<y\leq\frac{3}{2}k$, $A_{2}$ be the event $y\geq\frac{7}{4}k$, and $A_{3}$ be $\frac{3}{2}k<y<\frac{7}{4}k$.
-Consider the potential function
-$h(P_{t})=\begin{cases}0,&{}\text{if $v\in P_{t}$},\\ e\mu n^{k/2},&{}\text{if $A_{1}$},\\ e\mu n^{k/2}+1,&{}\text{if $A_{2}\cup A_{3}$}.\end{cases}$
-For $A_{1}$, we pessimistically consider the decrease of $h$ by generating $v$ and the increase of $h$ by kicking out a corresponding solution of $y$. Hence, we have
-$E[h$ ${}(P_{t})-h(P_{t+1})\mid P_{t},A_{1}]\geq\frac{1}{\mu}\frac{1}{n^{y-k}}\mathopen{}\mathclose{{}\left(1-\frac{1}{n}}\right)^{n-(y-k)}e\mu n^{k/2}-\frac{1}{2}\cdot 1\geq 1-\frac{1}{2}=\frac{1}{2}.$
-For $A_{2}$, $h$ will not increase, and there are $y-k\in[\frac{3}{4}k,k]$ ones in the block of a corresponding solution of $y$.
-We pessimistically consider the decrease of $h$ by flipping $y-k-\frac{1}{2}k$ ones and keeping other bits unchanged of the corresponding solution.
-Hence, we have
-$E[h$ ${}(P_{t})-h(P_{t+1})\mid P_{t},A_{2}]\geq\frac{1}{\mu}\frac{\binom{y-k}{y-k-k/2}}{n^{y-k-k/2}}\mathopen{}\mathclose{{}\left(1-\frac{1}{n}}\right)^{n-(y-k-k/2)}\frac{1}{2}$
-[…]
-where the last inequality uses $\frac{3}{4}k\leq y-k\leq k$.
-For $A_{3}$, $h$ will not increase, and there are $y-k\in(\frac{1}{2},\frac{3}{4}k)$ ones in the block of a corresponding solution of $y$.
-We pessimistically consider the decrease of $h$ by flipping $\frac{1}{4}k$ ones and keeping other bits unchanged of the corresponding solution.
-Hence, we have
-$E[h$ ${}(P_{t})-h(P_{t+1})\mid P_{t},A_{3}]\geq\frac{1}{\mu}\frac{\binom{y-k}{k/4}}{n^{k/4}}\mathopen{}\mathclose{{}\left(1-\frac{1}{n}}\right)^{n-k/4}\frac{1}{2}\geq\frac{1}{2e\mu}\frac{\binom{k/2}{k/4}}{n^{k/4}}.$
-Therefore, we have
-$E[h$ ${}(P_{t+1})-h(P_{t})\mid P_{t},v\notin P_{t}]=\sum_{i=1}^{3}\Pr[A_{i}]E[h(P_{t+1})-h(P_{t})\mid P_{t},A_{i}]$
-[…]
-[…]
-where the third inequality uses $\binom{k/2}{k/4}\leq\binom{\frac{3}{4}k}{k/2}$ and $n^{k/2}\geq n^{k/4}$, and the last inequality uses $k\geq 1$.
-Since the maximum $h$ is $e\mu n^{k/2}+1$, we know that the Pareto front point $v$ will be reached in at most
-$(e\mu n^{k/2}+1)\frac{{4e\mu n^{k/2}\sqrt{k}}}{4^{k/4}}=O\mathopen{}\mathclose{{}\left(\frac{\mu k^{1/2}}{2^{k/2}}\mu n^{k}}\right)$
-in expectation, and will survive in all future populations.
-Consider the process of covering all non-inner Pareto front points in the order of covering the $1$-st level, …, $(m/2)$-th level.
-Since the 00-th level points are already covered from our assumption, and since there are at most $M$ non-inner Pareto front points, we know that the remaining levels are covered in at most $O(\frac{\mu k^{1/2}}{2^{k/2}}M\mu n^{k})$ iterations in expectation.
-With the same argument in the proof of Lemma 7, we know that the expected number of iterations to cover the full Pareto front is at most $eM\mu n^{k}$.
-Hence, we proved this lemma.
-∎
 Combining Lemmas 12 and 13, we have the runtime of the SMS-EMOA with stochastic population update in the following theorem.
@@
 Consider using SMS-EMOA with $\mu\geq 2(M+1)$ and stochastic population update to optimize the $m$OJZJ problem.
-Then after at most $\frac{1}{2}em^{k+1}k^{k}+en\mu M+O(\min\{\frac{\mu k^{1/2}}{2^{k/2}},1\}M\mu n^{k})$ iterations in expectation, the full Pareto front is covered.
-Theorem 14 shows that the best runtime for the SMS-EMOA with stochastic population update is $O(\min\{\frac{Mk^{1/2}}{2^{k/2}},1\}M^{2}n^{k})$ when we choose $\mu=\Theta(M)$.
-Recall that the runtime of the original SMS-EMOA is $O(M^{2}n^{k})$ for $\mu=\Theta(M)$ from Theorem 8.
-We know the asymptotic speed-up is $\Theta(\max\{\frac{2^{k/2}}{Mk^{1/2}},1\})$ with $M=(2n/m-2k+3)^{m/2}$.
-Hence, when the objective number is not small, like at least the same size of $k$, then the speed-up is only $\Theta(1)$.
-We note that it is not contradictory to the nice speed-up obtained for two objectives in [BZLQ23] but adds a limitation situation for the stochastic population update.
-6 Heavy-Tailed Mutation Helps
+Then after at most $e\mu(mk)^{k}(1+\ln m)+en\mu M+O(\min\{\frac{\mu k^{1/2}}{2^{k/2}},1\}M\mu n^{k})$ iterations in expectation, the full Pareto front is covered.
+Comparing the runtime guarantees or Theorem 8 (classic SMS-EMOA) and Theorem 14 (SMS-EMOA with stochastic selection), we see that stochastic selection can at most lead to a speed-up by a factor of order $2^{k/2}/\mu k^{1/2}$.
+Now $\mu$ is at least $\Omega(M)=\Omega((2n/m-2k+3)^{m/2})$.
+Consequently, the advantage of stochastic selection is rapidly decreasing with growing numbers of objectives and vanishes, e.g., when $m\geq k$.
+Heavy-Tailed Mutation Helps
 In the previous section, we saw that the advantages of a stochastic population update do not well generalize from bi-objective optimization to many-objective optimization.
@@
 We shall prove that the $k^{\Omega(k)}$ factor speed-up observed in bi-objective optimization extends to many objectives.
-6.1 Heavy-Tailed Mutation
-Different from the standard bit-wise mutation, the heavy-tailed mutation flips each bit of the solution with probability $\alpha/n$ independently, where $\alpha$ follows a power-law distribution with parameter $\beta$ defined as follows.
+Heavy-Tailed Mutation
+Different from the standard bit-wise mutation operator, which flips each bit independently with probability $1/n$, the heavy-tailed mutation operator proposed in (Doerr et al.
+2017) flips each bit independently with probability $\alpha/n$, where $\alpha$ follows a power-law distribution with parameter $\beta$.
+The number $\alpha$ is sampled anew in each application of the heavy-tailed mutation operator.
+The underlying power-law is defined as follows.
 Definition 15.
 Let $n\in\mathbb{N}$ and $\beta>1$.
-We say $\alpha$ follows the power-law distribution if for all $i\in[1..n/2]$, $\Pr[\alpha=i]=\mathopen{}\mathclose{{}\left(\sum_{j=1}^{n/2}k^{-\beta}}\right)^{-1}i^{-\beta}$.
-This power-law distribution has a good feature that a point with Hamming distance of $j$ of a given bit-string will be generated with a probability of $\Omega(j^{-\beta})$.
-It will facilitate the jump with a larger size and facilitate escaping from the local optimum.
-The heavy-tailed mutation has resulted in an asymptotic performance speed-up by a factor of $k^{\Omega(k)}$ of the $(1+1)$ EA on the single-objective Jump function with gap size $k$ [DLMN17].
-For the bi-objective OJZJ, the speed-up of $k^{\Omega(k)}$ is also witnessed when it is applied to the GSEMO [DZ21] and to the NSGA-II [DQ23a].
-Several other positive theoretical results exist for this heavy-tailed mutation, or more generally, other heavy-tailed parameter choises [FQW18, WQT18, QGWF21, COY21, DELQ22, ABD22, DGI22, DR23].
+We say $\alpha$ follows a power-law distribution with (negative) exponent $\beta$ if for all $i\in[1..n/2]$, we have $\Pr[\alpha=i]=\left(\sum_{j=1}^{n/2}k^{-\beta}\right)^{-1}i^{-\beta}$.
+The power-law feature extends to the heavy-tailed mutation operator in the sense that it generates an offspring with Hamming distance $j$ from the parent with probability $\Omega(j^{-\beta})$.
+This facilitates larger jumps in the search space, and thus escaping from a local optimum.
+The heavy-tailed mutation operator has resulted in asymptotic performance gains by a factor of $k^{\Omega(k)}$ for the ${(1+1)}$ EA optimizing single-objective (classic and generalized) Jump functions with gap size $k$ (Doerr et al.
+2017; Bambury, Bultel, and Doerr 2021).
+For the bi-objective OJZJ benchmark, again a speed-up of $k^{\Omega(k)}$ was proven when optimized via the GSEMO (Zheng and Doerr 2023d) and the NSGA-II (Doerr and Qu 2023a).
+Several other positive theoretical results exist for this heavy-tailed mutation, or more generally, other heavy-tailed parameter choices (Friedrich, Quinzan, and Wagner 2018; Wu, Qian, and Tang 2018; Quinzan et al.
+2021; Corus, Oliveto, and Yazdani 2021; Dang et al.
+2022; Antipov, Buzdalov, and Doerr 2022; Doerr, Ghannane, and Ibn Brahim 2022; Doerr and Rajabi 2023).
 Prior to this work, no theoretical analysis of the heavy-tailed mutation operator in many-objective optimization existed.
-6.2 Runtime
+Runtime
 We now analyze the runtime of the SMS-EMOA using the heavy-tailed mutation operator instead of bit-wise mutation, the standard choice for this algorithm.
-Since the heavy-tailed mutation does not change the survival selection of the original SMS-EMOA, we easily have the following corollary about the survival of the individuals with rank one from Lemma 4.
+Since the heavy-tailed mutation does not change the survival selection of the original SMS-EMOA, we immediately have the following result on the survival of the individuals.
 Corollary 16.
-Consider using SMS-EMOA with heavy-tailed mutation and with $\mu\geq M$ to optimize the $m$OJZJ problem.
-For any $x\in F_{1}$ for the current generation $t$, we have that there is $y$ with $f(y)=f(x)$ in $P_{t+1}$.
-Specifically, any Pareto front point will be maintained in all future populations once it is reached.
-With similar proof ideas in Lemma 5 to 7 but noticing the probability of generating a solution with a specific Hamming distance, we obtain the following lemmas for reaching the first inner Pareto optimum (Lemma 17), covering all inner Pareto front points (Lemma 18), and covering the remaining Pareto front points (Lemma 19).
-Lemma 17.
-Let $k\leq n^{\prime}/2$.
-Consider using SMS-EMOA with heavy-tailed mutation and with $\mu\geq M$ to optimize the $m$OJZJ problem.
-Then after at most $C_{n/2}^{\beta}\binom{n}{k}mk^{\beta}$ iterations in expectation, the population (and also the populations afterwards) contains at least one inner Pareto optimum.
-Proof.
-We use similar arguments of Lemma 5 except following probability calculations.
-The probability to generate a $y$ with $|y_{B_{i}}|_{1}\in[k..n^{\prime}-k]$ from $x$ with $A_{i}<k$ or with $A_{i}>n^{\prime}-k$ is at least $\binom{n}{k}^{-1}P_{k}^{\beta}\geq\binom{n}{k}^{-1}\mathopen{}\mathclose{{}\left(C_{n/2}^{\beta}}\right)^{-1}k^{-\beta}$.
-Hence such $y$ is generated in expected $\binom{n}{k}C_{n/2}^{\beta}k^{\beta}$ iterations.
-Hence the expected iterations to reach an inner Pareto optimum is at most $\frac{m}{2}\binom{n}{k}C_{n/2}^{\beta}k^{\beta}=\binom{n}{k}C_{n/2}^{\beta}mk^{\beta}$ iterations.
-∎
-Lemma 18.
-Let $k\leq n^{\prime}/2$.
-Consider using SMS-EMOA with heavy-tailed mutation and with $\mu\geq M$ to optimize the $m$OJZJ problem.
-Assume that the current population contains at least one inner Pareto optimum.
-Then after at most $\frac{e\beta}{\beta-1}n\mu M$ iterations in expectation, all inner Pareto front points are covered.
-Proof.
-We use similar arguments of Lemma 6 except following probability calculation.
-The probability to generate an offspring with its function value of an uncovered inner Pareto front point from a neighbor covered by the population is at least $\frac{1}{n\mu}P_{1}^{\beta}=\frac{\beta-1}{e\beta n\mu}$.
-∎
-Lemma 19.
-Let $k\leq n^{\prime}/2$.
-Consider using SMS-EMOA with heavy-tailed mutation and with $\mu\geq M$ to optimize the $m$OJZJ problem.
-Assume that the current population covers all inner Pareto front points.
-Then after at most $\binom{n}{k}C_{n/2}^{\beta}\mu Mk^{\beta}$ iterations in expectation, the full Pareto front is covered.
-Proof.
-We use similar arguments of Lemma 7 except following probability calculation.
-The probability to generate any point in $j$-th level condition on that all $(j-1)$-th level points are covered is at least
-$\frac{1}{\mu}\binom{n}{k}^{-1}P_{k}^{\beta}\geq\binom{n}{k}^{-1}\mathopen{}\mathclose{{}\left(C_{n/2}^{\beta}}\right)^{-1}\frac{1}{\mu}k^{-\beta}.\qed$
-Considering the whole process for the full coverage of the Pareto front from Lemmas 17 to 19, we have the following runtime result for the SMS-EMOA with heavy-tailed mutation.
+The assertion of Lemma 4 extends to the SMS-EMOA with heavy-tailed mutation.
+With proof ideas similar to those in Lemma 5 to 7, but noticing the different probabilities of generating solutions with a specific Hamming distance, we obtain the following runtime guarantee.
 Theorem 17.
-Let $k\leq n^{\prime}/2$.
-Consider using SMS-EMOA with $\mu\geq M$ to optimize the $m$OJZJ problem.
+Let $k\leq n^{\prime}/2$ and $\beta>1$.
+Consider using SMS-EMOA with $\mu\geq M$ and with heavy-tailed mutation to optimize the $m$OJZJ problem.
 Then after at most $O(M\mu k^{\beta-0.5-k}n^{k})$ iterations or function evaluations in expectation, the full Pareto front is covered.
-Theorem 20 shows the best asymptotic upper bound of the runtime of the SMS-EMOA with heavy-tailed mutation is $O(M^{2}n^{k}k^{\beta-0.5-k})$.
-Comparing with the runtime of $O(M^{2}n^{k})$ for $\mu=\Theta(M)$ for the original SMS-EMOA from Theorem 8, we see a speed-up of $k^{k+0.5-\beta}$, which is also the speed-up reached for the single objective and two objectives.
-7 SMS-EMOA Also Performs Well for Bi-objective Optimization
+Compared to the runtime guarantee of Theorem 8 for the original SMS-EMOA, the guarantee of Theorem 17 above for the SMS-EMOA with heavy-tailed mutation is by a factor of asymptotically $k^{k+0.5-\beta}$ stronger.
+The SMS-EMOA Also Performs Well for Bi-objective Optimization
 The above sections discussed the runtime of the SMS-EMOA for many objectives, and showed that it performs well for many objectives, which is different from the original NSGA-II.
-Since the only theory paper [BZLQ23] merely considers its performance for the bi-objective OJZJ, in this section, we will analyze its runtime for other well-known bi-objective benchmarks to have a more comprehensive theoretical understanding.
-7.1 OneMinMax and LOTZ
-OneMinMax [GL10] (also the CountingOnesCountingZeroes [${}^{+}$02]) and LOTZ [LTZ04] (also the weighted version WLPTNO [QYZ13]) are the bi-objective counterpart of the well-analyzed single-objective OneMax and LeadingOnes respectively.
-For OneMinMax, two objectives count the number of ones and zeros in a given bit-string respectively.
-For LOTZ, the first objective counts the continuous ones starting from the first bit position, and the second one counts the continuous zeros starting from the last bit position.
-Both benchmarks with a simple and clear structure facilitate the fundamental theoretical understanding of the MOEAs.
-We note that there are other benchmarks [QTZ16, LZZZ16, DOSS23b], but here we focus on the two kinds of most popular benchmarks for analysis.
-Here are formal definitions of the OneMinMax and LOTZ.
-Definition 21 ([GL10]).
+Since the only theory paper (Bian et al.
+2023) on the SMS-EMOA merely considers its performance on the bi-objective OJZJ problem, to broaden our understanding of this algorithm we now analyze its runtime for the two most prominent bi-objective benchmarks OneMinMax and LOTZ.
+OneMinMax and LOTZ
+The OneMinMax benchmark introduced by Giel and Lehre (2010) (and also the CountingOnesCountingZeroes benchmark proposed by Laumanns et al.
+(2002)) are bi-objective counterparts of the famous single-objective OneMax benchmark.
+Likewise, the LOTZ benchmark defined by Laumanns, Thiele, and Zitzler (2004) (also the weighted version WLPTNO of Qian, Yu, and Zhou (2013)) are bi-objective counterparts of the classic single-objective LeadingOnes benchmark.
+For OneMinMax, the two objectives count the number of ones and zeros in a given bit-string, respectively.
+For LOTZ, the first objective is the number of contiguous ones starting from the first bit position, and the second objective is the number of contiguous zeros starting from the last bit position.
+Both benchmarks with their simple and clear structure facilitate the fundamental theoretical understanding of MOEAs.
+We note that there are other benchmarks used in the mathematical analysis of MOEAs (see, e.g.
+Horoba and Neumann 2008; Brockhoff et al.
+2009; Qian, Tang, and Zhou 2016; Li et al.
+2016; Dang et al.
+2023b), but clearly OneMinMax and LOTZ are the most prominent ones.
+Here are the formal definitions of OneMinMax and LOTZ.
+Definition 18 (Giel and Lehre (2010)).
 Let $n\in\mathbb{N}$ be the problem size. The OneMinMax function :$:\{0,1\}^{n}\rightarrow\mathbb{R}$ is defined by
-$\textsc{OneMinMax}(x)=\mathopen{}\mathclose{{}\left(\sum_{i=1}^{n}(1-x_{i}),\sum_{i=1}^{n}x_{i}}\right)$
+$\textsc{OneMinMax}(x)=\left(\sum_{i=1}^{n}(1-x_{i}),\sum_{i=1}^{n}x_{i}\right)$
 for any $x=(x_{1},\dots,x_{n})\in\{0,1\}^{n}$.
-Definition 22 ([LTZ04]).
+Definition 19 (Laumanns, Thiele, and Zitzler (2004)).
 Let $n\in\mathbb{N}$ be the problem size. The LOTZ function :$:\{0,1\}^{n}\rightarrow\mathbb{R}$ is defined by
-$\textsc{LOTZ}(x)=\mathopen{}\mathclose{{}\left(\sum_{i=1}^{n}\prod_{j=1}^{i}x_{j},\sum_{i=n}^{1}\prod_{j=n}^{i}x_{j}}\right)$
+$\textsc{LOTZ}(x)=\left(\sum_{i=1}^{n}\prod_{j=1}^{i}x_{j},\sum_{i=1}^{n}\prod_{j=i}^{n}(1-x_{j})\right)$
 for any $x=(x_{1},\dots,x_{n})\in\{0,1\}^{n}$.
-7.2 Runtime
-We now establish the runtime of the SMS-EMOA on the OneMinMax and LOTZ.
-Note that the Pareto front sizes for OneMinMax and LOTZ are $n+1$.
-Hence, with the similar proof of Lemma 4, we know that for $\mu\geq n+1$ and for a specific function value $u$ reached by $F_{1}$, one individual in the non-empty $\cup_{i\geq 2}F_{i}$ or one individual $x$ in $F_{1}$ with $\Delta(x,F_{1})=0$ will be removed and $u$ will remain covered in the next generation.
-We have the following formal description.
-Lemma 23.
-Consider using SMS-EMOA with $\mu\geq n+1$ to optimize the LOTZ or OneMinMax problem.
-For any $x\in F_{1}$ for the current generation $t$, we have that there is $y$ with $f(y)=f(x)$ in $P_{t+1}$.
-Specifically, any Pareto front point will be maintained in all future populations once it is reached.
-Theorem 24 shows the runtime of SMS-EMOA on OneMinMax.
-The proof considers the probability to reach an uncovered Pareto front point from a specific covered one.
-From Lemma 23 that any covered Pareto front point will be kept, we then turn such probability to the runtime argument.
+Runtime
+We now analyze the runtime of the SMS-EMOA on the OneMinMax and LOTZ benchmarks.
+Note that the Pareto front, and more generally the size of any set of pairwise non-dominated solutions, for OneMinMax and LOTZ have sizes at most $n+1$.
+Hence we can use Lemma 4 with $M=n+1$.
+Theorem 20 below gives an upper bound for the runtime of SMS-EMOA on OneMinMax.
+This bound (in terms of fitness evaluations) is of the same asymptotic order as the runtime guarantee for the NSGA-II (Zheng, Liu, and Doerr 2022).
+Knowing from Lemma 4 that we cannot lose Pareto points, the proof of our results consists of adding the waiting times for finding a new Pareto point from an already existing neighboring one.
 Theorem 20.
-Consider using SMS-EMOA with $\mu\geq n+1$ to optimize the OneMinMax problem with problem size $n$. Then after at most $2e\mu n(\ln n+1)$ iterations in expectation, the full Pareto front is covered.
-Proof.
-Since each search point in $\{0,1\}^{n}$ is Pareto optimal, from Lemma 23, we know that any Pareto front point covered by the population $P_{t}$ will be maintained in all future generations.
-Assume that $(i,n-i)$ for some $i\in[0..n]$ is covered in $P_{t}$.
-Then the probability to reach a solution with function value $(i+1,n-i-1)$ is at least
-$\frac{1}{\mu}\frac{n-i}{n}\mathopen{}\mathclose{{}\left(1-\frac{1}{n}}\right)^{n-1}\geq\frac{n-i}{e\mu n},$
-and the probability to reach a solution with function value $(i-1,n-i+1)$ is at least
-$\frac{1}{\mu}\frac{i}{n}\mathopen{}\mathclose{{}\left(1-\frac{1}{n}}\right)^{n-1}\geq\frac{i}{e\mu n}.$
-Thus the expected numbers of iterations to cover $(i+1,n-i-1)$ and $(i-1,n-i+1)$ are at most $\frac{e\mu n}{n-i}$ and at most $\frac{e\mu n}{i}$ respectively.
-Hence, the expected number of iterations to cover the full Pareto front is
-$\Bigg{(}\sum_{j=i}^{1}\frac{e\mu n}{n-j}\Bigg{)}+\mathopen{}\mathclose{{}\left(\sum_{j=i}^{n-1}\frac{e\mu n}{j}}\right)\leq 2\sum_{j=1}^{n-1}\frac{e\mu n}{j}\leq 2e\mu n(\ln n+1).\qed$
-We now establish the runtime of SMS-EMOA on LOTZ, see Theorem 25.
+Consider using the SMS-EMOA with $\mu\geq n+1$ to optimize the OneMinMax problem with problem size $n$.
+Then after at most $2e\mu n(\ln n+1)$ iterations (or $\mu+2e\mu n(\ln n+1)$ fitness evaluations) in expectation, the full Pareto front is covered.
+Our runtime guarantee for LOTZ, see Theorem 21 below, again coincides with the guarantee for the NSGA-II (Zheng, Liu, and Doerr 2022).
 In the proof, the process is divided into two stages.
 The first stage is to cover a Pareto front point for the first time and the second stage is to cover the full Pareto front starting from one covered Pareto front point.
-For the first stage, we pessimistically consider the time to reach $1^{n}$ and for the second stage, we consider the time to reach $01^{i}0^{n-i},i=n-1,\dots,$ 0 one by one.
+For the first stage, we pessimistically consider the time to reach $1^{n}$.
+For the second stage, we consider the time to generate $01^{i}0^{n-i},i=n-1,\dots,$ 0 one after the other.
+In both stages, we heavily rely on Lemma 4 asserting that important progress is not lost, which allows to add waiting times of successive progress events.
 Theorem 21.
-Consider using SMS-EMOA with $\mu\geq n+1$ to optimize the LOTZ problem with problem size $n$. Then after at most $2e\mu n^{2}$ iterations in expectation, the full Pareto front is covered.
-Proof.
-We first consider the time to reach $1^{n}$ in the first stage.
-Let $S=\arg\max_{x\in P_{t}}f_{1}(x)$ and $S^{\prime}=\arg\max_{x\in R_{t}}f_{1}(x)$.
-Then there is at least one individual $x\in S^{\prime}$ such that $x\in F_{1}$ (If not, there is $y\notin S^{\prime}$, such that $x\prec y$ for all $x\in S^{\prime}$.
-That is, $f_{1}(y)\geq f_{1}(x),f_{2}(y)\geq f_{2}(x),$ and at least one of the inequality is strict.
-Since $y\notin S$, we know that $f_{1}(y)<f_{1}(x)$, which contradicts $f_{1}(y)\geq f_{1}(x)$).
-From Lemma 23, we know that $f(x)$ (also $f_{1}(x)$) will be covered by $P_{t+1}$.
-Hence, to reach $1^{n}$, we only need to reach such $S^{\prime}$ with $f_{1}(S^{\prime})=n$.
-We calculate
-$\Pr[f_{1}(S^{\prime})>f_{1}(S)\mid f_{1}(S)]\geq\frac{1}{\mu}\mathopen{}\mathclose{{}\left(1-\frac{1}{n}}\right)^{n-1}\frac{1}{n}\geq\frac{1}{e\mu n}$
-where we consider choosing one individual of $S$ and flipping the critical bit value from 00 to $1$.
-Since $f_{1}(S)$ will not decrease, we know that the expected iterations to witness such $S^{\prime}$ is at most $e\mu n$.
-Since there are at most $n$ such $S^{\prime}$ with $f_{1}(S^{\prime})>f_{1}(S)$ to reach $1^{n}$, we know that $1^{n}$ will be reached in at most $e\mu n^{2}$ in expectation.
-Once $1^{n}$ is reached, we consider the time to reach $1^{n-1}0,\dots,10^{n-1},0^{n}$ one by one.
-Assume that $1^{i}0^{n-i}$ with $i\in[1..n]$ is reached.
-Then to reach $1^{i-1}0^{n-i+1}$, we only need to choose $1^{i}0^{n-i}$ from $P_{t}$, and flip the $i$-th bit from $1$ to 00 with other bits unchanged, which happens with probability
-$\frac{1}{\mu}\mathopen{}\mathclose{{}\left(1-\frac{1}{n}}\right)^{n-1}\frac{1}{n}\geq\frac{1}{e\mu n}.$
-From Lemma 23, we know that any Pareto front point will be maintained in all future generations once reached.
-Hence, $1^{i}0^{n-i}\in P_{\tau}$ for all $\tau\geq t$.
-Thus, the expected number of iterations to reach $1^{i-1}0^{n-i+1}$ is at most $e\mu n$.
-Since we have $n$ other Pareto optima to reach, we know that the expected number of iterations to cover all remaining Pareto front points is at most $e\mu n^{2}$.
-In summary, the full Pareto front will be covered in at most $2e\mu n^{2}$ iterations in expectation. ∎
-Note that the runtime of the GSEMO and NSGA-II are $O(n^{2}\ln n)$ and $O(n^{3})$ for the OneMinMax and LOTZ respectively.
-Hence, Theorems 24 and 25 shows the SMS-EMOA achieves the same asymptotic runtime as GSEMO and NSGA-II.
-8 Conclusion
+Consider using the SMS-EMOA with $\mu\geq n+1$ to optimize the LOTZ problem with problem size $n$.
+Then after at most $2e\mu n^{2}$ iterations (or fitness evaluations) in expectation, the full Pareto front is covered.
+Note that the runtimes of the GSEMO on OneMinMax and LOTZ are $O(n^{2}\ln n)$ and $O(n^{3})$, respectively (Giel and Lehre 2010; Laumanns, Thiele, and Zitzler 2004).
+Hence, they agree with our bounds for the SMS-EMOA when using the most interesting population size $\mu=\Theta(n)$.
+Conclusion
 Motivated by the observed difficulty of the NSGA-II for many objectives, this paper resorted to the SMS-EMOA, a variant of the steady-state NSGA-II, and proved that, different from the NSGA-II, it efficiently solves the $m$OJZJ problem.
-Noting that the SMS-EMOA also employs non-dominated sorting, but replaces the crowding distance with the hypervolume, this result together with the one of Wietheger and Doerr [WD23] supports the conclusion that non-dominated sorting is a good building block for MOEAs, but the crowding distance has deficiencies for more than two objectives.
-We also showed that the stochastic population update proposed in [BZLQ23] for the bi-objective SMS-EMOA becomes less effective for many objectives.
+Noting that the SMS-EMOA also employs non-dominated sorting, but replaces the crowding distance with the hypervolume, this result together with the one of Wietheger and Doerr (2023) supports the conclusion that non-dominated sorting is a good building block for MOEAs, but the crowding distance has deficiencies for more than two objectives.
+We also showed that the stochastic population update proposed in (Bian et al.
+2023) for the bi-objective SMS-EMOA becomes less effective for many objectives.
 All these results in a very rigorous manner support the general knowledge that multi-objective optimization becomes increasingly harder with growing numbers of objectives.

### v3

@@
 To this aim, we first propose a many-objective counterpart, the $m$-objective $m$OJZJ problem, of the bi-objective OJZJ benchmark, which is the first many-objective multimodal benchmark used in a mathematical runtime analysis.
-We prove that SMS-EMOA computes the full Pareto front of this benchmark in an expected number of $O(M^{2}n^{k})$ iterations, where $n$ denotes the problem size (length of the bit-string representation), $k$ the gap size (a difficulty parameter of the problem), and $M=(2n/m-2k+3)^{m/2}$ the size of the Pareto front.
-This result together with the existing negative result on the original NSGA-II shows that in principle, the general approach of the NSGA-II is suitable for many-objective optimization, but the crowding distance as tie-breaker has deficiencies.
+We prove that SMS-EMOA computes the full Pareto front of this benchmark in an expected number of $O(\mu Mn^{k})$ iterations, where $n$ denotes the problem size (length of the bit-string representation), $k$ the gap size (a difficulty parameter of the problem), $M=(2n/m-2k+3)^{m/2}$ the size of the Pareto front, and $\mu$ the population size (at least the same size as the largest incomparable set).
+This result together with the existing negative result for the original NSGA-II shows that, in principle, the general approach of the NSGA-II is suitable for many-objective optimization, but the crowding distance as tie-breaker has deficiencies.
 We obtain three additional insights on the SMS-EMOA.
-Different from a recent result for the bi-objective OJZJ benchmark, the stochastic population update often does not help for $m$OJZJ.
-It results in a $1/\Theta(\min\{Mk^{1/2}/2^{k/2},1\})$ speed-up, which is $\Theta(1)$ for large $m$ such as $m>k$.
-On the positive side, we prove that heavy-tailed mutation still results in a speed-up of order $k^{0.5+k-\beta}$.
-Finally, we conduct the first runtime analyses of the SMS-EMOA on the bi-objective OneMinMax and LOTZ benchmarks and show that it has a performance comparable to the GSEMO and the NSGA-II.
-Introduction
+Different from a recent result for the bi-objective OJZJ benchmark, a recently proposed stochastic population update often does not help for $m$OJZJ.
+It at most results in a speed-up by a factor of order $2^{k}/\mu$, which is $\Theta(1)$ for large $m$, such as $m>k$.
+On the positive side, we prove that heavy-tailed mutation irrespective of the number $m$ of objectives results in a speed-up of order $k^{0.5+k-\beta}/e^{k}$, the same advantage as previously shown for the bi-objective case.
+Finally, we conduct the first runtime analyses of the SMS-EMOA on the classic bi-objective OneMinMax and LOTZ benchmarks and show that the SMS-EMOA has a performance comparable to the GSEMO and the NSGA-II.
+Our main technical insight, a general condition ensuring that the SMS-EMOA does not lose Pareto-optimal objective values, promises to be useful also in other runtime analyses of this algorithm.
+1 Introduction
 The NSGA-II is the most widely-applied multiobjective evolutionary algorithm (MOEA).
-Non-dominated sorting and crowding distance are its two major features differentiating it from basic MOEAs such as the GSEMO or the $(\mu+1)$ SIBEA.
-Zheng, Liu, and Doerr (2022) conducted the first runtime analysis of the NSGA-II (see (Zheng and Doerr 2023a) for the journal version).
-This work quickly inspired many interesting follow-up results in bi-objective optimization (Zheng and Doerr 2022; Bian and Qian 2022; Doerr and Qu 2023a, b, c; Dang et al.
-2023a, b; Cerf et al.
-2023).
-In contrast to these positive results for two objectives, Zheng and Doerr (2023b) proved that for $m\geq 3$ objectives the NSGA-II needs at least exponential time (in expectation and with high probability) to cover the full Pareto front of the $m$-objective OneMinMax benchmark, a simple many-objective version of the basic OneMax problem where all search points are Pareto optimal.
-They claimed that the main reason for this low efficiency is the independent computation of the crowding distance in each objective.
-A very recent work showed that the NSGA-III, a successor algorithm of the NSGA-II aimed at better coping with many objectives, can efficiently solve the $3$-objective OneMinMax problem (Wietheger and Doerr 2023).
-Since apparently practitioners much prefer the NSGA-II (more than 17,000 citations on Google scholar only in the last five years) over the NSGA-III (5,035 citations since its publication in 2013), it remains an interesting question whether there are variants of the NSGA-II which better cope with many objectives.
-With the SMS-EMOA, an interesting variant of the NSGA-II was proposed by Beume, Naujoks, and Emmerich (2007).
+Non-dominated sorting and crowding distance are its two main building blocks differentiating it from basic MOEAs such as the GSEMO or the $(\mu+1)$ SIBEA.
+Zheng, Liu, and Doerr [ZLD22] conducted the first runtime analysis of the NSGA-II (see [ZD23a] for the journal version).
+This work quickly inspired many interesting follow-up results in bi-objective optimization [ZD22, BQ22, DQ23a, DQ23b, DQ23c, DOSS23a, DOSS23b, CDH+23].
+In contrast to these positive results for two objectives, Zheng and Doerr [ZD23b] proved that for $m\geq 3$ objectives the NSGA-II needs at least exponential time (in expectation and with high probability) to cover the full Pareto front of the $m$-objective OneMinMax benchmark, a simple many-objective version of the basic OneMax problem.
+They claimed that the main reason for this low efficiency is the independent computation of the crowding distance in each objective, which suggested that the NSGA-II would have similar difficulties on other many-objective problems.
+A very recent work showed that the NSGA-III, a successor algorithm of the NSGA-II aimed at better coping with many objectives, can efficiently solve the $3$-objective OneMinMax problem [WD23].
+Since apparently practitioners much prefer the NSGA-II (more than 17,000 citations on Google scholar only in the last five years) over the NSGA-III (5,035 citations since its publication in 2013), it remains an interesting question whether there are other variants of the NSGA-II which also cope well with many objectives.
+With the SMS-EMOA, an interesting variant of the NSGA-II was proposed by Beume, Naujoks, and Emmerich [BNE07].
 This algorithm is a steady-state variant of the NSGA-II (that is, in each iteration only a single offspring is generated and possibly integrated into the population) that further replaces the crowding distance as secondary selection criterion with the classic hypervolume contribution.
-Many empirical works (see the almost 2,000 papers citing (Beume, Naujoks, and Emmerich 2007)) confirmed the good performance of the SMS-EMOA for many-objective optimization.
+Many empirical works (see the almost 2,000 papers citing [BNE07]) confirmed the good performance of the SMS-EMOA for many-objective optimization.
 The first mathematical runtime analysis of the SMS-EMOA was conducted very recently by Bian et al.
-(2023), who proved that its expected runtime on the bi-objective OJZJ problem is $O(n^{k+1})$.
+[BZLQ23], who proved that its expected runtime on the bi-objective OJZJ problem is $O(n^{k+1})$.
 They also proposed a stochastic population update mechanism and proved that it has the often superior runtime of $O(n^{k+1}\min\{1,n/2^{k/4}\})$.
 Zheng et al.
-(2024) proved that the SMS-EMOA has an expected runtime of $O(n^{4})$ on the bi-objective DLTB problem.
+[ZLDD24] proved that the SMS-EMOA has an expected runtime of $O(n^{4})$ on the bi-objective DLTB problem.
 Our Contributions: This paper conducts the first mathematical runtime analysis of the SMS-EMOA for more than two objectives.
-We first define the $m$OJZJ benchmark, an $m$-objective counterpart of the bi-objective OJZJ problem (Zheng and Doerr 2023d), which is the problem analyzed in the first runtime analysis for the SMS-EMOA (Bian et al.
-2023).
+We first define the $m$OJZJ benchmark, an $m$-objective counterpart of the bi-objective OJZJ problem [ZD23c], which is the problem analyzed in the first runtime analysis for the SMS-EMOA [BZLQ23].
 We note that the $m$OJZJ problem is the first multimodal many-objective benchmark proposed for a theoretical analysis, to the best of our knowledge.
-We prove that the SMS-EMOA covers the full Pareto front of this benchmark in an expected number of $O(M^{2}n^{k})$ iterations, where $n$ is the problems size, $k$ the gap size (a difficulty parameter of the problem), $m$ the number of objectives, and $M=(2n/m-2k+3)^{m/2}$ the size of the Pareto front.
+We prove in Theorem 9 that the SMS-EMOA can compute the full Pareto front of this benchmark in an expected number of $O(\mu Mn^{k})$ iterations, where $n$ is the problems size, $k$ the gap size (a difficulty parameter of the problem that is usually small), $m$ the number of objectives, $M=(2n/m-2k+3)^{m/2}$ the size of the Pareto front, and where the population size $\mu$ has to be at least as large as the largest set of incomparable solutions (which is not larger than $(2n/m+1)^{m/2}$).
 We recall that the original NSGA-II needs at least exponential time to optimize the $m$OneMinMax problem, which is a special case of $m$OJZJ with gap size $k=1$.
-Since the SMS-EMOA employs non-dominated sorting, but replaces the crowding distance in the original NSGA-II by the hypervolume contribution, our result in a similar fashion as the analysis of the NSGA-III in (Wietheger and Doerr 2023) suggests that the general approach of the NSGA-II is suitable for many-objective optimization and that it is only the crowding distance as tie-breaker which is not appropriate for more than two objectives.
-We then analyze whether the better performance of the SMS-EMOA on the bi-objective OJZJ problem achieved via a new stochastic population update (Bian et al.
-2023) extends to the $m$-objective $m$OJZJ problem.
-Unfortunately, we shall observe that only a speed-up of order $1/\Theta(\min\{Mk^{1/2}/2^{k/2},1\})$ is obtained, which is $\Theta(1)$ when $m$ is large, e.g., $m>k$.
+Since the SMS-EMOA employs non-dominated sorting, but replaces the crowding distance in the original NSGA-II by the hypervolume contribution, our result in a similar fashion as the analysis of the NSGA-III in [WD23] suggests that the general approach of the NSGA-II is suitable for many-objective optimization and that it is only the crowding distance as tie-breaker which is not appropriate for more than two objectives.
+We then analyze whether the better performance of the SMS-EMOA on the bi-objective OJZJ problem achieved via a new stochastic population update [BZLQ23] extends to the $m$-objective $m$OJZJ problem.
+Unfortunately, we shall observe that only a speed-up of order at most $2^{k}/\mu$ is obtained, which is $\Theta(1)$ when $m$ is large, e.g., $m>k$.
 On the positive side, we show that the advantage of heavy-tailed mutation is preserved.
-We analyze the SMS-EMOA with heavy-tailed mutation on the $m$OJZJ benchmark and prove that a speed-up of order $k^{0.5+k-\beta}$ is achieved.
-This is the same speed-up as observed for single-objective and bi-objective Jump problems (Doerr et al.
-2017; Zheng and Doerr 2023d).
+We analyze the SMS-EMOA with heavy-tailed mutation on the $m$OJZJ benchmark and prove that a speed-up of order $k^{0.5+k-\beta}/e^{k}$ is achieved.
+This is the same speed-up as observed for single-objective and bi-objective Jump problems [DLMN17, ZD23c].
 We note that this is the first theoretical work to support the usefulness of heavy-tailed mutation in many-objective optimization.
-Finally, since so far the performance of the SMS-EMOA was only analyzed on the bi-objective OJZJ problem (Bian et al.
-2023), we conduct mathematical runtime analyses of the SMS-EMOA also on the two most prominent bi-objective benchmarks OneMinMax and LOTZ.
-We prove that the SMS-EMOA finds the Pareto fronts of these benchmarks in an expected number of iterations of at most $2e(n+1)n(\ln n+1)$ for OneMinMax and at most $2en^{2}(n+1)$ for LOTZ.
+Finally, since so far the performance of the SMS-EMOA was only analyzed on the bi-objective OJZJ problem [BZLQ23] (and, in a work [ZLDD24] parallel to this, on the not overly prominent DLTB benchmark), we conduct mathematical runtime analyses of the SMS-EMOA also on the two most prominent bi-objective benchmarks OneMinMax and LOTZ.
+We prove that the SMS-EMOA finds the Pareto fronts of these benchmarks in an expected number of at most $2e(n+1)n(\ln n+1)$ iterations for OneMinMax and at most $2en^{2}(n+1)$ for LOTZ.
 These are the same asymptotic runtimes (in terms of fitness evaluations) as known for the GSEMO and the NSGA-II.
-Preliminaries
-In many-objective optimization, one tries to find good solutions for a problem containing several, but at least three objectives.
+The technical key to all these results are two simple, but general results ensuring that the SMS-EMOA (different from the NSGA-II in many-objective optimization) cannot forget Pareto-optimal solution values when the population size is sufficiently large (Lemmas 4 and 20).
+We are convinced that these will find applications in future runtime analyses of the SMS-EMOA.
+In fact, Lemma 4 has already been used in the subsequent work [WD24].
+This paper is organized as follows.
+Section 2 briefly introduces some basic concepts for the multi-objective optimization and the mathematical notations that will be used in this paper.
+The many-objective OJZJ benchmark is proposed in Section 3.
+Sections 4 to 6 analyze the runtime of the original SMS-EMOA, the SMS-EMOA with stochastic population update, and the one with heavy-tailed mutation on this many-objective benchmark.
+The runtime for two well-known bi-objective functions is conducted in Section 7.
+Section 8 concludes this work.
+2 Preliminaries
+In multi-objective optimization, one tries to find good solutions for a problem containing several objectives.
 This work considers pseudo-Boolean maximization, hence our problem is described by a function $f=(f_{1},\dots,f_{m}):\{0,1\}^{n}\rightarrow\mathbb{R}^{m}$.
-Here $m>2$ is the number of objectives and $n\in\mathbb{N}$ is called the problem size.
-As in bi-objective optimization (that is, the case $m=2$), usually not all solutions are comparable.
-We say that a solution $x\in\{0,1\}^{n}$ dominates a solution $y$, denoted as $x\succ y$, if $f_{i}(x)\geq f_{i}(y)$ for all $i\in\{1,\dots,m\}$, and at least one of these inequalities is strict.
-For a given many-objective function $f$, we say that $x\in\{0,1\}^{n}$ is a Pareto optimum if and only if there is no $y\in\{0,1\}^{n}$ that dominates $x$ w.r.t.
-$f$.
-The set of all Pareto optima is called Pareto set, and the set of all function values of the Pareto optima is called Pareto front.
-The goal for an MOEA is to find the full Pareto front as much as possible, that is, to compute a not too large set $S$ of solutions such that $f(S)$ equals or approximates well the Pareto front.
+Here $m$ is the number of objectives and $n\in\mathbb{N}$ is called the problem size.
+We use the term bi-objective optimization when $m=2$, and many-objective optimization when $m\geq 3$.
+The literature is not uniform in from which number of objectives to speak of many objectives.
+We include already $m=3$ into the many-objective setting since the drastically different behavior of the NSGA-II detected in [ZD23b] starts already at $m=3$.
+Different from single-objective optimization, usually not all solutions are comparable in multi-objective problems.
+We say that a solution $x\in\{0,1\}^{n}$ dominates a solution $y$ (with regard to $f$), denoted as $x\succ y$, if $f_{i}(x)\geq f_{i}(y)$ for all $i\in\{1,\dots,m\}$, and at least one of these inequalities is strict.
+For a given multi-objective function $f$, we say that $x\in\{0,1\}^{n}$ is a Pareto optimum if and only if there is no $y\in\{0,1\}^{n}$ that dominates $x$ (with regard to $f$).
+The set of all Pareto optima is called Pareto set, and the set of all function values of Pareto optima is called Pareto front.
+The goal for an MOEA is to find the full Pareto front as good as possible, that is, to compute a not too large set $S$ of solutions such that $f(S)$ equals or approximates well the Pareto front.
 As common in the mathematical runtime analysis of MOEAs, we call the runtime of an algorithm the number of function evaluations until its population $P$ covers the full Pareto front, that is, $f(P)$ contains the Pareto front.
-We refer to (Neumann and Witt 2010; Auger and Doerr 2011; Jansen 2013; Zhou, Yu, and Qian 2019; Doerr and Neumann 2020) for general introductions to the mathematical runtime analysis of evolutionary algorithms, and to (Brockhoff 2011) for a discussion of runtime analyses of MOEAs.
-Mathematical Notations.
+We refer to [NW10, AD11, Jan13, ZYQ19, DN20] for general introductions to the mathematical runtime analysis of evolutionary algorithms, and to [Bro11] for a discussion of runtime analyses of MOEAs.
+Mathematical notation.
 For $a,b\in\mathbb{N}$ with $a\leq b$, we denote the set of integers in the interval $[a,b]$ by $[a..b]=\{a,a+1,\dots,b\}$.
 For a point $x\in\{0,1\}^{n}$, let $x_{[a..b]}:=(x_{a},x_{a+1},\dots,x_{b})$ and $|x|_{1}$ be the number of ones in $x$.
-The $m$-Objective Jump Function
-As mentioned earlier, the NSGA-II was proven to have enormous difficulties in optimizing many-objective problems (Zheng and Doerr 2023b).
-In that paper, the $m$-objective counterpart $m$OneMinMax of the bi-objective OneMinMax benchmark was proposed and analyzed.
-Since the first runtime work of the SMS-EMOA so far (Bian et al.
-2023) analyzed the bi-objective jump problem OJZJ, whose special case with gap size $k=1$ is (essentially) the OneMinMax problem, we shall propose and work with an $m$-objective version of OJZJ as well.
+3 The $m$-Objective Jump Problem
+As mentioned earlier, the NSGA-II was proven to have enormous difficulties in optimizing many-objective problems [ZD23b].
+In that paper, the $m$-objective counterpart $m$OneMinMax of the bi-objective OneMinMax benchmark [GL10] was proposed and analyzed.
+Since the first runtime work of the SMS-EMOA so far [BZLQ23] analyzed the bi-objective jump problem OJZJ, whose special case with gap size $k=1$ is (essentially) the OneMinMax problem, we shall propose and work with an $m$-objective version of OJZJ.
 Again, its special case $k=1$ will be (essentially) equal to the $m$OneMinMax problem.
-With this, our results are comparable both the ones in (Zheng and Doerr 2023b) and (Bian et al.
-2023).
-$m$OJZJ
+With this, our results are comparable both the ones in [ZD23b] and [BZLQ23].
+3.1 $m$OJZJ
 We first recall the definition of the $m$OneMinMax problem.
@@
 On each of these, a bi-objective OneMinMax problem is defined.
-We note that this general block construction goes back to the seminal paper of (Laumanns, Thiele, and Zitzler 2004).
-Definition 1 (Zheng and Doerr (2023b)).
-Let $m$ be the number of objectives and be even, and the problem size $n$ be a multiple of $m/2$.
+We note that this general approach to construct many-objective problems from bi-objective ones goes back to the seminal paper of [LTZ04].
+Definition 1 ([ZD23b]).
+Let $m$ be the even number of objectives.
+Let the problem size $n$ be a multiple of $m/2$.
 Let $n^{\prime}=2n/m\in\mathbb{N}$.
-For any $x=(x_{1},\dots,x_{n})$, the $m$-objective function mOneMinMax $f:\{0,1\}^{n}\rightarrow\mathbb{R}^{m}$ is defined by
-$f_{i}(x)$ =$=\begin{cases}\textsc{OneMax}(\bar{x}_{[\frac{i-1}{2}n^{\prime}+1..\frac{i+1}{2}n^{\prime}]}),&\text{if $i$ is odd}\\ \textsc{OneMax}({x}_{[\frac{i-2}{2}n^{\prime}+1..\frac{i}{2}n^{\prime}]}),&\text{else,}\end{cases}$
+For any $x=(x_{1},\dots,x_{n})$, the $m$-objective function mOneMinMax is the function $f:\{0,1\}^{n}\rightarrow\mathbb{R}^{m}$ defined by
+$f_{i}(x)$ =$=\begin{cases}\textsc{OneMax}(\bar{x}_{[\frac{i-1}{2}n^{\prime}+1..\frac{i+1}{2}n^{\prime}]}),&\text{if $i$ is odd,}\\ \textsc{OneMax}({x}_{[\frac{i-2}{2}n^{\prime}+1..\frac{i}{2}n^{\prime}]}),&\text{else,}\end{cases}$
 where $\bar{x}=(1-x_{1},\dots,1-x_{n})$ and the function $\textsc{OneMax}:\{0,1\}^{n^{\prime}}\rightarrow\mathbb{R}$ is defined by
@@
 for any $y\in\{0,1\}^{n^{\prime}}$.
-In order to include $m$OneMinMax as a special case of $m$OJZJ, we define the $m$-objective OJZJ in a similar manner, that is, we divide the $n$ bit positions into $m/2$ blocks and define a OJZJ problem in each block.
+We define the $m$-objective OJZJ in a similar manner, that is, we divide the $n$ bit positions into $m/2$ blocks and define a OJZJ problem in each block.
 Definition 2.
 […]
-$f_{i}(x)$ =$=\begin{cases}\textsc{Jump}_{n^{\prime},k}(x_{[\frac{i-1}{2}n^{\prime}+1..\frac{i+1}{2}n^{\prime}]}),&\text{if $i$ is odd}\\ \textsc{Jump}_{n^{\prime},k}(\bar{x}_{[\frac{i-2}{2}n^{\prime}+1..\frac{i}{2}n^{\prime}]}),&\text{else,}\end{cases}$
+$f_{i}(x)$ =$=\begin{cases}\textsc{Jump}_{n^{\prime},k}(x_{[\frac{i-1}{2}n^{\prime}+1..\frac{i+1}{2}n^{\prime}]}),&\text{if $i$ is odd,}\\ \textsc{Jump}_{n^{\prime},k}(\bar{x}_{[\frac{i-2}{2}n^{\prime}+1..\frac{i}{2}n^{\prime}]}),&\text{else,}\end{cases}$
 where $\bar{x}=(1-x_{1},\dots,1-x_{n})$ and the function $\textsc{Jump}_{n^{\prime},k}:\{0,1\}^{n^{\prime}}\rightarrow\mathbb{R}$ is defined by
-[…]
+$\textsc{Jump}_{n^{\prime},k}(y)$ =$=\begin{cases}k+\textsc{OneMax}(y),&\text{if $|y|_{1}\leq n^{\prime}-k$ or $y=1^{n^{\prime}}$}\\ n^{\prime}-\textsc{OneMax}(y),&\text{else}\end{cases}$
 for any $y\in\{0,1\}^{n^{\prime}}$.
 We note that the function $\textsc{Jump}_{n^{\prime},k}$ used in the definition above is the famous Jump benchmark.
-It was first defined by Droste, Jansen, and Wegener (2002) and has quickly become the most employed multimodal benchmark in the theory of randomized search heuristics, leading to many fundamental results on how these algorithms cope with local optima (Jansen and Wegener 2002; Dang et al.
-2018; Doerr 2021; Benbaki, Benomar, and Doerr 2021; Hevia Fajardo and Sudholt 2022; Rajabi and Witt 2022; Witt 2023; Doerr et al.
-2024).
+It was first defined by Droste, Jansen, and Wegener [DJW02] and has quickly become the most employed multimodal benchmark in the theory of randomized search heuristics, leading to many fundamental results on how these algorithms cope with local optima [JW02, DFK+18, Doe21, BBD21, HS22, RW22, Wit23, DEJK24].
 […]
-Characteristics
+3.2 Characteristics
 We now give more details on this $m$OJZJ function.
-Let $B_{i}:=[(i-1)n^{\prime}+1..in^{\prime}]$, $i\in[1..m/2]$, be the $i$-th block in the partition of $n$ bit positions.
-From Definition 2, we know that the bit values in the block $B_{i}$ only influence the objectives $f_{2i-1}$ and $f_{2i}$.
+Let $B_{i}:=[(i-1)n^{\prime}+1..in^{\prime}]$, $i\in[1..m/2]$, be the $i$-th block of the $n$ bit positions.
+From Definition 2, we know that the bit values in block $B_{i}$ only influence the objectives $f_{2i-1}$ and $f_{2i}$.
 Figure 1 plots the objective values of $f_{2i-1}$ and $f_{2i}$ relative to the number of ones in this block.
-Obviously, $m$OJZJ is multimodal with respect to the definition of multimodality of multiobjective problems in (Zheng and Doerr 2023d).
+Obviously, $m$OJZJ is multimodal with respect to the definition of multimodality of multiobjective problems in [ZD23c].
 Figure 1: The objective values of $f_{2i-1}$ and $f_{2i}$ in $m$OJZJ w.r.t. $|x_{B_{i}}|_{1}$, the number of ones in the block $B_{i}$.
-It is not difficult to see that the Pareto set is $S^{*}=\{x\in\{0,1\}^{n}\mid|x_{B_{i}}|_{1}\in[k..n-k]\cup\{0,n\},i\in[1..m/2]\}$. Hence, the Pareto front is
+It is not difficult to see that the Pareto set is
+$S^{*}=\{x\in\{0,1\}^{n}\mid\forall i\in[1..m/2],|x_{B_{i}}|_{1}\in[k..n-k]\cup\{0,n\}\}.$
+Hence, the Pareto front is
 […]
@@
 and the Pareto front size is $M:=|F^{*}|=(n^{\prime}-2k+3)^{m/2}$.
-We recall at this point that we defined the runtime of an MOEA as the time to cover the full Pareto front $F^{*}$, that is, the first time to reach a population $P$ with $F^{*}\subseteq f(P)$.
-Note that for $k>1$, not all points are Pareto optimal and not all sets of mutually non-dominated points are the subsets of the Pareto front.
-However, we have the following result showing that the maximum number of mutually non-dominated points is at most the size of the Pareto front.
-Due to the limited space, all mathematical proofs could only be sketched or had to be omitted completely.
-They can be found in the preprint (Zheng and Doerr 2023c).
+What is harder to determine is the size of a largest set of incomparable, that is, pairwise not weakly dominating, individuals.
+One could be tempted to believe that this number is equal to the size of the Pareto front, but this is not true.
+[^1] In fact, in the conference version [ZD24] of this work we made this wrong claim.
+We sincerely apologize to the research community for this mistake.
+Since this number is important in the analysis of MOEAs, for example, it is the best upper bound on the population size of the GSEMO algorithm, we now discuss this question in more detail.
+To this aim, let us define $\overline{M}$ as the maximum size of a set of incomparable individuals (with respect to the $m$OJZJ problem for some parameters $n$, $k$, and $m$).
+By definition, $\overline{M}\geq M$, since the Pareto front is witnessed by a set of incomparable solutions.
+The following example shows that $\overline{M}$ can be considerably larger than $M$.
+Let $n$ be large, $m=4$, and $k=n^{\prime}/2$.
+Then $F^{*}=\{(a_{1},n^{\prime}+2k-a_{1},a_{2},n^{\prime}+2k-a_{2})\mid a_{1},a_{2}\in\{n^{\prime}/2,n^{\prime},3n^{\prime}/2\}\}$ and hence $|F^{*}|=9$.
+We now construct a set of pairwise incomparable solutions with much larger cardinality.
+For $i=1,\dots,k-1$ consider the individual $x^{(i)}=1^{i}0^{n^{\prime}-i}1^{n^{\prime}/2+i}0^{n^{\prime}/2-i}$, where we use the word notation known from theoretical computer science to describe bit-strings.
+In this notation, for example, $1^{i}$ is a string consisting of $i$ ones, and the concatenation of two strings $u$ and $v$ is denoted by $uv$.
+We have $f(x^{(i)})=(k+i,i,n^{\prime}/2-i,k+n^{\prime}/2-i)$.
+Consequently, for $i,j\in[1..k-1]$ with $i\neq j$, the two individuals $x^{(i)}$ and $x^{(j)}$ are incomparable.
+This shows that $\overline{M}\geq k-1=n/4-1$, which can be considerably larger than $M=9$.
+Clearly, this example is on the extreme side.
+Also, noting that all our runtimes will be at least of order $n^{k}$, such large values of $k$ lead to instance that are too difficult to be solved anyway.
+For the more typical case that $k$ is small, the following trivial upper bound is relatively tight, as wittnessed by its closeness to the lower bound $M$.
 Lemma 3.
-Let $k\leq n^{\prime}/2$ and $P$ be a set of solutions such that any two are not weakly dominated (w.r.t. $m$OJZJ). Then $|P|\leq M$.
-For ease of reading of the proof sketches in the following sections, we call a Pareto optimum $x$ an inner Pareto optimum if for all blocks $B_{i}=[(i-1)n^{\prime}+1..in^{\prime}],i\in[1..m/2]$, we have $x_{B_{i}}\notin\{0^{n^{\prime}},1^{n^{\prime}}\}$.
-We call $u=(u_{1},n^{\prime}+2k-u_{1},\dots,u_{m/2},n^{\prime}+2k-u_{m/2})$ and $v=(v_{1},n^{\prime}+2k-v_{1},\dots,v_{m/2},n^{\prime}+2k-v_{m/2})$ neighbors if there exists $i\in[1..m/2]$ such that $u_{i}-v_{i}=1$ and $u_{j}=v_{j}$ for all $j\neq i$.
-The SMS-EMOA Can Optimize $m$OJZJ
-As discussed before, the original NSGA-II with non-dominated sorting and crowding distance needs at least exponential runtime to cover the full Pareto front of $m$OneMinMax, $m\geq 3$, (Zheng and Doerr 2023b), which is a special case of $m$OJZJ.
-In this section, we analyze the runtime of the SMS-EMOA on this problem and show that it does not encounter such problems.
-Algorithm Description
-The SMS-EMOA is a steady-state variant (that is, the offspring size is smaller than the parent size) of the NSGA-II.
-Like the NSGA-II, the SMS-EMOA works with a population of fixed size $\mu$.
-However, in each iteration, only one offspring $x^{\prime}$ is generated from the parent population $P_{t}$.
+Let $k\leq n^{\prime}/2$ and $P$ be a set of solutions pairwise incomparable solutions (w.r.t. $m$OJZJ). Then $|P|\leq\overline{M}\leq(n^{\prime}+1)^{m/2}$.
+Proof.
+Let $P^{\prime}$ be a set of pairwise incomparable solutions of maximal cardinality.
+By definition, $P^{\prime}$ contains no two solutions with identical objective value.
+Consequently, $|P|\leq\overline{M}=|P^{\prime}|\leq f(\{0,1\}^{n})=(n^{\prime}+1)^{m/2}$.
+∎
+For the ease of reading of the proofs in the following sections, we call a Pareto optimum $x$ an inner Pareto optimum if for all blocks $B_{i},i\in[1..m/2]$, we have $x_{B_{i}}\notin\{0^{n^{\prime}},1^{n^{\prime}}\}$.
+This is equivalent to saying that $|x_{B_{i}}|_{1}\in[k..n^{\prime}-k]$ for all $i\in[1..m/2]$.
+We call two objective vectors $u=(u_{1},n^{\prime}+2k-u_{1},\dots,u_{m/2},n^{\prime}+2k-u_{m/2})$ and $v=(v_{1},n^{\prime}+2k-v_{1},\dots,v_{m/2},n^{\prime}+2k-v_{m/2})$ neighbors if there exists an $i\in[1..m/2]$ such that $|u_{i}-v_{i}|=1$ and $u_{j}=v_{j}$ for all $j\neq i$.
+4 The SMS-EMOA Can Optimize $m$OJZJ
+As discussed before, the original NSGA-II with non-dominated sorting and crowding distance needs at least exponential time to cover the full Pareto front of $m$OneMinMax, $m\geq 3$, [ZD23b], which is a special case of $m$OJZJ.
+In this section, we analyze the runtime of the SMS-EMOA on this problem, and more generally, the $m$OJZJ problem for $k\in[1..n^{\prime}/2]$, and show that it does not encounter such problems.
+4.1 Algorithm Description
+The SMS-EMOA is a steady-state variant (that is, the offspring population size is much smaller than the parent population size) of the NSGA-II.
+Like the NSGA-II, the SMS-EMOA works with a (parent) population of fixed size $\mu$.
+However, in each iteration, only one offspring $x^{\prime}$ is generated from the population $P_{t}$.
 From the $\mu+1$ individuals in the combined parent and offspring population $R_{t}=P_{t}\cup\{x^{\prime}\}$, a single individual is removed.
 To this aim, the SMS-EMOA like the NSGA-II uses non-dominated sorting, that is, it partitions $R_{t}$ into fronts $F_{1},\dots,F_{i^{*}}$, where $F_{i}$ contains all non-dominated individuals in $R_{t}\setminus(\bigcup_{j=1}^{i-1}F_{j})$.
-Different from the original NSGA-II that now uses the crowding distance as the secondary criterion for the removal of individuals, the SMS-EMOA removes the individual with the smallest hypervolume contribution in the critical front (which here always is the last front $F_{i^{*}}$).
+Different from the original NSGA-II, which uses the crowding distance as the secondary criterion for the removal of individuals, the SMS-EMOA removes the individual with the smallest hypervolume contribution in the critical front (which here always is the last front $F_{i^{*}}$).
 The hypervolume of a set $S$ of individuals w.r.t. a reference point $r$ in the objective space is defined as
-$\operatorname{HV}_{r}(S)=\mathcal{L}\left(\,\bigcup_{u\in S}\{h\in\mathbb{R}^{m}\mid r\leq h\leq f(u)\}\right),$
+$\operatorname{HV}_{r}(S)=\mathcal{L}\mathopen{}\mathclose{{}\left(\,\bigcup_{u\in S}\{h\in\mathbb{R}^{m}\mid r\leq h\leq f(u)\}}\right),$
 where $\mathcal{L}$ is the Lebesgue measure. The hypervolume contribution of an individual $x\in F_{i^{*}}$ is calculated via
 $\Delta_{r}(x,F_{i^{*}}):=\operatorname{HV}_{r}(F_{i^{*}})-\operatorname{HV}_{r}(F_{i^{*}}\setminus\{x\}).$
-Algorithm 1 gives the pseudocode of the SMS-EMOA.
+Algorithm 1 gives the pseudocode of the SMS-EMOA regarded in this work.
+We generate the offspring by mutating a single, randomly chosen parent via bit-wise mutation with mutation rate $\frac{1}{n}$.
+We note that all our proofs are robust to adding crossover with constant rate $p<1$, that is, if we could generate the offspring via some crossover operator with probability $p$ and via mutation otherwise, all our runtime guarantees would remain valid except for an additional $\frac{1}{1-p}$ factor.
 1:Initialize $P_{0}$ by generating $\mu$ solutions uniformly at random from $\{0,1\}^{n}$
@@
 4:   Generate $x^{\prime}$ by flipping each bit of $x$ independently with probability $1/n$
-5:   Use fast-non-dominated-sort() (Deb et al. 2002) to partition $R_{t}=P_{t}\cup\{x^{\prime}\}$ into $F_{1},\dots,F_{i^{*}}$
+5:   Use fast-non-dominated-sort() [DPAM02] to partition $R_{t}=P_{t}\cup\{x^{\prime}\}$ into $F_{1},\dots,F_{i^{*}}$
 6:   Calculate $\Delta_{r}(z,F_{i^{*}})$ for all $z\in F_{i^{*}}$ and find $D=\arg\min_{z\in F_{i^{*}}}{\Delta_{r}(z,F_{i^{*}})}$
@@
 Algorithm 1 SMS-EMOA
-Runtime of SMS-EMOA
+4.2 Runtime of SMS-EMOA
 We now analyze the runtime of the SMS-EMOA, that is, the time until its population covers the full Pareto front of $m$OJZJ.
-We start by proving the following result ensuring that certain individuals survive in the population or are replaced by at least as good individuals.
-We formulate this result in manner more general than what we need since we expect this general version to be useful in other analyses of the SMS-EMOA.
-The key to the proof is that (i) if there is a point with a rank larger than one, then all solutions with rank one (that is, in $F_{1}$) will survive, and (ii) if $F_{1}$ contains two or more solutions with same function value, then only such a solution can be removed from $F_{1}$.
-If the population size is large enough, by the pigeon-hole principle, one of these two cases comes true.
+We start by proving the following general result ensuring that if the population size is at least the maximum size of a set of incomparable solutions, then individuals survive in the population or are replaced by at least as good individuals.
+We formulate this result in manner more general than what we need since we expect this version to be useful in future analyses of the SMS-EMOA (and this happened already in [WD24]).
+The key to the proof is first noting that by a simple domination argument we can concentrate on individuals in $F_{1}$.
+For these, we argue that (i) if the combined parent and offspring population contains an individual not in $F_{1}$, then all solutions in $F_{1}$ will survive, and (ii) if $F_{1}$ contains two or more solutions with same function value, then only such a solution can be removed from $F_{1}$.
+If the population size is large enough, by the pigeon-hole principle, one of these two cases is always satisfied.
 Lemma 4.
 Consider any $m$-objective optimization problem.
-Let $M\in\mathbb{N}$ be such that any set $S$ of pairwise non-dominated solutions having different objective values satisfies $|S|\leq M$.
-Consider solving this problem via the SMS-EMOA with population size $\mu\geq M$ and using a reference point $r$ such that $\operatorname{HV}_{r}(\{x\})>0$ for any individual $x\in\{0,1\}^{n}$.
+Let $\overline{M}\in\mathbb{N}$ be such that any set $S$ of incomparable solutions satisfies $|S|\leq\overline{M}$.
+Consider solving this problem via the SMS-EMOA with population size $\mu\geq\overline{M}$ and using a reference point $r$ such that $\operatorname{HV}_{r}(\{x\})>0$ for any individual $x\in\{0,1\}^{n}$.
 Then the following is true.
@@
 In particular, if $R_{t}$ contains a Pareto optimum $x$, then all future generations contain a solution $y$ with $f(y)=f(x)$.
-Recall that an inner Pareto optimum is a Pareto optimum with $x_{B_{i}}\notin\{0^{n^{\prime}},1^{n^{\prime}}\}$ for all blocks $B_{i}=[(i-1)n^{\prime}+1..in^{\prime}],i\in[1..m/2]$, as defined before.
-While it is very likely that at least one initial solution is an inner Pareto optimum, this does not happen with probability one, and hence in the following lemma we estimate the time until the population contains an inner Pareto optimum.
+Proof.
+Let $t\in\mathbb{N}$ and $x\in R_{t}$.
+By induction, it suffices to show the lemma for $s=t+1$, that is, we now show that $P_{t+1}$ contains an individual $y$ such that $y\succeq x$.
+Without loss of generality, we may assume that $x$ is contained in the first front $F_{1}$ of the non-dominated sorting of $R_{t}$ – if not, then there is some $\tilde{x}\in F_{1}$ with $\tilde{x}\succ x$ and showing the existence of a $y\in P_{t+1}$ with $y\succeq\tilde{x}$ immediately gives the desired $y$ with $y\succeq x$.
+Hence let us assume that $x\in F_{1}$.
+If $|F_{1}|\leq\mu$, then all individuals of $F_{1}$ are taken into the next generation.
+Consequently $y:=x\in P_{t+1}$ is the desired individual.
+If $|F_{1}|=\mu+1$, then by our assumption $\mu\geq\overline{M}$ and the fact that $F_{1}$ is a set of incomparable solutions, we see that $F_{1}$ contains at least two individuals with identical objective value (pigeon-hole principle).
+Since such individuals have a hypervolume contribution of zero, and since only such individuals have a hypervolume contribution of zero, we know that the single individual $z^{\prime}$ removed from $R_{t}$ is an individual such that there is at least one more individual $z^{\prime\prime}\in R_{t}$ with $f(z^{\prime})=f(z^{\prime\prime})$.
+If $z^{\prime}=x$, then $y:=z^{\prime\prime}\in P_{t+1}$ is the desired individual.
+If $z^{\prime}\neq x$, then $x$ survives into $P_{t+1}$ and thus trivially $y:=x$ is the desired individual.
+∎
+Recall that an inner Pareto optimum is a Pareto optimum with $x_{B_{i}}\notin\{0^{n^{\prime}},1^{n^{\prime}}\}$ for all blocks $B_{i},i\in[1..m/2]$, as defined before.
+While it is very likely that at least one initial solution is an inner Pareto optimum, this does not happen with probability one, and hence in Lemma 6 we estimate the time until the population contains an inner Pareto optimum.
 Since this time usually is much smaller than the time to generate all remaining Pareto optima, we do not care that this is estimate could easily be improved.
@@
 By Lemma 4, an individual with this $f_{i}$ value will remain in the population.
-Hence a total of at most $m$ such block changes (applied to the right individual) suffice to obtain an inner Pareto optimum.
+Hence a total of at most $m/2$ such block changes (applied to the right individual) suffice to obtain an inner Pareto optimum.
+In the proof, we need the following elementary observation, which extends an analogous result from [BZLQ23] for the special case that $n^{\prime}=n$.
+Lemma 5.
+Let $k,n^{\prime},n\in\mathbb{N}$ such that $k\leq n^{\prime}/2$ and $n^{\prime}\leq n$. Then $\binom{n^{\prime}-i}{k-i}/n^{k-i}$ increases in $i\in[0..k-1]$.
+Proof.
+Recall that in the proof of [BZLQ23, Theorem 1], it was already proven that $\binom{n-i}{k-i}/n^{k-i}$ increases in $i$. Hence
+$\frac{\binom{n^{\prime}-i}{k-i}}{n^{k-i}}=\frac{\binom{n^{\prime}-i}{k-i}}{(n^{\prime})^{k-i}}\mathopen{}\mathclose{{}\left(\frac{n^{\prime}}{n}}\right)^{k-i}$
+is the product of two expressions both increasing in $i$, and hence is increasing as well. ∎
+Now we state and prove our result on the time to see first inner Pareto optimum.
 Lemma 6.
 Let $k\leq n^{\prime}/2$.
-Consider using SMS-EMOA with $\mu\geq M$ to optimize the $m$OJZJ problem.
-Then after at most $e\mu(mk)^{k}(1+\ln m)$ iterations in expectation, the population (and also the populations afterwards) contains at least one inner Pareto optimum.
+Consider using the SMS-EMOA with $\mu\geq\overline{M}$ to optimize the $m$OJZJ problem.
+Then after at most $e\mu(mk/2)^{k}(1+\ln m)$ iterations in expectation, the population (and also the populations afterwards) contains at least one inner Pareto optimum.
+Proof.
+Denote by $f=(f_{1},\dots,f_{m})$ the $m$OJZJ function.
+This lemma is already proven if the initial population has at least one inner Pareto optimum.
+Now we consider that the initial population $P$ contains no inner Pareto optimum.
+For any $x\in\{0,1\}^{n},$ let $I(x)=\{i\in[1..m/2]\mid|x_{B_{i}}|_{1}\in[k..n^{\prime}-k]\}$ and $L(x)=|I(x)|$.
+Let $L(P)=\max\{L(x)\mid x\in P\}$.
+Then $L(P)\geq 0$ and we consider the first time to reach a population $P$ with $L(P)=m/2$, that is, to reach an inner Pareto optimum.
+We first show that $L(P)$ cannot decrease.
+Let $S_{P}=\{x\in P\mid L(x)=L(P)\}$.
+Let $x\in S_{P}$.
+From Lemma 4, we know that for the next population, there is $y$ such that $y\succeq x$.
+Hence, $f_{2i-1}(y)\geq f_{2i-1}(x)$ for any $i\in[1..m/2]$.
+If $|x_{B_{i}}|_{1}\geq k$, then $f_{2i-1}(x)=k+|x_{B_{i}}|_{1}$.
+Thus $f_{2i-1}(y)\geq k+|x_{B_{i}}|_{1}\geq 2k$ and hence $|y_{B_{i}}|_{1}\geq k$ from the definition of mOJZJ.
+An analogous argument based on $f_{2i}$ shows that $|x_{B_{i}}|_{1}\leq n^{\prime}-k$ implies $|y_{B_{i}}|_{1}\leq n^{\prime}-k$.
+Hence $L(y)\geq L(x)$, which shows that $L(P)$ does not decrease.
+Now we consider the increase of $L(P)$ when $L(P)<m/2$.
+Let $x\in P$ with $L(x)=L(P)$.
+The probability to select $x$ as a parent is $1/\mu$.
+For a specific $i\in[1..m/2]\setminus I(x)$, we know that $|x_{B_{i}}|_{1}\in[0..k-1]\cup[n-k+1..n]$.
+W.l.o.g., let $|x_{B_{i}}|_{1}\in[0..k-1]$.
+Then the probability of generating an individual $y$ with $|y_{B_{i}}|\in[k..n^{\prime}-k]$ from $x$ is at least
+$\binom{n^{\prime}-|x_{B_{i}}|_{1}}{k-|x_{B_{i}}|_{1}}\mathopen{}\mathclose{{}\left(\frac{1}{n}}\right)^{k-|x_{B_{i}}|_{1}}\mathopen{}\mathclose{{}\left(1-\frac{1}{n}}\right)^{n-(k-|x_{B_{i}}|_{1})}\geq\binom{n^{\prime}}{k}\frac{1}{n^{k}}\frac{1}{e}\geq\mathopen{}\mathclose{{}\left(\frac{n^{\prime}}{k}}\right)^{k}\frac{1}{en^{k}}=\frac{1}{e}\mathopen{}\mathclose{{}\left(\frac{2}{mk}}\right)^{k},$ (1)
+where the first inequality is from Lemma 5, and the last equality uses $n^{\prime}=2n/m$.
+Note that $|[1..m/2]\setminus I(x)|=m/2-L(P)$.
+Hence, the probability of increasing $L(P)$ in one iteration is at least
+$\frac{m/2-L(P)}{\mu}\frac{1}{e}\mathopen{}\mathclose{{}\left(\frac{2}{mk}}\right)^{k},$
+and thus the expected number of iterations to witness such an improvement is at most $e\mu(mk/2)^{k}/(m/2-L(P))$.
+Hence, the expected number of iterations to generate an inner Pareto optimum is at most
+$\sum_{i=0}^{m/2-1}\frac{e\mu}{m/2-i}\mathopen{}\mathclose{{}\left(\frac{mk}{2}}\right)^{k}\leq e\mu\mathopen{}\mathclose{{}\left(\frac{mk}{2}}\right)^{k}(1+\ln m),$
+where the last inequality uses $\sum_{j-1}^{m/2}\frac{1}{j}\leq 1+\ln(m/2)\leq 1+\ln m$. ∎
 In the next lemma, we consider the stage of covering all inner Pareto front points once at least one such point is in the population.
@@
 Let $k\leq n^{\prime}/2$.
-Consider using SMS-EMOA with $\mu\geq M$ to optimize the $m$OJZJ problem.
+Consider using the SMS-EMOA with $\mu\geq\overline{M}$ to optimize the $m$OJZJ problem.
 Assume that the current population contains at least one inner Pareto optimum.
-Then after at most $en\mu M$ iterations in expectation, all inner Pareto front points are covered.
+Then after at most $en\mu(n^{\prime}-2k+1)^{m/2}$ iterations in expectation, all inner Pareto front points are covered.
+Proof.
+Let $u\in F^{*}$ be an uncovered inner Pareto front point that has some neighbor covered by the population $P$.
+Since such a neighbor is also a Pareto front point, we know that it is maintained in all future generations.
+The probability to generate an offspring with function value $u$ is at least $\mathopen{}\mathclose{{}\left(1-\frac{1}{n}}\right)^{n-1}\frac{1}{n\mu}\geq\frac{1}{en\mu}$.
+Hence the expected number of iterations to witness such an offspring is at most $en\mu$.
+From Lemma 4, we know that once such an offspring is generated, $u$ will be covered in all future populations.
+Noting that there are at most $(n^{\prime}-2k+1)^{m/2}$ inner Pareto front points, we know that after an expected number of at most $en\mu(n^{\prime}-2k+1)^{m/2}$ iterations, all inner Pareto front points are covered.
+∎
 The last stage is to cover the remaining Pareto front points.
-The following lemma shows the runtime.
+The following lemma bounds the runtime of this phase.
 The key of the proof is that we can divide the Pareto front into several levels and any individual in the $i$-th level can be generated from a point in the $(i-1)$-th level by flipping the right $k$ bits.
@@
 Let $k\leq n^{\prime}/2$.
-Consider using SMS-EMOA with $\mu\geq M$ to optimize the $m$OJZJ problem.
+Consider using the SMS-EMOA with $\mu\geq\overline{M}$ to optimize the $m$OJZJ problem.
 Assume that the current population covers all inner Pareto front points.
-Then after at most $eM\mu n^{k}$ iterations in expectation, the full Pareto front is covered.
-Summing up the runtime of all stages from Lemmas 5 to 7, we have the following theorem for the runtime of the full coverage of the Pareto front.
+Then after at most $e((n^{\prime}-2k+3)^{m/2}-(n^{\prime}-2k+1)^{m/2})\mu n^{k}$ iterations in expectation, the full Pareto front is covered.
+Proof.
+We say that a block $x_{B_{i}}$ of an individual $x$ is extreme if $x_{B_{i}}\in\{(0,\dots,0),(1,\dots,1)\}$.
+We divide the Pareto front points into $m/2+1$ levels $L_{0},L_{1},\dots,L_{m/2}$, where $L_{j}$ consists of the objective values of all Pareto optima having exactly $j$ extreme blocks.
+Obviously, all inner Pareto front points belong to the 00-th level.
+Note that any point in $j$-th level ($j\geq 1$) can be generated from an individual having a certain objective value in $L_{j-1}$ via flipping specific $k$ bits, leaving the other bits unchanged.
+Hence, the probability to generate any point in $j$-th level condition on that all $(j-1)$-th level points are covered is at least
+$\frac{1}{\mu}\mathopen{}\mathclose{{}\left(1-\frac{1}{n}}\right)^{n-k}\mathopen{}\mathclose{{}\left(\frac{1}{n}}\right)^{k}\geq\frac{1}{e\mu n^{k}}.$
+From Lemma 4, we know that such a Pareto front point is reached in at most $e\mu n^{k}$ iterations in expectation, and will survive in all future populations.
+Consider the process of covering all non-inner Pareto front points in the order of increasing level.
+Since the 00-th level points are already covered from our assumption, and since there are at most $((n^{\prime}-2k+3)^{m/2}-(n^{\prime}-2k+1)^{m/2})$ non-inner Pareto front points, we know that the remaining levels are covered in at most $e((n^{\prime}-2k+3)^{m/2}-(n^{\prime}-2k+1)^{m/2})\mu n^{k}$ iterations in expectation.
+∎
+Summing up the runtimes of the three stages from Lemmas 6 to 8, we obtain the following theorem on the runtime of the full coverage of the Pareto front.
 Theorem 9.
-Let $k\leq n^{\prime}/2$.
-Consider using SMS-EMOA with $\mu\geq M$ to optimize the $m$OJZJ problem.
-Then after at most $e\mu(mk)^{k}(1+\ln m)+e\mu Mn+e\mu Mn^{k}$ iterations ($\mu+e\mu(mk)^{k}(1+\ln m)+e\mu Mn+e\mu Mn^{k}$ function evaluations) in expectation, the full Pareto front is covered.
-Since $k\leq n^{\prime}/2=n/m$ and $M\geq 2^{m/2}$, we easily see that both runtime expressions are $O(\mu Mn^{k})$, even when allowing $k$, $m$ and $\mu$ to depend on $n$.
-Runtime of the GSEMO on $m$OJZJ
+Consider the $m$OJZJ problem with problem size $n$, number of objectives $m$, and gap size $k$.
+Write $n^{\prime}=2n/m$ for the block length.
+Assume that $k\leq n^{\prime}/2$.
+Denote by $M=(n^{\prime}-2k+3)^{m/2}$ the size of the Pareto front and by $\overline{M}$ the size of the largest incomparable set (for which we know that $\overline{M}\leq(2n/m+1)^{m/2}$).
+Consider using the SMS-EMOA with $\mu\geq\overline{M}$ to solve this problem.
+Then after at most $e\mu(mk/2)^{k}(1+\ln m)+e\mu Mn^{k}=O(\mu Mn^{k})$ iterations ($\mu+e\mu(mk/2)^{k}(1+\ln m)+e\mu Mn^{k}=O(\mu Mn^{k})$ function evaluations) in expectation, the full Pareto front is covered.
+Since $k\leq n^{\prime}/2=n/m$ and $M\geq 2^{m/2}$, we easily see that both runtime expressions in the theorem are $O(\mu Mn^{k})$, even when allowing $k$, $m$ and $\mu$ to depend on $n$.
+4.3 Runtime of the GSEMO on $m$OJZJ
 Since our arguments above can easily be extended to analyze the runtime of the GSEMO on the $m$OJZJ, we quickly do so for reasons of completeness.
-The GSEMO is a multiobjective counterpart of the ${(1+1)}$ EA.
-The initial population consists of a single randomly generated solution.
-In each iteration, a solution is uniformly at random picked from the population to generate an offspring via standard bit-wise mutation.
+The GSEMO is the multiobjective counterpart of the single-objective $(1+1)$ evolutionary algorithm.
+The initial population of the GSEMO consists of a single randomly generated solution.
+In each iteration, a solution is picked uniformly at random from the population to generate an offspring via standard bit-wise mutation.
 If this offspring is not dominated by any solution in the population, it is added to the population and all solutions weakly dominated by it are removed.
 It is not difficult to see that all solutions in the population of the GSEMO are mutually non-dominated.
-Hence, the population size of the GSEMO is at most $M$ by Lemma 3.
+Hence, the population size of the GSEMO is at most $\overline{M}$ by Lemma 3.
 When analyzing the runtime of the GSEMO on $m$OJZJ, the main difference is that SMS-EMOA requires a statement like Lemma 4 to ensure that previous progress is not lost via unlucky selection decisions.
@@
 Consider using the GSEMO to optimize the $m$OJZJ problem.
-Then after an expected number of at most $eM(mk)^{k}(1+\ln m)+eM^{2}n+eM^{2}n^{k}$ iterations (or $1+eM(mk)^{k}(1+\ln m)+eM^{2}n+eM^{2}n^{k}$ fitness evaluations), the full Pareto front is covered.
-Reduced Impact of Stochastic Population Update
-Bian, Zhou, Li, and Qian (2023) proposed a stochastic population update mechanism for the SMS-EMOA and proved that, somewhat unexpectedly given the state of the art, it can lead to significant performance gains.
+Then after an expected number of at most $e\overline{M}(mk/2)^{k}(1+\ln m)+eM\overline{M}n^{k}$ iterations (or $1+e\overline{M}(mk/2)^{k}(1+\ln m)+eM\overline{M}n^{k}$ fitness evaluations), the full Pareto front is covered.
+5 Reduced Impact of Stochastic Population Update
+Bian et al.
+[BZLQ23] proposed a stochastic population update mechanism for the SMS-EMOA and proved, somewhat unexpectedly given the state of the art, that it can lead to significant performance gains.
 More precisely, it was proven that the classic SMS-EMOA with reasonable population size $\mu=\Theta(n)$ solves the bi-objective OJZJ problem in an expected number of $O(n^{k+2})\cap\Omega(n^{k+1})$ function evaluations, whereas for the SMS-EMOA with stochastic population update, $O(n^{k+2}\min\{1,2^{-k/4}n\})$ suffice.
 Hence, for $k=\omega(\log n)$, a super-polynomial speed-up was shown.
-In this section, we extend the analysis of (Bian et al.
-2023) to $m$ objectives.
+In this section, we extend the analysis of [BZLQ23] to $m$ objectives.
 Unfortunately, we will observe that the larger population sizes necessary here reduce the impact of the stochastic population update.
-As in (Bian et al.
-2023), we have no proven tight lower bounds for the SMS-EMOA with stochastic population update, but our upper-bound proofs suggest that the reduced effect of the stochastic population update on the runtime guarantee is real, that is, the impact on the true runtime is diminishing with the larger population sizes necessary in the many-objective setting.
-Stochastic Population Update
+As in [BZLQ23], we have no proven tight lower bounds for the SMS-EMOA with stochastic population update, but our upper-bound proofs suggest that the reduced effect of the stochastic population update on the runtime guarantee is real, that is, the impact on the true runtime is diminishing with the larger population sizes necessary in the many-objective setting.
+As a side result, we simply and moderately improve the analysis of how the stochastic population update helps to traverse the valley of low-quality solutions of the OJZJ benchmarks.
+Whereas in [BZLQ23] a complex drift argument was used to show a speed-up of $\mu/2^{k/4}$, we give a simple waiting time argument that yields a speed-up of $\mu/(2^{k}-1)$, both in the bi-objective case of [BZLQ23] and in the general case.
+5.1 Stochastic Population Update
 The rough idea of the stochastic selection proposed by Bian et al.
-(2023) is that a random half of the individuals survive into the next generation regardless of their quality.
+[BZLQ23] is that a random half of the individuals survive into the next generation regardless of their quality.
 The individual to be discarded is chosen from the remaining individuals according to non-dominated sorting and hypervolume contribution as in the classic SMS-EMOA.
-This approach resembles the random mixing of two acceptance operators in the Move Acceptance Hyper-Heuristic studied recently (Lehre and Özcan 2013; Lissovoi, Oliveto, and Warwicker 2023; Doerr et al.
-2023).
+This approach resembles the random mixing of two acceptance operators in the Move Acceptance Hyper-Heuristic studied recently [LÖ13, LOW23, DDLS23].
 More precisely, after generating the offspring $x^{\prime}$, the SMS-EMOA with stochastic selection chooses from $P_{t}\cup\{x^{\prime}\}$ a set $R^{\prime}$ of $\lfloor(\mu+1)/2\rfloor$ solutions randomly with replacement.
 The individual $z^{\prime}$ to be removed from $P_{t}\cup\{x^{\prime}\}$ is then determined via non-dominated sorting and hypervolume contribution applied to $R^{\prime}$ only.
-We note that with this mechanism, even the worst solution enters the next generation with probability at least $1/2$.
-Runtime
+We note that with this mechanism, any solution enters the next generation with probability at least $1/2$.
 To analyze the runtime of the SMS-EMOA with stochastic population update, we first derive and formulate separately two insights on the survival of solutions.
@@
 Lemma 11.
-Consider using the SMS-EMOA with stochastic population update to optimize the $m$OJZJ problem.
+Consider using the SMS-EMOA with stochastic population update to optimize any specific problem.
 For any iteration $t$ and $x\in R_{t}$, we have that there is a $y$ with $f(y)=f(x)$ in $P_{t+1}$ with probability at least $1/2$.
+Proof.
+We simply calculate the probability of $x$ being removed after the population update.
+Since $x$ can be removed only if it is selected in $R^{\prime}$, we know that it will be removed with probability at most $\frac{\lfloor(\mu+1)/2\rfloor}{\mu+1}\leq 1/2$, and hence $f(x)$ will be maintained in $P_{t+1}$ with probability at least $1/2$.
+∎
 Lemma 12.
-Consider using the SMS-EMOA with stochastic population update and with $\mu\geq 2(M+1)$ to optimize the $m$OJZJ problem.
+Consider using the SMS-EMOA with stochastic population update and with $\mu\geq 2\overline{M}+1$ to optimize any specific problem, where $\overline{M}$ is the maximum size of a set of incomparable solutions for this problem.
 If at some time $t$ the combined parent and offspring population $R_{t}$ of the SMS-EMOA contains some solution $x$, then at any later time $s>t$ the parent population $P_{s}$ contains a solution $y$ such that $y\succeq x$.
-Since the proofs of Lemmas 5 and 6 for the classic SMS-EMOA mostly relied on elementary properties of standard bit-wise mutation and on the survival guarantee of Lemma 4, we can now use the survival guarantee of Lemma 11 to obtain analogous results for the SMS-EMOA with stochastic selection (at the price of requiring essentially twice the population size).
+Proof.
+If $x\notin R^{\prime}$, then it survives automatically into the next generation.
+If $x\in R^{\prime}$, then it is subject to a selection as in the standard SMS-EMOA.
+Since $R^{\prime}=\lfloor(\mu+1)/2\rfloor>\overline{M}$, Lemma 4 gives the desired $y$.
+∎
+5.2 Runtime Analysis for the SMS-EMOA With Stochastic Population Update
+We now analyze the runtime of the SMS-EMOA with stochastic population update on the many-objective OJZJ problem.
+We start by distilling why the SMS-EMOA profits from the stochastic population update and tightening the analysis of [BZLQ23].
+Our analysis is also significantly simpler than the previous one, which might easy future works in this direction.
+The core of the analysis in [BZLQ23] is proving that the SMS-EMOA can profit from the stochastic population update by accepting solutions in the valley of low-quality solutions of the bi-objective OJZJ benchmark.
+When extracted from the long proof in [BZLQ23], this would give the following lemma.
+Lemma 13 ([BZLQ23]).
+Consider using the SMS-EMOA with stochastic population update to optimize the bi-objective OJZJ benchmark.
+Assume that $\mu\geq 2(M+1)$.
+Assume that at some time $t$, the population contains an individual with $k$ ones.
+Then the expected number of additional iterations to generate the all-zero solution is at most $\min\{e\mu n^{k},(e\mu n^{k/2}+1)2e\mu^{2}n^{k/2}/2^{k/4}\}=e\mu n^{k}\min\{1,(1+o(1))2e\mu/2^{k/4}\}$.
+The proof of this lemma relies on a technical application of the additive drift theorem of He and Yao [HY01], in particular, with a non-intuitive potential function.
+Since the valleys of low-quality solutions of the OJZJ benchmark have the same shape regardless of the number of objectives, we could use the elementary lemmas from the previous subsection to extend the lemma above to the many-objective setting.
+Since we are not fully satisfied with the complex proof of Lemma 13, we propose an alternative one that appears both simpler and gives a tighter result.
+We formulate it in the many-objective setting, but note that it includes the bi-objective setting.
+In terms of the bound, our result replaces the factor of $(1+o(1))2^{-k/4}$, which is the term responsible for the runtime improvement, by a factor of $2^{-k+1}$, so essentially the fourth power of the previous factor.
+Lemma 14.
+Consider using the SMS-EMOA with stochastic population update to optimize the $m$-objective OJZJ benchmark, $m\geq 2$.
+Assume that $\mu\geq 2\overline{M}+1$.
+Assume that at some time $t$, the population contains an individual $x$ with $k$ ones in the $i$-th block.
+Then the expected number of additional iterations to generate an individual $z$ with only zeros in the $i$-th block and with $f_{j}(z)=f_{j}(x)$ for all $j\in[1..m]\setminus\{2i-1,2i\}$ is at most $e\mu n^{k}\min\{1,4e\mu/2^{k}\}$.
+Proof.
+The proof of this lemma consists of a simple waiting time argument, that is, we compute the probability $p$ that the desired solution $z$ is generated from $x$ in at most two iterations and then bound the expected waiting time by $2/p$.
+Let $x$ be as in the assumptions of this lemma.
+Let $X$ be the set of indices of ones it has in the $i$-th block.
+Let $Y\subseteq X$.
+Let $E_{Y}$ be the event that in the first iteration, (i) $x$ is selected as parent for mutation, (ii) it is mutated into an individual $y$ that agrees with $x$ in all bits except for exactly the ones in $Y$, and (iii) this $y$ is taken into the next parent population (despite being dominated by $x$ when $Y\notin\{\emptyset,X\}$), and in the next iteration, (iv) this $y$ is selected as parent for mutation and (v) mutated into a solution $z$ that agrees with $x$ in all bits except for exactly the ones in $X$.
+The probabilities for these steps to happen as described are $1/\mu$ for step (i) and (iv), $(1/n)^{|Y|}(1-1/n)^{n-|Y|}$ for step (ii), at least $1/2$ for step (iii) by Lemma 11, and $(1/n)^{k-|Y|}(1-1/n)^{n-(k-|Y|)}$ for step (v), noting that here exactly the bits in $X\setminus Y$ have to be flipped.
+Multiplying these probabilities, we see that $\Pr[E_{Y}]\geq\tfrac{1}{2}(1-1/n)^{2n-k}\mu^{-2}n^{-k}\geq\tfrac{1}{2}e^{-2}\mu^{-2}n^{-k}$.
+There are $2^{k}$ disjoint events $E_{Y}$, each ensuring that after two iterations the population contains an individual $z$ as desired.
+Hence the probability $p$ to obtain such a $z$ within two iterations is at least $p\geq\sum_{Y}\Pr[E_{Y}]\geq 2^{k}\cdot\tfrac{1}{2}e^{-2}\mu^{-2}n^{-k}$.
+By Lemma 12, the population always contains an individual satisfying the assumptions on $x$ and having the same objective value as $x$.
+Hence the time $T$ to generate the desired individual $z$ is stochastically dominated by two times a geometric distribution with success rate $p$, hence $E[T]\leq 2/p\leq 4e^{2}\mu^{2}n^{k}/2^{k}$.
+∎
+Since the proofs of Lemmas 6 and 7 for the classic SMS-EMOA mostly relied on elementary properties of standard bit-wise mutation and on the survival guarantee of Lemma 4, we can now use the survival guarantee of Lemma 12 to obtain analogous results for the SMS-EMOA with stochastic selection (at the price of requiring essentially twice the population size).
 This yields the following estimates for the time to obtain at least one inner Pareto optimum and the time to cover all inner Pareto front points starting from a population with at least one inner Pareto optimum.
 Lemma 15.
-Let $k\leq n^{\prime}/2$. Consider using the SMS-EMOA with stochastic population update and with $\mu\geq 2(M+1)$ to optimize the $m$OJZJ problem. Then
-  • after at most $e\mu(mk)^{k}(1+\ln m)$ iterations in expectation, the population (and also the populations afterwards) contains at least one inner Pareto optimum;
-  • after another at most $en\mu M$ iterations in expectation, all inner Pareto front points are covered.
+Let $k\leq n^{\prime}/2$. Consider using the SMS-EMOA with stochastic population update and with $\mu\geq 2\overline{M}+1$ to optimize the $m$OJZJ problem. Then
+  • after at most $e\mu(mk/2)^{k}(1+\ln m)$ iterations in expectation, the population (and also the populations afterwards) contains at least one inner Pareto optimum;
+  • after another at most $en\mu(n^{\prime}-2k+1)^{m/2}$ iterations in expectation, all inner Pareto front points are covered.
 Now we consider the runtime for the full coverage of the Pareto front after all inner Pareto front points are covered.
-Similar to the proof of Lemma 7 for the original SMS-EMOA, we partition the Pareto front points into $m/2+1$ levels based on the number of extreme blocks they contain.
-The difference here is that we consider the effect of the gap points, similar to Bian et al.
-(2023).
+This analysis follows precisely the proof of Lemma 8 for the original SMS-EMOA except that we replace the estimate for expected time to obtain an extremal block, $e\mu n^{k}$ in that lemma, by the potentially lower number $e\mu n^{k}\min\{1,4e\mu/2^{k}\}$.
+This immediately gives the following result.
 Lemma 16.
 Let $k\leq n^{\prime}/2$.
-Consider using the SMS-EMOA with stochastic population update and with $\mu\geq 2(M+1)$ to optimize the $m$OJZJ problem.
+Consider using the SMS-EMOA with stochastic population update and with $\mu\geq 2\overline{M}+1$ to optimize the $m$OJZJ problem.
 Assume that the current population covers all inner Pareto front points.
-Then after at most $O(\min\{\frac{\mu k^{1/2}}{2^{k/2}},1\}M\mu n^{k})$ iterations in expectation, the full Pareto front is covered.
-Combining Lemmas 12 and 13, we have the runtime of the SMS-EMOA with stochastic population update in the following theorem.
+Then after at most $e\mu((n^{\prime}-2k+3)^{m/2}-(n^{\prime}-2k+1)^{m/2})n^{k}\min\{1,4e\mu/2^{k}\}$ iterations in expectation, the full Pareto front is covered.
+Combining Lemmas 15 and 16, we have the runtime of the SMS-EMOA with stochastic population update in the following theorem.
+It includes the case of two objectives, and then, with $\overline{M}=M=(n-2k+3)$, improves the previous result for this case, Theorem 2 in [BZLQ23], in particular by replacing the $2^{k/4}$ term by $2^{k}$.
 Theorem 17.
 Let $k\leq n^{\prime}/2$.
-Consider using SMS-EMOA with $\mu\geq 2(M+1)$ and stochastic population update to optimize the $m$OJZJ problem.
-Then after at most $e\mu(mk)^{k}(1+\ln m)+en\mu M+O(\min\{\frac{\mu k^{1/2}}{2^{k/2}},1\}M\mu n^{k})$ iterations in expectation, the full Pareto front is covered.
-Comparing the runtime guarantees or Theorem 8 (classic SMS-EMOA) and Theorem 14 (SMS-EMOA with stochastic selection), we see that stochastic selection can at most lead to a speed-up by a factor of order $2^{k/2}/\mu k^{1/2}$.
+Consider using SMS-EMOA with $\mu\geq 2\overline{M}+1$ and stochastic population update to optimize the $m$OJZJ problem.
+Then after at most $e\mu(mk/2)^{k}(1+\ln m)+{en\mu(n^{\prime}-2k+1)^{m/2}+e\mu((n^{\prime}-2k+3)^{m/2}-(n^{\prime}-2k+1)^{m/2})n^{k}\min\{1,4e\mu/2^{k}\}}=O(\mu Mn^{k}\min\{1,4e\mu/2^{k}\})$ iterations in expectation, the full Pareto front is covered.
+Comparing the runtime guarantees of Theorem 9 (classic SMS-EMOA) and Theorem 17 (SMS-EMOA with stochastic selection), we see that stochastic selection can at most lead to a speed-up by a factor of order $2^{k}/\mu$.
 Now $\mu$ is at least $\Omega(M)=\Omega((2n/m-2k+3)^{m/2})$.
 Consequently, the advantage of stochastic selection is rapidly decreasing with growing numbers of objectives and vanishes, e.g., when $m\geq k$.
-Heavy-Tailed Mutation Helps
-In the previous section, we saw that the advantages of a stochastic population update do not well generalize from bi-objective optimization to many-objective optimization.
+6 Heavy-Tailed Mutation Helps
+In the previous section, we saw that the advantage of the stochastic population update does not generalize well from bi-objective to many-objective optimization.
 We now regard another design choice that so far was only analyzed in bi-objective optimization, namely a heavy-tailed mutation operator.
-We shall prove that the $k^{\Omega(k)}$ factor speed-up observed in bi-objective optimization extends to many objectives.
-Heavy-Tailed Mutation
-Different from the standard bit-wise mutation operator, which flips each bit independently with probability $1/n$, the heavy-tailed mutation operator proposed in (Doerr et al.
-2017) flips each bit independently with probability $\alpha/n$, where $\alpha$ follows a power-law distribution with parameter $\beta$.
+For this, we shall prove that the $k^{\Omega(k)}$ factor speed-up observed in bi-objective optimization extends to many objectives.
+6.1 Heavy-Tailed Mutation
+Different from the standard bit-wise mutation operator, which flips each bit independently with probability $1/n$, the heavy-tailed mutation operator proposed in [DLMN17] flips each bit independently with probability $\alpha/n$, where $\alpha$ follows a power-law distribution with parameter $\beta$.
 The number $\alpha$ is sampled anew in each application of the heavy-tailed mutation operator.
@@
 Let $n\in\mathbb{N}$ and $\beta>1$.
-We say $\alpha$ follows a power-law distribution with (negative) exponent $\beta$ if for all $i\in[1..n/2]$, we have $\Pr[\alpha=i]=\left(\sum_{j=1}^{n/2}k^{-\beta}\right)^{-1}i^{-\beta}$.
-The power-law feature extends to the heavy-tailed mutation operator in the sense that it generates an offspring with Hamming distance $j$ from the parent with probability $\Omega(j^{-\beta})$.
+We say that $\alpha$ follows a power-law distribution with (negative) exponent $\beta$ if for all $i\in[1..n/2]$, we have $\Pr[\alpha=i]=\mathopen{}\mathclose{{}\left(C_{n/2}^{\beta}}\right)^{-1}i^{-\beta}$, where $C_{n/2}^{\beta}=\sum_{j=1}^{n/2}j^{-\beta}$.
+The power-law in the choice of $\alpha$ extends to the offspring distribution in the way that an offspring has Hamming distance $j$ from the parent with probability $\Omega(j^{-\beta})$ [DLMN17].
 This facilitates larger jumps in the search space, and thus escaping from a local optimum.
-The heavy-tailed mutation operator has resulted in asymptotic performance gains by a factor of $k^{\Omega(k)}$ for the ${(1+1)}$ EA optimizing single-objective (classic and generalized) Jump functions with gap size $k$ (Doerr et al.
-2017; Bambury, Bultel, and Doerr 2021).
-For the bi-objective OJZJ benchmark, again a speed-up of $k^{\Omega(k)}$ was proven when optimized via the GSEMO (Zheng and Doerr 2023d) and the NSGA-II (Doerr and Qu 2023a).
-Several other positive theoretical results exist for this heavy-tailed mutation, or more generally, other heavy-tailed parameter choices (Friedrich, Quinzan, and Wagner 2018; Wu, Qian, and Tang 2018; Quinzan et al.
-2021; Corus, Oliveto, and Yazdani 2021; Dang et al.
-2022; Antipov, Buzdalov, and Doerr 2022; Doerr, Ghannane, and Ibn Brahim 2022; Doerr and Rajabi 2023).
+We recall that for standard bit-wise mutation, the offspring has Hamming distance $j$ from the parent with probability is $j^{-\Omega(j)}$ only.
+Since this observation was formulated in [DLMN17, Lemma 5.1] in a slightly impractical manner (with implicit constants independent of $\beta$, but only applying to distances in $(\beta-1,n/2]\cup\{1\}$), we first give a formulation that is more useful for most purposes.
+It treats $\beta$ as a constant, which makes a lot of sense in the light of all previous results, which gave the best performances for constant values of $\beta$.
+Lemma 19.
+Let $\beta>1$.
+Then there is a $c>0$ such that the following holds.
+Let $n\in\mathbb{N}$.
+Let $x\in\{0,1\}^{n}$ and $y$ be generated by applying the heavy-tailed mutation (with parameter $\beta$) to $x$.
+Let $H(x,y)$ be the Hamming distance between $x$ and $y$, let $j\in[1..n/2]$, and let $P_{j}^{\beta}:=P_{j}^{\beta}(n):=\Pr[H(x,y)=j]$.
+Then
+$P_{j}^{\beta}\geq cj^{-{\beta}}.$
+Proof.
+If $\beta-1>\sqrt{n}$, then any $c>0$ such that $c\leq\min\{P_{j}^{\beta}(n)\mid n\in[1..(\beta-1)^{2}],j\in[1..n/2]\}$ trivially satisfied our claim.
+Hence, in the following, let $\beta-1\leq\sqrt{n}$.
+If $j\in(\beta-1,n/2]\cup\{1\}$, then [DLMN17, Lemma 5.1 (1) and (3)] gives our result.
+If $j\in(1,\beta-1]$, then $j\leq\beta-1\leq\sqrt{n}$ and we conclude from [ZD23d, Lemma 16] that $P_{j}^{\beta}\geq\frac{\beta-1}{2\sqrt{2\pi}e^{8\sqrt{2}+13}\beta}j^{-\beta}$.
+Hence $c=\frac{\beta-1}{2\sqrt{2\pi}e^{8\sqrt{2}+13}\beta}$ suffices in this case.
+∎
+The heavy-tailed mutation operator has resulted in asymptotic performance gains by a factor of $k^{\Omega(k)}$ for the $(1+1)$ EA optimizing single-objective (classic and generalized) Jump functions with gap size $k$ [DLMN17, BBD24].
+For the bi-objective OJZJ benchmark, again a speed-up of $k^{\Omega(k)}$ was proven when optimized via the GSEMO [ZD23c] and the NSGA-II [DQ23a].
+Several other positive theoretical results exist for this heavy-tailed mutation, or more generally, other heavy-tailed parameter choices [FQW18, WQT18, QGWF21, COY21, DELQ22, ABD22, DR23, DGI24].
 Prior to this work, no theoretical analysis of the heavy-tailed mutation operator in many-objective optimization existed.
-Runtime
+6.2 Runtime
 We now analyze the runtime of the SMS-EMOA using the heavy-tailed mutation operator instead of bit-wise mutation, the standard choice for this algorithm.
@@
 The assertion of Lemma 4 extends to the SMS-EMOA with heavy-tailed mutation.
-With proof ideas similar to those in Lemma 5 to 7, but noticing the different probabilities of generating solutions with a specific Hamming distance, we obtain the following runtime guarantee.
+With proof ideas similar to those in Lemma 6 to 8, we prove the following three runtime estimates.
+The main differences are the different probabilities for generating a certain solution.
+However, these require somewhat different computations in the following lemma, which therefore appears slightly technical.
+We note that we did not optimize the result here, but were content with arriving at a bound that is small compared to the time needed to generate the extremal Pareto optima.
+Lemma 21.
+Let $k\leq n^{\prime}/2$.
+Consider using the SMS-EMOA with heavy-tailed mutation and with $\mu\geq\overline{M}$ to optimize the $m$OJZJ problem.
+Then after at most $O\mathopen{}\mathclose{{}\left((em/2)^{k}k^{\beta}\mu\ln m}\right)$ (for $m\geq 4$) or $O\mathopen{}\mathclose{{}\left(\binom{n}{k}k^{\beta}\mu/(n-k+1)}\right)$ (for $m=2$) iterations in expectation, the population (and also the populations afterwards) contains at least one inner Pareto optimum.
+Proof.
+We use the same general as in Lemma 6 except for the following probability estimates.
+From Lemma 19, we know that the probability to generate a $y$ with $|y_{B_{i}}|_{1}\in[k..n^{\prime}-k]$ from $x$ with $|x_{B_{i}}|_{1}\in[0..k-1]$ is at least
+$\Omega\mathopen{}\mathclose{{}\left(\frac{\binom{n^{\prime}-|x_{B_{i}}|_{1}}{k-|x_{B_{i}}|_{1}}(k-|x_{B_{i}}|_{1})^{-\beta}}{\binom{n}{k-|x_{B_{i}}|_{1}}}}\right).$ (2)
+Let first $m\geq 4$. For all $j\in[0..k-1]$, we compute
+$\begin{split}&\mathopen{}\mathclose{{}\left(\frac{\binom{n^{\prime}-(j+1)}{k-(j+1)}}{\binom{n}{k-(j+1)}(k-(j+1))^{\beta}}}\right)\bigg{/}\mathopen{}\mathclose{{}\left(\frac{\binom{n^{\prime}-j}{k-j}}{\binom{n}{k-j}(k-j)^{\beta}}}\right)\\ ={}&{}\mathopen{}\mathclose{{}\left(\frac{(n^{\prime}-(j+1))!(k-(j+1))!(n-k+j+1)!}{(k-(j+1))!(n^{\prime}-k)!n!(k-(j+1))^{\beta}}}\right)\\ &{}\bigg{/}\mathopen{}\mathclose{{}\left(\frac{(n^{\prime}-j)!(k-j)!(n-k+j)!}{(k-j)!(n^{\prime}-k)!n!(k-j)^{\beta}}}\right)\\ ={}&{}\frac{(n^{\prime}-(j+1))!(n-k+j+1)!(k-j)^{\beta}}{(n^{\prime}-j)!(n-k+j)!(k-(j+1))^{\beta}}=\frac{n-k+j+1}{n^{\prime}-j}\mathopen{}\mathclose{{}\left(\frac{k-j}{k-(j+1)}}\right)^{\beta}\\ \geq{}&{}\frac{n-k+1}{n^{\prime}}\geq\frac{mn^{\prime}/2-n^{\prime}/2+1}{n^{\prime}}>1,\end{split}$ (3)
+where the antepenultimate inequality uses $j\geq 0$, and the last inequality uses $m\geq 4$. Hence the expression in (2) is smallest for $|x_{B_{i}}|_{1}=0$, and hence at least of the order of
+$\frac{\binom{n^{\prime}}{k}}{\binom{n}{k}}k^{-\beta}\geq\mathopen{}\mathclose{{}\left(\frac{n^{\prime}}{en}}\right)^{k}k^{-\beta}=\mathopen{}\mathclose{{}\left(\frac{2}{em}}\right)^{k}k^{-\beta}.$ (4)
+For $m=2$, we have $n=n^{\prime}$. Since
+$\binom{n-j}{k-j}=\binom{n-j}{n-k}\geq\binom{n-(k-1)}{n-k}=n-k+1$
+for $j\in[0..k-1]$, and
+$\frac{n-k+1}{\binom{n}{k-|x_{B_{i}}|_{1}}}(k-|x_{B_{i}}|_{1})^{-\beta}\geq\frac{n-k+1}{\binom{n}{k}}k^{-\beta},$ (5)
+we can estimate the expression in (2) by $\Omega\mathopen{}\mathclose{{}\left((n-k+1)k^{-\beta}/\binom{n}{k}}\right)$.
+Replacing (1) by (4) or (5) in the proof of Lemma 6, we obtain that the expected number of iterations to reach an inner Pareto optimum is $O\mathopen{}\mathclose{{}\left((em/2)^{k}k^{\beta}\mu\ln m}\right)$ for $m\geq 4$ and is $O\mathopen{}\mathclose{{}\left(\binom{n}{k}k^{\beta}\mu/(n-k+1)}\right)$ for $m=2$.
+∎
+Lemma 22.
+Let $k\leq n^{\prime}/2$.
+Consider using the SMS-EMOA with heavy-tailed mutation and with $\mu\geq\overline{M}$ to optimize the $m$OJZJ problem.
+Assume that the current population contains at least one inner Pareto optimum.
+Then after at most $\frac{e\beta}{\beta-1}n\mu(n^{\prime}-2k+1)^{m/2}$ iterations in expectation, all inner Pareto front points are covered.
+Proof.
+We use similar arguments as in the proof of Lemma 7 except for the following probability calculation.
+The probability to generate an offspring with function value equal to the one of an uncovered inner Pareto front point from a neighbor covered by the population is at least $\frac{1}{n\mu}P_{1}^{\beta}\geq\frac{\beta-1}{e\beta n\mu}$, where $P_{1}^{\beta}\geq\frac{\beta-1}{e\beta}$ is from [ZD23d, Lemma 16].
+∎
+Lemma 23.
+Let $k\leq n^{\prime}/2$.
+Consider using the SMS-EMOA with heavy-tailed mutation and with $\mu\geq\overline{M}$ to optimize the $m$OJZJ problem.
+Assume that the current population covers all inner Pareto front points.
+Then after at most $\binom{n}{k}C_{n/2}^{\beta}\mu k^{\beta}((n^{\prime}-2k+3)^{m/2}-(n^{\prime}-2k+1)^{m/2})$ iterations in expectation, the full Pareto front is covered.
+Proof.
+We reuse the arguments of Lemma 8 except for the following probability calculation.
+From Lemma 19, we know that the probability to generate any point in $j$-th level condition on that all $(j-1)$-th level points are covered is at least
+$\frac{1}{\mu}\binom{n}{k}^{-1}P_{k}^{\beta}=\Omega\mathopen{}\mathclose{{}\left(\binom{n}{k}^{-1}\frac{1}{\mu}k^{-\beta}}\right).\qed$
+Considering the whole process for the full coverage of the Pareto front from Lemmas 21 to 23, we have the following runtime result for the SMS-EMOA with heavy-tailed mutation (where we estimated $\binom{n}{k}\leq n^{k}/(k!)$ and $k!\geq\sqrt{2\pi}k^{k+0.5}e^{-k}$).
 Theorem 24.
 Let $k\leq n^{\prime}/2$ and $\beta>1$.
-Consider using SMS-EMOA with $\mu\geq M$ and with heavy-tailed mutation to optimize the $m$OJZJ problem.
-Then after at most $O(M\mu k^{\beta-0.5-k}n^{k})$ iterations or function evaluations in expectation, the full Pareto front is covered.
-Compared to the runtime guarantee of Theorem 8 for the original SMS-EMOA, the guarantee of Theorem 17 above for the SMS-EMOA with heavy-tailed mutation is by a factor of asymptotically $k^{k+0.5-\beta}$ stronger.
-The SMS-EMOA Also Performs Well for Bi-objective Optimization
-The above sections discussed the runtime of the SMS-EMOA for many objectives, and showed that it performs well for many objectives, which is different from the original NSGA-II.
-Since the only theory paper (Bian et al.
-2023) on the SMS-EMOA merely considers its performance on the bi-objective OJZJ problem, to broaden our understanding of this algorithm we now analyze its runtime for the two most prominent bi-objective benchmarks OneMinMax and LOTZ.
-OneMinMax and LOTZ
-The OneMinMax benchmark introduced by Giel and Lehre (2010) (and also the CountingOnesCountingZeroes benchmark proposed by Laumanns et al.
-(2002)) are bi-objective counterparts of the famous single-objective OneMax benchmark.
-Likewise, the LOTZ benchmark defined by Laumanns, Thiele, and Zitzler (2004) (also the weighted version WLPTNO of Qian, Yu, and Zhou (2013)) are bi-objective counterparts of the classic single-objective LeadingOnes benchmark.
+Consider using SMS-EMOA with $\mu\geq\overline{M}$ and with heavy-tailed mutation to optimize the $m$OJZJ problem.
+Then after at most $O(M\mu k^{\beta-0.5-k}(en)^{k})$ iterations or function evaluations in expectation, the full Pareto front is covered.
+Compared to the runtime guarantee of Theorem 9 for the original SMS-EMOA, the guarantee of Theorem 24 above for the SMS-EMOA with heavy-tailed mutation is by a factor of asymptotically $k^{k+0.5-\beta}/e^{k}$ stronger.
+Compared to the SMS-EMOA with stochastic population update (Theorem 17), the speed-up is by a factor of $\min\{k^{k+0.5-\beta}/e^{k},4\mu k^{0.5-\beta}(k/2)^{k}/e^{k-1}\}$, which is super-exponential w.r.t.
+$k$.
+7 Runtime Analysis for Two Classic Bi-objective Benchmarks
+In the sections above, we discussed the runtime of the SMS-EMOA for the many-objective OJZJ benchmark and showed a good performance, different from the original NSGA-II.
+Since the only two previous theory papers [BZLQ23, ZLDD24] on the SMS-EMOA merely considers its performance on the bi-objective OJZJ and DLTB problems, to broaden our understanding of this algorithm, we now analyze its runtime for the two most prominent bi-objective benchmarks, OneMinMax and LOTZ.
+7.1 OneMinMax and LOTZ
+The OneMinMax benchmark introduced by Giel and Lehre [GL10] (and likewise the similar CountingOnesCountingZeroes benchmark previously proposed by Laumanns et al.
+[LTZ+02]) are bi-objective counterparts of the famous single-objective OneMax benchmark.
+Analogously, the LOTZ benchmark defined by Laumanns, Thiele, and Zitzler [LTZ04] (also the weighted version WLPTNO of Qian, Yu, and Zhou [QYZ13]) are bi-objective counterparts of the classic single-objective LeadingOnes benchmark.
 For OneMinMax, the two objectives count the number of ones and zeros in a given bit-string, respectively.
 For LOTZ, the first objective is the number of contiguous ones starting from the first bit position, and the second objective is the number of contiguous zeros starting from the last bit position.
-Both benchmarks with their simple and clear structure facilitate the fundamental theoretical understanding of MOEAs.
-We note that there are other benchmarks used in the mathematical analysis of MOEAs (see, e.g.
-Horoba and Neumann 2008; Brockhoff et al.
-2009; Qian, Tang, and Zhou 2016; Li et al.
-2016; Dang et al.
-2023b), but clearly OneMinMax and LOTZ are the most prominent ones.
-Here are the formal definitions of OneMinMax and LOTZ.
-Definition 18 (Giel and Lehre (2010)).
-Let $n\in\mathbb{N}$ be the problem size. The OneMinMax function :$:\{0,1\}^{n}\rightarrow\mathbb{R}$ is defined by
-$\textsc{OneMinMax}(x)=\left(\sum_{i=1}^{n}(1-x_{i}),\sum_{i=1}^{n}x_{i}\right)$
-for any $x=(x_{1},\dots,x_{n})\in\{0,1\}^{n}$.
-Definition 19 (Laumanns, Thiele, and Zitzler (2004)).
-Let $n\in\mathbb{N}$ be the problem size. The LOTZ function :$:\{0,1\}^{n}\rightarrow\mathbb{R}$ is defined by
-$\textsc{LOTZ}(x)=\left(\sum_{i=1}^{n}\prod_{j=1}^{i}x_{j},\sum_{i=1}^{n}\prod_{j=i}^{n}(1-x_{j})\right)$
-for any $x=(x_{1},\dots,x_{n})\in\{0,1\}^{n}$.
-Runtime
-We now analyze the runtime of the SMS-EMOA on the OneMinMax and LOTZ benchmarks.
-Note that the Pareto front, and more generally the size of any set of pairwise non-dominated solutions, for OneMinMax and LOTZ have sizes at most $n+1$.
-Hence we can use Lemma 4 with $M=n+1$.
+Both benchmarks with their simple and clear structure have greatly facilitated the understanding of how MOEAs work.
+We note that there are other benchmarks used in the mathematical analysis of MOEAs, see, e.g., [HN08, BFH+09, QTZ16, LZZZ16, DOSS23b], but clearly OneMinMax and LOTZ are the most prominent ones.
+Definition 25 ([GL10, LTZ04]).
+Let $n\in\mathbb{N}$ be the problem size. The OneMinMax and LOTZ benchmarks are the functions $\{0,1\}^{n}\rightarrow\mathbb{R}^{2}$ defined by
+$\textsc{OneMinMax}(x)$ $=\mathopen{}\mathclose{{}\left(\sum_{i=1}^{n}(1-x_{i}),\sum_{i=1}^{n}x_{i}}\right)$
+$\textsc{LOTZ}(x)$ $=\mathopen{}\mathclose{{}\left(\sum_{i=1}^{n}\prod_{j=1}^{i}x_{j},\sum_{i=1}^{n}\prod_{j=i}^{n}(1-x_{j})}\right)$
+for all $x=(x_{1},\dots,x_{n})\in\{0,1\}^{n}$.
+7.2 Runtime
+We now prove runtime guarantees for the SMS-EMOA on the OneMinMax and LOTZ benchmarks.
+Note that the Pareto front, and, more generally, the size of any set of incomparable solutions for OneMinMax and LOTZ have sizes at most $n+1$.
+Hence we can use Lemma 4 with $\overline{M}=n+1$.
 Theorem 26 below gives an upper bound for the runtime of SMS-EMOA on OneMinMax.
-This bound (in terms of fitness evaluations) is of the same asymptotic order as the runtime guarantee for the NSGA-II (Zheng, Liu, and Doerr 2022).
-Knowing from Lemma 4 that we cannot lose Pareto points, the proof of our results consists of adding the waiting times for finding a new Pareto point from an already existing neighboring one.
+This bound (in terms of fitness evaluations) is of the same asymptotic order as the runtime guarantee $\frac{2e^{2}}{e-1}\mu n(\ln n+1)$ for the NSGA-II [ZD23a, Theorem 2].
+Due to the more complex population dynamics, the leading constant in the result in [ZD23a] is inferior (58% larger).
+More critically, the result in [ZD23a] was proven for population sizes $\mu\geq 4(n+1)$ only, and it was disproven for $\mu=n+1$ [ZD23a, Theorem 12], whereas our result here holds for all $\mu\geq n+1$.
+Knowing from Lemma 4 that we cannot lose Pareto points, the proof of our result consists of adding the waiting times for finding a new Pareto point from an already existing neighboring one.
 Theorem 26.
@@
 Then after at most $2e\mu n(\ln n+1)$ iterations (or $\mu+2e\mu n(\ln n+1)$ fitness evaluations) in expectation, the full Pareto front is covered.
-Our runtime guarantee for LOTZ, see Theorem 21 below, again coincides with the guarantee for the NSGA-II (Zheng, Liu, and Doerr 2022).
-In the proof, the process is divided into two stages.
-The first stage is to cover a Pareto front point for the first time and the second stage is to cover the full Pareto front starting from one covered Pareto front point.
-For the first stage, we pessimistically consider the time to reach $1^{n}$.
-For the second stage, we consider the time to generate $01^{i}0^{n-i},i=n-1,\dots,$ 0 one after the other.
-In both stages, we heavily rely on Lemma 4 asserting that important progress is not lost, which allows to add waiting times of successive progress events.
+Proof.
+Since each search point in $\{0,1\}^{n}$ is Pareto optimal, from Lemma 4, we know that any Pareto front point covered by the population $P_{t}$ will be maintained in all future generations.
+Assume that for some $i\in[0..n]$, the population $P_{t}$ contains a solution with objective value $(i,n-i)$.
+Then the probability to reach a solution with value $(i+1,n-i-1)$ is at least
+$\frac{1}{\mu}\frac{n-i}{n}\mathopen{}\mathclose{{}\left(1-\frac{1}{n}}\right)^{n-1}\geq\frac{n-i}{e\mu n},$
+and the probability to reach a solution with function value $(i-1,n-i+1)$ is at least
+$\frac{1}{\mu}\frac{i}{n}\mathopen{}\mathclose{{}\left(1-\frac{1}{n}}\right)^{n-1}\geq\frac{i}{e\mu n}.$
+Thus the expected numbers of iterations to cover $(i+1,n-i-1)$ and $(i-1,n-i+1)$ are at most $\frac{e\mu n}{n-i}$ and at most $\frac{e\mu n}{i}$, respectively.
+Hence, the expected number of iterations to cover the full Pareto front is
+$\Bigg{(}\sum_{j=i}^{1}\frac{e\mu n}{n-j}\Bigg{)}+\mathopen{}\mathclose{{}\left(\sum_{j=i}^{n-1}\frac{e\mu n}{j}}\right)\leq 2\sum_{j=1}^{n-1}\frac{e\mu n}{j}\leq 2e\mu n(\ln n+1).\qed$
+We now turn to the LOTZ benchmark.
+Our runtime guarantee, see Theorem 27 below, is again of the same asymptotic order of magnitude as the guarantee $\frac{2e^{2}}{e-1}\mu n^{2}$ proven for the NSGA-II [ZD23a, Theorem 8], and again, the guarantee for the NSGA-II has a weaker leading constant and requires a larger population size of $\mu\geq 4(n+1)$.
+In the proof, we first regard the time to reach the individual $1^{n}$, and then consider the times to generate the individuals $01^{i}0^{n-i},i=n-1,\dots,$ 0 one after the other.
+In both stages, we heavily rely on Lemma 4 asserting that important progress is not lost.
 Theorem 27.
 Consider using the SMS-EMOA with $\mu\geq n+1$ to optimize the LOTZ problem with problem size $n$.
-Then after at most $2e\mu n^{2}$ iterations (or fitness evaluations) in expectation, the full Pareto front is covered.
-Note that the runtimes of the GSEMO on OneMinMax and LOTZ are $O(n^{2}\ln n)$ and $O(n^{3})$, respectively (Giel and Lehre 2010; Laumanns, Thiele, and Zitzler 2004).
-Hence, they agree with our bounds for the SMS-EMOA when using the most interesting population size $\mu=\Theta(n)$.
-Conclusion
+Then after at most $2e\mu n^{2}$ iterations (or $\mu+2e\mu n^{2}$ fitness evaluations) in expectation, the full Pareto front is covered.
+Proof.
+Denote by $f$ the LOTZ function.
+We first consider the time to reach $1^{n}$.
+Let $X_{t}$ denote the maximum $f_{1}$-value of an individual in $P_{t}$.
+Assume that at some $t$, we have $X_{t}<n$.
+Let $x\in P_{t}$ be an individual with $f_{1}(x)=X_{t}$.
+Then with probability at least $\frac{1}{\mu}(1-\frac{1}{n})^{n-1}\frac{1}{n}$, this $x$ is chosen as parent and mutated to an individual $y$ with $f_{1}(y)>X_{t}$.
+By Lemma 4, $P_{t+1}$ contains an individual $z$ with $z\succeq y$.
+In particular, we have $f_{1}(z)\geq f_{1}(y)>X_{t}$ and hence $X_{t+1}>X_{t}$.
+We have just seen that the probability to increase $X_{t}$ by at least one is at least $\frac{1}{\mu}(1-\frac{1}{n})^{n-1}\frac{1}{n}\geq\frac{1}{e\mu n}$.
+Consequently, the expected waiting time for this event is at most $e\mu n$.
+After at most $n$ such increases, hence after an expected time of at most $e\mu n^{2}$, we have reached the state $X_{t}=n$, that is, $1^{n}\in P_{t}$.
+Once $1^{n}$ is in the population, we consider the time to reach the Pareto optima $1^{n-1}0,\dots,10^{n-1},0^{n}$ one by one.
+Assume that $1^{i}0^{n-i}$ with $i\in[1..n]$ is in the population.
+Then to reach $1^{i-1}0^{n-i+1}$, we only need to choose $1^{i}0^{n-i}$ as parent and flip the $i$-th bit from $1$ to 00, with the other bits unchanged.
+This happens with probability
+$\frac{1}{\mu}\mathopen{}\mathclose{{}\left(1-\frac{1}{n}}\right)^{n-1}\frac{1}{n}\geq\frac{1}{e\mu n}.$
+From Lemma 4, we know that any Pareto front point will be maintained in all future generations once reached.
+Hence, $1^{i}0^{n-i}\in P_{\tau}$ for all $\tau\geq t$.
+Thus, the expected number of iterations to reach $1^{i-1}0^{n-i+1}$ is at most $e\mu n$.
+Since we have at most $n$ such Pareto optima to reach, we know that the expected number of iterations to cover all remaining Pareto front points is at most $e\mu n^{2}$.
+In summary, the full Pareto front will be covered in at most $2e\mu n^{2}$ iterations in expectation. ∎
+8 Conclusion
 Motivated by the observed difficulty of the NSGA-II for many objectives, this paper resorted to the SMS-EMOA, a variant of the steady-state NSGA-II, and proved that, different from the NSGA-II, it efficiently solves the $m$OJZJ problem.
-Noting that the SMS-EMOA also employs non-dominated sorting, but replaces the crowding distance with the hypervolume, this result together with the one of Wietheger and Doerr (2023) supports the conclusion that non-dominated sorting is a good building block for MOEAs, but the crowding distance has deficiencies for more than two objectives.
-We also showed that the stochastic population update proposed in (Bian et al.
-2023) for the bi-objective SMS-EMOA becomes less effective for many objectives.
+Noting that the SMS-EMOA also employs non-dominated sorting, but replaces the crowding distance with the hypervolume, this result together with the one of Wietheger and Doerr [WD23] for the NSGA-III supports our conclusion that non-dominated sorting is a good building block for MOEAs, but the crowding distance has deficiencies for more than two objectives.
+We also showed that the stochastic population update proposed in [BZLQ23] for the bi-objective SMS-EMOA becomes less effective for many objectives.
 All these results in a very rigorous manner support the general knowledge that multi-objective optimization becomes increasingly harder with growing numbers of objectives.
-On the positive side, we showed that the advantages of heavy-tailed mutation, previously observed in single- and bi-objective optimization, remain when increasing the number of objectives.
+On the positive side, we showed that the advantages of heavy-tailed mutation, previously observed in single- and bi-objective optimization, remain unchanged when increasing the number of objectives.
 Given that our first result shows advantages of the SMS-EMOA and there is only one previous work performing rigorous runtime analyses for this algorithm, we extended our knowledge in this direction by proving competitive runtime guarantees for this algorithm on the two most prominent bi-objective benchmarks.
+They show that the SMS-EMOA performs here at least at well as the NSGA-II.
+Subsequent Work
+In [WD24], the runtime of several MOEAs on many-objective problems was analyzed.
+Concerning the runtime of the classic SMS-EMOA on the $m$OJZJ benchmark, our guarantee of $O(\mu Mn^{k})$ was improved to $O(\mu mn^{k})$.
+To prove this result, our general technical lemma for the survival of individuals in the population (Lemma 4) was used.
+The improvement then stemmed from detecting a faster way how all the MOEAs regarded in that work discover the Pareto front.
+The work [WD24] did not regard the SMS-EMOA with stochastic population update and with heavy-tailed mutation.
+Since both shorten the time to find an additional extremal block (see the proof of Lemma 8), it is clear that analogous improvements are possible for the SMS-EMOA with stochastic population update and heavy-tailed mutation, giving the bounds $O(\mu mn^{k}\min\{1,\mu/2^{k}\})$ and $O(\mu mk^{\beta-0.5-k}n^{k})$ respectively.
 Acknowledgments